### v3

# Towards An Analytical Framework for Dynamic Potential Games

Abstract. Potential game is an emerging notion and framework for studying $N$-player games, especially with heterogeneous players. In this paper, we build an analytical framework for dynamic potential games. We prove that a game is a dynamic potential game if and only if each player’s value function can be decomposed as a potential function and a residual term which is solely dependent on other players’ policies. This decomposition is consistent with the result in the static setting and enables us to identify and analyze an important and new class of dynamic potential games called the distributed game. Moreover, we prove that a game is a dynamic potential game if the value function has a symmetric Jacobian. This generalizes the differential characterization for static potential games by replacing the classical derivative with a new notation of functional derivative with respect to Markov policies. For a general class of continuous-time stochastic games, we explicitly characterize their potential functions. In particular, we show that the potential function of linear-quadratic games can be studied through a system of linear ODEs. Furthermore, under a rank condition on control coefficients, we prove a linear-quadratic game is a Markov potential game if and only if all players have identical cost functions.
Key words. Dynamic potential game, closed-loop Nash equilibrium, linear derivative, distributed game, stochastic differential game, linear-quadratic potential game
AMS subject classifications. 91A14, 91A06, 91A15
1 Introduction
General-sum $N$-player games are of paramount importance in various fields, spanning economics, engineering, biology, and ecology (see e.g., [17, 5]). These games model strategic interactions among multiple players, where each player aims to optimize her individual objective based on the observed system state and actions of other players. $N$-player games are notoriously difficult to analyze.
One well-established paradigm for analyzing $N$-player games is the mean field framework [8, 28]. Its ingenious yet simple aggregation ideas enables efficiently approximating the Nash equilibrium of games with a large number of players. However, such an approximation typically requires that players are (approximately) homogeneous who interact weakly and symmetrically. As a result, the mean field approach is not appropriate for general games with heterogeneous players and/or with general forms of interactions.
An alternative and emerging framework for studying $N$-player games, especially with heterogeneous players, is the potential game introduced by [37]. A game is a (static) potential game if there exists a potential function such that whenever one player unilaterally deviates from their action, the change of the potential function equals to the change of that player’s value function. Once the potential function is identified, the daunting task of finding Nash equilibrium of the game is simplified to an optimization problem of finding the global optimum of the potential function.
This introduction of potential functions for static games has since been adopted naturally for dynamic games. The key distinctions between dynamic potential games and static ones lie in the dependence of the potential function on the set of admissible polices, and the interconnection and interaction among all players through the state dynamics. This makes constructing a potential function or verifying its existence in dynamic games significantly more challenging. Indeed, [29] shows that a dynamic game where the cost at each state forms a static potential game may not be a dynamic potential game.
One of the main culprits for the slow progress on dynamic potential games is the lack of appropriate analytical tools and limited understanding of the game structure for a dynamic potential game. The vast majority of existing literature on dynamic potential games focuses on algorithms for computing Nash equilibria of dynamic potential games [35, 34, 31, 33, 44, 49, 12, 16, 29, 32, 38, 48]. These algorithmic studies are, however, built without verifying a priori whether the game under consideration is actually a dynamic potential game. In discrete-time games, separability of value functions has been identified in [29]. With this separability property, one class of dynamic games known as team Markov games has been analyzed and remains the primary example for potential games studies [30, 46].
Our work.
This paper focuses on closed-loop dynamic potential games, which have a wide range of applications including economic modelling and analysis of reinforcement learning algorithms. In closed-loop dynamic games, players optimize their objectives over functions of the time and the system state, known as policies. This is in contrast to open-loop dynamic potential games [14, 15], where the strategies are functions of the time and system noise.
  • We prove that a game is a dynamic potential game if and only if each player’s value function can be decomposed in two components: a potential function and a residual term solely dependent on other players’ policies (Theorem 2.2). It generalizes similar results for static games [37, 45] and discrete Markov games [29] to general dynamic games with arbitrary state dynamics and policy classes.
  • We apply this decomposition characterization to identify a new and important class of dynamic potential games called distributed games (Section 3). In these games, each player’s state and polices are decoupled, allowing for the explicit construction of dynamic potential functions based on the games’ cost functions (Theorems 3.1 and 3.2). This characterization applies to distributed games with measurable coefficients and heterogeneous players. It also underscores the critical distinction between static and dynamic potential games, even in this decoupled setting. Specifically, constructing dynamic potential functions requires considering static potential games with enlarged action sets that include players’ state variables as additional actions (Remark 3.1).
    To the best of our knowledge, this is the first systematic characterization of potential functions for continuous-time distributed games with heterogeneous players. It enables the analysis of closed-loop Nash equilibria in these (previously intractable) large-scale games using stochastic control techniques, as detailed in Remark 3.1.
  • We introduce a notion of linear derivative to characterize the sensitivity of the value function with respect to unilateral deviation of policies (Definition 4.1). Leveraging this concept, we prove that an arbitrary dynamic game is a dynamic potential game if the value functions have a symmetric Jacobian formed by their second-order derivatives (Theorem 4.3). Moreover, we construct the potential function using first-order derivatives of value functions.
    The notation of linear derivatives is crucial for extending the differential characterization, originally established for static potential games [37, 45, 2], to dynamic games with continuous state and action spaces and infinite-dimensional policy classes. It addresses the challenge of defining derivatives over policy classes lacking a natural metric structure; see Remark 4.2 for details.
  • For continuous-time stochastic games, where the state dynamics involves controlled drift and diffusion coefficients, we establish the linear differentiability of value functions and characterize the dynamic potential function via two approaches: one probabilistic approach where the potential function is represented via sensitivity processes of the state and control processes with respect to policies, and another where the potential function is expressed via a system of linear partial differential equations (PDEs). The probabilistic approach applies to games with twice differentiable policies and model coefficients (see Remark 4.3), while the PDE approach accommodates more general forms of policies and coefficients, such as Hölder continuous state coefficients and policies (see Remark 4.4).
  • We show that the probabilistic and PDE characterizations are consistent for linear-quadratic (LQ) games, and reduce to a system of ordinary differential equations (Theorem 4.6). In particular, we prove that under a rank condition on control coefficients, a linear-quadratic game is a Markov potential game if and only if all players have identical cost functions (Proposition 4.8).
    To the best of our knowledge, this is the first necessary and sufficient condition for multi-dimensional LQ Markov potential games. It implies that a strong coupling between the players’ actions requires the homogeneity of cost functions for a potential game. It also suggests that the strength of interactions among players through dynamics and policies is crucial in determining whether a dynamic game is a potential game. These are intriguing insights about the structure of dynamic potential games (Remark 4.5).
Our approach.
One of the key challenges to develop the general analytical framework for the dynamic potential game is to develop an appropriate notion of the functional derivative with respect to polices for possibly continuous time/state spaces and general policy classes. For instance, in the discrete time setting of [29, 22] with finite-dimensional parameterizations for policies, Fréchet derivatives are adopted to show that the game is a potential game if the second-order derivatives of the value functions are symmetric. However, neither Fréchet derivative nor parametrization is appropriate for the general framework: the set of admissible policies becomes infinite-dimensional and may not be a normed space; for games with stochastic policies, the admissible polices take values in the space of probability measures and hence do not form a vector space. To overcome the above difficulties, we define the derivative of value function along a convex combination of two policies. Compared with the Fréchet derivative, such a directional derivative is easier to compute and well-defined under weaker conditions. It also eliminates the need to select a norm over the policy class, providing more flexibility in the analysis (see Remark 4.2 for more details).
In contrast to the mean field approach and its graphon extensions, which either require homogeneity in players’ state dynamics and symmetric interaction among players and policy classes or at least some specific asymptotic structures of the cost functions (see e.g., [19, 3, 26, 6]), the potential game framework offers a new perspective on constructing (closed-loop) Nash equilibrium for finite-player games that is non-asymptotic in nature. It also accommodates heterogeneity among players to allow for different action sets and state dynamics, different dependencies on individual behavior, and different interaction strengths between pairs of players (see Theorem 3.2 and Proposition 3.2). We emphasize that our proposed framework characterizes closed-loop dynamic potential games with a finite number of heterogeneous players. This contrasts with potential mean field games, which involve infinitely many homogeneous players and are associated with McKean-Vlasov control problems (see, for example, [4, 11, 18, 9] and their references for more details).
For instance, in Section 3, we construct potential functions for closed-loop $N$-player distributed games, arising from multi-vehicle coordination [1], energy market [40], crowd motion modelling [4, 9], and interacting particle models in biology [39] and physics [9]. The results accommodate players with heterogeneous characteristics and interactions, facilitating the construction of distributed closed-loop Nash equilibria in large-scale games using stochastic control techniques.
Notation.
For each $n\in{\mathbb{N}}$, we denote by ${\mathbb{S}}^{n}$ the space of $n\times n$ symmetric matrices. For each $T>0$, $n\in{\mathbb{N}}$, $\alpha\in(0,1]$, probability space $(\Omega,\mathcal{F},\mathbb{P})$ and Euclidean space $(E,|\cdot|)$, we introduce the following spaces:
  • $\mathcal{S}^{\infty}([t,T];E)$, $t\in[0,T]$, is the space of measurable processes $X:\Omega\times[t,T]\rightarrow E$ such that ${\mathbb{E}}[\sup_{s\in[t,T]}|X_{s}|^{q}]<\infty$ for all $q\geq 1$;
  • $\mathcal{H}^{\infty}([t,T];E)$, $t\in[0,T]$, is the space of measurable processes $X:\Omega\times[t,T]\rightarrow{\mathbb{R}}^{n}$ such that ${\mathbb{E}}[\int_{t}^{T}|X_{s}|^{q}{\mathrm{d}}s]<\infty$ for all $q\geq 1$;
  • $C^{i+\alpha/2,j+\alpha}([0,T]\times{\mathbb{R}}^{n};E)$, $i,j\in{\mathbb{N}}\cup\{0\}$, is the space of functions $u:[0,T]\times{\mathbb{R}}^{n}\rightarrow E$ such that $\|u\|_{{i+\alpha/2,j+\alpha}}\coloneqq\sum_{\ell=0}^{i}\|\partial^{\ell}_{t}u\|_{0}+\sum_{\ell=0}^{j}\|\partial^{\ell}_{x}u\|_{0}+[\partial^{i}_{t}u]_{\alpha/2}+[\partial^{j}_{x}u]_{\alpha}<\infty$, where $\|u\|_{0}\coloneqq\sup_{(t,x)\in[0,T]\times{\mathbb{R}}^{n}}|u(t,x)|$ and $[u]_{\alpha}\coloneqq\sup_{(t,x),(t,x^{\prime})\in[0,T]\times{\mathbb{R}}^{n}}\frac{|u(t,x)-u(t^{\prime},x^{\prime})|}{(|t-t^{\prime}|^{1/2}+|x-x^{\prime}|)^{\alpha}}$. Similar definitions extend to $C^{i+\alpha/2,j+\alpha,k+\alpha}([0,T]\times{\mathbb{R}}^{n}\times{\mathbb{R}}^{m};E)$ for functions with two space variables.
Throughout the paper, unless otherwise stated, proofs of theorems are deferred to Section 5.
2 (Dynamic) Potential Game and Its Nash Equilibrium
To introduce the mathematical framework for dynamic potential games, let us start by some basic notions for the game and associated policies. Consider a game $\mathcal{G}=(I_{N},\mathcal{S},(A_{i})_{i\in I_{N}},\pi^{(N)},(V_{i})_{i\in I_{N}})$, where $I_{N}=\{1,\ldots,N\}$, $N\in\mathbb{N}$, is a finite set of players; $\mathcal{S}$ is a topological space representing the state space of the underlying dynamics; $A_{i}$ is a Borel subset of a topological vector space representing the action set of player $i$; $\pi^{(N)}=\prod_{i\in I_{N}}\pi_{i}$ is the set of joint policy profiles of all players, where $\pi_{i}$ is a set of Borel measurable functions $\phi_{i}:\mathcal{S}\rightarrow A_{i}$ representing the admissible policies of player $i$; and $V_{i}:\mathcal{S}\times\pi^{(N)}\rightarrow{\mathbb{R}}$ is the value function of player $i$, where $V_{i}^{s}(\phi)\coloneqq V_{i}(s,\phi)$ is player $i$’s expect cost if the state dynamics starts at the state $s\in\mathcal{S}$ and all players take the policy profile $\phi\in\pi^{(N)}$. For each $i\in I_{N}$, we denote by $\pi^{(N)}_{-i}=\prod_{j\in I_{N}\setminus\{i\}}\pi_{j}$ the sets of policy profiles of all players except player $i$. The elements of $\pi^{(N)}$ and $\pi^{(N)}_{-i}$ are denoted as ${\phi}=(\phi_{i})_{i\in I_{N}}$, and $\phi_{-i}=(\phi_{j})_{j\in I_{N}\setminus\{i\}}$, respectively.
This game includes discrete-time and continuous-time dynamic games; and its value function $V_{i}^{s}(\phi)$ may depend on the underlying state system, whose time and space variables are referred collectively as the state variable. Note that one may simply view the static game as a stateless dynamic game.
The focus of this paper is on a class of dynamic games called dynamic potential games, defined through two related yet distinct notions of Nash equilibrium: closed-loop and Markovian types.
Definition 2.1 (Nash equilibrium).
A policy profile $\phi\in\pi^{(N)}$ is a closed-loop Nash equilibrium for $\mathcal{G}$ with initial state $s_{0}\in\mathcal{S}$ if
$V^{s_{0}}_{i}((\phi_{i},\phi_{-i}))\leq V^{s_{0}}_{i}((\phi^{\prime}_{i},\phi_{-i}))\quad\forall i\in I_{N},\phi^{\prime}_{i}\in\pi_{i}.$
A policy profile $\phi\in\pi^{(N)}$ is a Markov Nash equilibrium for $\mathcal{G}$ if
$V^{s}_{i}((\phi_{i},\phi_{-i}))\leq V^{s}_{i}((\phi^{\prime}_{i},\phi_{-i}))\quad\forall s\in\mathcal{S},i\in I_{N},\phi^{\prime}_{i}\in\pi_{i}.$
Definition 2.1 is consistent with the concept of closed-loop and Markov equilibrium introduced in stochastic differential games, as described in [10]. Note that a closed-loop equilibrium $\phi$ can depend on the initial state $s_{0}$ of the system dynamics, while Markov equilibrium requires that the policy profile $\phi$ be a closed-loop Nash equilibrium for all initial states.
Definition 2.2 (Dynamic potential game).
A dynamic game $\mathcal{G}$ with initial state $s_{0}\in\mathcal{S}$ is a closed-loop potential game (CLPG), if there exists $\Phi^{s_{0}}:\pi^{(N)}\rightarrow{\mathbb{R}}$, called a potential function, such that for all $i\in I_{N}$, $\phi_{i},\phi_{i}^{\prime}\in\pi_{i}$ and $\phi_{-i}\in\pi^{(N)}_{-i}$,
$\Phi^{s_{0}}((\phi^{\prime}_{i},\phi_{-i}))-\Phi^{s_{0}}((\phi_{i},\phi_{-i}))=V^{s_{0}}_{i}((\phi^{\prime}_{i},\phi_{-i}))-V^{s_{0}}_{i}((\phi_{i},\phi_{-i})).$ (2.1)
A game $\mathcal{G}$ is a Markov potential game (MPG), if there exists a potential function $\Phi:\mathcal{S}\times\pi^{(N)}\rightarrow{\mathbb{R}}$ such that for all $s_{0}\in\mathcal{S}$, $\Phi(s_{0},\cdot)$ is a potential function of $\mathcal{G}$ with initial state $s_{0}\in\mathcal{S}$.
Intuitively, a game $\mathcal{G}$ is a dynamic potential game if there exists a potential function such that whenever one player unilaterally deviates from their policy, the change of the potential function equals to the change of that player’s value function. Definition 2.2 enables finding a Nash equilibrium of a potential game by optimizing its potential function, as shown in the following proposition.
Proposition 2.1.
If there exists $s_{0}\in\mathcal{S}$ such that $\mathcal{G}$ with initial state $s_{0}$ is a CLPG with potential function $\Phi^{s_{0}}$, then any $\phi^{*}\in\operatorname*{arg\,min}_{\phi\in\pi^{(N)}}\Phi^{s_{0}}(\phi)$ is a closed-loop Nash equilibrium of $\mathcal{G}$ with initial state $s_{0}$. Similarly, if $\mathcal{G}$ is an MPG with potential function $\Phi$ and $\phi^{*}\in\pi^{(N)}$ satisfies $\phi^{*}\in\operatorname*{arg\,min}_{\phi\in\pi^{(N)}}\Phi(s,\phi)$ for all $s\in\mathcal{S}$, then $\phi^{*}$ is an Markov Nash equilibrium of $\mathcal{G}$.
Proposition 2.1 shows the importance of an explicit criterion for the existence or the construction of a potential function for a given potential game. The most known example of MPG is the team Markov game, where the potential function corresponds to the players’ value function.
Example 2.1 (Team Markov Game).
This is a game where all players have a common interest. That is, in this game $\mathcal{G}=(I_{N},\mathcal{S},(A_{i})_{i\in I_{N}},\pi^{(N)},(V_{i})_{i\in I_{N}})$, $V_{i}=V_{j}$ for all $i,j\in I_{N}$; see [30, 46]. In this setting, $\mathcal{G}$ is an MPG with a potential function $\Phi\equiv V_{1}$.
Moreover, a game is a potential game if and only if the value function of each player can be decomposed into two terms: one that is common for all players, namely the potential function, and the other that may be different for each player but depends only on the polices of other players.
Theorem 2.2.
A game $\mathcal{G}=(I_{N},\mathcal{S},(A_{i})_{i\in I_{N}},\pi^{(N)},(V_{i})_{i\in I_{N}})$ with initial state $s_{0}\in\mathcal{S}$ is a CLPG with potential $\Phi^{s_{0}}$ if and only if for all $i\in I_{N}$, there exists $U^{s_{0}}_{i}:\pi^{(N)}_{-i}\rightarrow{\mathbb{R}}$ such that
$V^{s_{0}}_{i}((\phi_{i},\phi_{-i}))=\Phi^{s_{0}}((\phi_{i},\phi_{-i}))+U^{s_{0}}_{i}(\phi_{-i}),\quad\forall\phi_{i}\in\pi_{i},\phi_{-i}\in\pi^{(N)}_{-i}.$ (2.2)
Furthermore, a game $\mathcal{G}$ is an MPG if and only if (2.2) holds for all $s_{0}\in S$ and $i\in I_{N}$.
Remark 2.1 (Dynamic vs static potential games).
The definition of dynamic potential game and the subsequent Proposition 2.1 are consistent with their respective counterparts for static potential games. Theorem 2.2 generalizes results from [45] for static games and [29] for discrete Markov games with stationary policies to general dynamic games with arbitrary state dynamics and policy classes.
The key distinctions between dynamic potential games and static ones lie in the dependence of the potential function on the set of admissible policy profiles, and the nonlinear relationship between value functions and policies through the state dynamics. This makes identifying a potential function in dynamic games, if one exists, significantly more challenging. The broader the class of admissible policy profiles $\pi^{(N)}$, the stricter the requirements on system dynamics and payoff functions to guarantee the existence of a potential function. Deriving and decomposing value functions for dynamic games, as in Theorem 2.2, is more challenging than for static ones, primarily due to the interconnection and interaction among all players that arise through the state dynamics.
In fact, a dynamic game in which the cost at each state forms a static potential game may not necessarily be an MPG [29]. As a result, there are few known examples of MPGs beyond the team Markov game. In the next section, we will identify and analyze an important (new) class of dynamic potential games called distributed games.
3 Distributed Game as Markov Potential Game
In a distributed game, each player’s control depends only on the state of herself, although her value function may depend on the joint state and policy profiles of all players. More precisely,
Definition 3.1 (Distributed game $\mathcal{G}_{\rm dist}$).
Let $T>0$, $N\in{\mathbb{N}}$, $(n_{i})_{i=1}^{N},(k_{i})_{i=1}^{N}\subset{\mathbb{N}}$. The dynamic game $\mathcal{G}_{\rm dist}=(I_{N},\mathcal{S},(A_{i})_{i\in I_{N}},{\pi}^{(N)}_{\rm dist},(V_{i})_{i\in I_{N}})$ with $I_{N}=\{1,\ldots,N\}$, $\mathcal{S}=[0,T]\times\prod_{i\in I_{N}}{\mathbb{R}}^{n_{i}}$ and $A_{i}$ is a Borel set in ${\mathbb{R}}^{k_{i}}$. Here the set ${\pi}^{(N)}_{\rm dist}=\prod_{i\in I_{N}}{\pi}^{\rm dist}_{i}$ contains distributed policy profiles, where ${\pi}^{\rm dist}_{i}$ contains measurable functions $\phi_{i}:[0,T]\times{\mathbb{R}}^{n_{i}}\rightarrow A_{i}$ such that $\sup_{(t,x_{i})\in[0,T]\times{\mathbb{R}}^{n_{i}}}\frac{|\phi_{i}(t,x_{i})|}{1+|x_{i}|}<\infty$, and for all $(t,x_{i})\in[0,T]\times{\mathbb{R}}^{n_{i}}$, player $i$’s state dynamics
${\mathrm{d}}X^{i}_{s}=b_{i}(s,X^{i}_{s},\phi_{i}(s,X^{i}_{s})){\mathrm{d}}s+\sigma_{i}(s,X^{i}_{s},\phi_{i}(s,X^{i}_{s})){\mathrm{d}}W_{s},\quad s\in[t,T];\quad X^{i}_{t}=x_{i},$ (3.1)
admits a unique square-integrable weak solution $X^{t,x_{i},\phi_{i}}$ on a probability space $(\Omega,\mathcal{F},\mathbb{P})$. Here $(b_{i},\sigma_{i}):[0,T]\times{\mathbb{R}}^{n_{i}}\times{\mathbb{R}}^{k_{i}}\rightarrow{\mathbb{R}}^{n_{i}}\times{\mathbb{R}}^{n_{i}\times n_{w}}$ are given measurable functions, and $W$ is an $n_{w}$-dimensional Brownian motion on $(\Omega,\mathcal{F},\mathbb{P})$. Let $n_{x}=\sum_{i=1}^{N}n_{i}$ and $n_{a}=\sum_{i=1}^{N}k_{i}$, and for each $t\in[0,T]$, $x=(x_{i})_{i\in I_{N}}\in\prod_{i\in I_{N}}{\mathbb{R}}^{n_{i}}$ and $\phi=(\phi_{i})_{i\in I_{N}}\in{\pi}^{(N)}_{\rm dist}$, let $X^{t,x,\phi}=(X^{t,x_{i},\phi_{i}})_{i\in I_{N}}$ be the joint state process, and let $\alpha^{t,x,\phi}=(\phi_{i}(\cdot,X^{t,x_{i},\phi_{i}}))_{i\in I_{N}}$ be the joint control process. Then the player $i$’s value function $V_{i}:[0,T]\times{\mathbb{R}}^{n_{x}}\times{\pi}^{(N)}_{\rm dist}\rightarrow{\mathbb{R}}$ is given by
$V^{t,x}_{i}(\phi)={\mathbb{E}}^{\mathbb{P}}\left[\int_{t}^{T}f_{i}(s,X^{t,x,\phi}_{s},\alpha^{t,x,\phi}_{s}){\mathrm{d}}s+g_{i}(X^{t,x,\phi}_{T})\right],$ (3.2)
where $f_{i}:[0,T]\times{\mathbb{R}}^{n_{x}}\times{\mathbb{R}}^{n_{a}}\rightarrow{\mathbb{R}}$ is a measurable function such that $\sup_{(t,x,a)\in[0,T]\times{\mathbb{R}}^{n_{x}}\times{\mathbb{R}}^{n_{a}}}\frac{|f_{i}(t,x,a)|}{1+|x|^{2}+|a|^{2}}<\infty$ and $g_{i}:{\mathbb{R}}^{n_{x}}\rightarrow{\mathbb{R}}$ is a measurable function such that $\sup_{x\in{\mathbb{R}}^{n_{x}}}\frac{|g_{i}(x)|}{1+|x|^{2}}<\infty$.
A distributed game $\mathcal{G}_{\rm dist}$ is an MPG if its cost functions admit a potential structure, as stated below.
Theorem 3.1.
Take the game $\mathcal{G}_{\rm dist}$ in Definition 3.1. Suppose that there exist measurable functions $F:[0,T]\times{\mathbb{R}}^{n_{x}}\times{\mathbb{R}}^{n_{a}}\rightarrow{\mathbb{R}}$, $G:{\mathbb{R}}^{n_{x}}\rightarrow{\mathbb{R}}$, $U_{f_{i}}:[0,T]\times\prod_{j\not=i}{\mathbb{R}}^{n_{j}}\times\prod_{j\not=i}{\mathbb{R}}^{k_{j}}\rightarrow{\mathbb{R}}$ and $U_{g_{i}}:\prod_{j\not=i}{\mathbb{R}}^{n_{j}}\rightarrow{\mathbb{R}}$, $i\in I_{N}$, such that for all $i\in I_{N}$ and $(t,x,a)\in[0,T]\times{\mathbb{R}}^{n_{x}}\times{\mathbb{R}}^{n_{a}}$,
$f_{i}(t,x,a)=F(t,x,a)+U_{f_{i}}(t,x_{-i},a_{-i}),\quad g_{i}(x)=G(x)+U_{g_{i}}(x_{-i})\quad$ (3.3)
with $x_{-i}=(x_{j})_{j\not=i}$ and $a_{-i}=(a_{j})_{j\not=i}$. Then the distributed game $\mathcal{G}_{\rm dist}$ is an MPG, with a potential function $\Phi:[0,T]\times{\mathbb{R}}^{n_{x}}\times{\pi}^{(N)}_{\rm dist}\rightarrow{\mathbb{R}}$ defined by
$\Phi^{t,x}(\phi)={\mathbb{E}}^{\mathbb{P}}\left[\int_{t}^{T}F(s,X^{t,x,\phi}_{s},\alpha^{t,x,\phi}_{s}){\mathrm{d}}s+G(X^{t,x,\phi}_{T})\right],$ (3.4)
with $X^{t,x,\phi}$ and $\alpha^{t,x,\phi}$ defined in (3.2).
Remark 3.1.
A few remarks regarding Theorem 3.1 are in place.
First, to the best of our knowledge, it is the first characterization of potential functions for general dynamic distributed games; and it applies to general distributed games with measurable policies and model coefficients defined by (3.1) and (3.2). For instance, it enables characterizing potential functions for dynamic games in crowd motion modeling, as shown in Example 3.1 below.
Secondly, it highlights analytically the crucial distinction between static and dynamic potential games. Indeed, to construct the functions $F$ and $G$ for a dynamic potential games, it is essential to construct associated static potential games by enlarging action sets with players’ state variables. In particular, it is sufficient to construct a static potential game with a potential function $F(t,\cdot)$ according to (3.3) with $f_{i}(t,\cdot)$ being player $i$’s value function and ${\mathbb{R}}^{n_{i}}\times{\mathbb{R}}^{k_{i}}$ player $i$’s action set, as well as another static potential game with a potential function $G$ for which $g_{i}$ is player $i$’s value function and ${\mathbb{R}}^{n_{i}}$ is player $i$’s action set.
Moreover, it allows for computing Nash equilibria of a distributed game with heterogeneous players using stochastic control techniques. As a special case of Theorem 2.1, finding an equilibrium of the game $\mathcal{G}_{\rm dist}$ can be translated to a distributed control problem, whose minimizer can be constructed using a stochastic maximum principle and a Hamilton-Jacobi-Bellman equation; see [24] for the special case where the state dynamics (3.1) has coefficients $b_{i}(t,x_{i},a_{i})=a_{i}$ and $\sigma_{i}(t,x_{i},a_{i})\equiv\sigma$.
Finally, it enables employing existing numerical algorithms for high-dimensional control problems (see [20] for a comprehensive survey) to minimize the potential function and compute equilibria for the game $\mathcal{G}_{\rm dist}$. The convergence of these algorithms can be ensured by adapting techniques for general control problems to the potential function $\Phi$ (see e.g., [21, 43]). Such equilibrium computations for distributed games via the potential function will be an interesting future research.
Example 3.1 (Crowd motion).
Take the game $\mathcal{G}_{\rm dist}$ in Definition 3.1. Suppose that for all $i,j\in I_{N}$, $n_{i}=n$ and there exists $c_{i}:[0,T]\times{\mathbb{R}}^{n}\times{\mathbb{R}}^{k_{i}}\rightarrow{\mathbb{R}}$, $h_{ij}:[0,T]\times{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ and $\ell_{i},r_{ij}:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ such that for all $(t,x,a)\in[0,T]\times{\mathbb{R}}^{n_{x}}\times{\mathbb{R}}^{n_{a}}$,
$f_{i}(t,x,a)=c_{i}(t,x_{i},a_{i})+\sum_{j\in I_{N}\setminus\{i\}}h_{ij}(t,x_{i}-x_{j}),\quad g_{i}(x)=\ell_{i}(x_{i})+\sum_{j\in I_{N}\setminus\{i\}}r_{ij}(x_{i}-x_{j}),$ (3.5)
and $h_{ij}(t,x)=h_{ji}(t,-x)$ and $r_{ij}(x)=r_{ji}(-x)$ for all $i\not=j$. Then $\mathcal{G}_{\rm dist}$ is an MPG and a potential function $\Phi$ is given by (3.4), with $F$ and $G$ defined as
$F(t,x,a)\coloneqq\sum_{i=1}^{N}c_{i}(t,x_{i},a_{i})+\frac{1}{2}\sum_{i,j\in I_{N},i\not=j}h_{ij}(t,x_{i}-x_{j}),\quad G(x)\coloneqq\sum_{i=1}^{N}\ell_{i}(x_{i})+\frac{1}{2}\sum_{i,j\in I_{N},i\not=j}r_{ij}(x_{i}-x_{j}).$
Note that in lieu of Remark 3.1, here $U_{f_{i}}(t,x_{-i},a_{-i})\coloneqq-\sum_{j\in I_{N}\setminus\{i\}}c_{j}(t,x_{j},a_{j})-\frac{1}{2}\sum_{k,\ell\in I_{N}\setminus\{i\},k\not=\ell}h_{k\ell}(t,x_{k}-x_{\ell})$ and $U_{g_{i}}(x_{-i})\coloneqq-\sum_{j\in I_{N}\setminus\{i\}}\ell_{j}(x_{j})-\frac{1}{2}\sum_{k,\ell\in I_{N}\setminus\{i\},k\not=\ell}r_{k\ell}(x_{k}-x_{\ell})$ in (3.3), both of which are independent of player $i$’s state and action.
This crowd motion game has been studied in [4] for pedestrian motion modelling, in [10] for animal flocking, and [9] for synchronization in neuroscience. We emphasize that the form of game here generalizes those in [4, 10, 9]: condition (3.5) removes the homogeneous and the symmetric interaction conditions imposed in their analysis. In particular, it allows each player to have different individual preferences specified by $c_{i}$ and $\ell_{i}$, as well as to be differently influenced by other players’ positions as indicated by the kernels $h_{ij}$ and $r_{ij}$.
Now, Theorem 3.1 can be extended with explicit constructions of $F$ and $G$ and hence explicit expression of the dynamic potential function $\Phi$, when the running and terminal costs are twice differentiable.
Theorem 3.2.
Take the game $\mathcal{G}_{\rm dist}$ in Definition 3.1. Suppose that for all $i\in I_{N}$ and $t\in[0,T]$, ${\mathbb{R}}^{n_{x}}\times{\mathbb{R}}^{n_{a}}\ni(x,a)\mapsto(f_{i}(t,x,a),g_{i}(x))\in{\mathbb{R}}^{2}$ is twice continuously differentiable, for all $i,j\in I_{N}$,
$\begin{pmatrix}\partial_{x_{i}x_{j}}f_{i}&\partial_{a_{i}x_{j}}f_{i}\\ \partial_{x_{i}a_{j}}f_{i}&\partial_{a_{i}a_{j}}f_{i}\end{pmatrix}=\begin{pmatrix}\partial_{x_{i}x_{j}}f_{j}&\partial_{a_{i}x_{j}}f_{j}\\ \partial_{x_{i}a_{j}}f_{j}&\partial_{a_{i}a_{j}}f_{j}\end{pmatrix},\quad\partial_{x_{i}x_{j}}g_{i}=\partial_{x_{i}x_{j}}g_{j},$ (3.6)
and for all $i\in I_{N}$, $\sup_{(t,x,a)\in[0,T]\times{\mathbb{R}}^{n_{x}}\times{\mathbb{R}}^{n_{a}}}\frac{|(\partial_{(x_{i},a_{i})}f_{i})(t,x,a)|+|(\partial_{x_{i}}g_{i})(x)|}{1+|x|+|a|}<\infty$. Then $\mathcal{G}_{\rm dist}$ is an MPG and a potential function $\Phi$ is given by (3.4), with $F$ and $G$ defined as
$\begin{split}F(t,x,a)&\coloneqq\sum_{i=1}^{N}\int_{0}^{1}\left(\begin{pmatrix}\partial_{x_{i}}f_{i}\\ \partial_{a_{i}}f_{i}\end{pmatrix}\big{(}t,\hat{x}+r(x-\hat{x}),\hat{a}+r(a-\hat{a})\big{)}\right)^{\top}\begin{pmatrix}x_{i}-\hat{x}_{i}\\ a_{i}-\hat{a}_{i}\end{pmatrix}{\mathrm{d}}r,\\ G(x)&\coloneqq\sum_{i=1}^{N}\int_{0}^{1}(\partial_{x_{i}}g_{i})(\hat{x}+r(x-\hat{x}))^{\top}(x_{i}-\hat{x}_{i}){\mathrm{d}}r\end{split}$ (3.7)
for some fixed $(\hat{x},\hat{a})\in{\mathbb{R}}^{n_{x}}\times{\mathbb{R}}^{n_{a}}$.
Theorem 3.2 is reminiscent of the differential characterizations of static potential games [37, 45]. However, as highlighted in Remark 3.1, static potential games and the dynamic ones are distinct. For instance, replacing (3.6) by the condition $(\partial_{a_{i}a_{j}}f_{i})(t,x,\cdot)=(\partial_{a_{i}a_{j}}f_{j})(t,x,\cdot)$ for all $(t,x)$ does not guarantee that $\mathcal{G}_{\rm dist}$ is an MPG.
In addition, Theorem 3.2 can be applied to identify a special class of dynamic games as an MFG with a quadratic form of potential function, if the costs of the game are quadratic and depend both on other players through their average behavior.
Example 3.2 (MPG with quadratic form of potential functions).
Take the game $\mathcal{G}_{\rm dist}$ in Definition 3.1. Suppose that $n_{i}=n$ and $k_{i}=k$ for all $i\in I_{N}$, and there exists $\bar{Q},Q_{i}:[0,T]\rightarrow{\mathbb{S}}^{n}$, $\bar{R},R_{i}:[0,T]\rightarrow{\mathbb{S}}^{k}$, $\bar{G},G_{i}\in{\mathbb{S}}^{n}$, $i\in I_{N}$, $\gamma,\kappa:[0,T]\rightarrow{\mathbb{R}}$ and $\eta\in{\mathbb{R}}$ such that for all $i\in I_{N}$ and $(t,x,a)\in[0,T]\times{\mathbb{R}}^{Nn}\times{\mathbb{R}}^{Nk}$,
$\begin{split}f_{i}(t,x,a)&=x_{i}^{\top}Q_{i}(t)x_{i}+\left(x_{i}-\gamma(t)\overline{x}^{(N-1)}_{-i}\right)^{\top}\bar{Q}(t)\left(x_{i}-\gamma(t)\overline{x}^{(N-1)}_{-i}\right)\\ &\quad+a_{i}^{\top}R_{i}(t)a_{i}+\left(a_{i}-\kappa(t)\overline{a}^{(N-1)}_{-i}\right)^{\top}\bar{R}(t)\left(a_{i}-\kappa(t)\overline{a}^{(N-1)}_{-i}\right),\\ g_{i}(x)&=x_{i}^{\top}G_{i}x_{i}+\left(x_{i}-\eta\overline{x}^{(N-1)}_{-i}\right)^{\top}\bar{G}\left(x_{i}-\eta\overline{x}^{(N-1)}_{-i}\right),\end{split}$ (3.8)
where $x_{i}\in{\mathbb{R}}^{n}$ and $a_{i}\in{\mathbb{R}}^{k}$ are player $i$’s state and action, respectively, and $\overline{x}^{(N-1)}_{-i}\in{\mathbb{R}}^{n}$ and $\overline{a}^{(N-1)}_{-i}\in{\mathbb{R}}^{k}$ are the average state and action of other players defined by:
$\overline{x}^{(N-1)}_{-i}=\frac{1}{N-1}\sum_{j=1,j\not=i}^{N}x_{j},\quad\overline{a}^{(N-1)}_{-i}=\frac{1}{N-1}\sum_{j=1,j\not=i}^{N}a_{j}.$
Then $\mathcal{G}_{\rm dist}$ is an MPG and a potential function $\Phi$ is defined by (3.4) with
$F(t,x,a)\coloneqq x^{\top}\tilde{Q}(t)x+a^{\top}\tilde{R}(t)a,\quad G(x)\coloneqq x^{\top}\tilde{G}x,$ (3.9)
where
$\tilde{Q}$ $=\begin{pmatrix}Q_{1}+\bar{Q}&-\frac{\gamma\bar{Q}}{N-1}&\dots&-\frac{\gamma\bar{Q}}{N-1}\\ -\frac{\gamma\bar{Q}}{N-1}&Q_{2}+\bar{Q}&\ddots&\vdots\\ \vdots&\ddots&\ddots&-\frac{\gamma\bar{Q}}{N-1}\\ -\frac{\gamma\bar{Q}}{N-1}&\dots&-\frac{\gamma\bar{Q}}{N-1}&Q_{N}+\bar{Q}\end{pmatrix},\quad\tilde{R}=\begin{pmatrix}R_{1}+\bar{R}&-\frac{\kappa\bar{R}}{N-1}&\dots&-\frac{\kappa\bar{R}}{N-1}\\ -\frac{\kappa\bar{R}}{N-1}&R_{2}+\bar{R}&\ddots&\vdots\\ \vdots&\ddots&\ddots&-\frac{\gamma\bar{R}}{N-1}\\ -\frac{\kappa\bar{R}}{N-1}&\dots&-\frac{\kappa\bar{R}}{N-1}&R_{N}+\bar{R}\end{pmatrix},$
$\tilde{G}$ $=\begin{pmatrix}G_{1}+\bar{G}&-\frac{\eta\bar{G}}{N-1}&\dots&-\frac{\eta\bar{G}}{N-1}\\ -\frac{\eta\bar{G}}{N-1}&G_{2}+\bar{G}&\ddots&\vdots\\ \vdots&\ddots&\ddots&-\frac{\eta\bar{G}}{N-1}\\ -\frac{\eta\bar{G}}{N-1}&\dots&-\frac{\eta\bar{G}}{N-1}&G_{N}+\bar{G}\end{pmatrix}.$
We reiterate that this is a game with heterogeneity among players, who can have different action sets, different state dynamics (i.e., different $b_{i}$ and $\sigma_{i}$), and different dependence on their individual behavior (i.e., different $Q_{i},R_{i}$ and $G_{i}$). This is in contrast to the classical $N$-player mean field games in [7, 10], which requires that all players have homogeneous state dynamics and cost functions. In particular, even when $N\rightarrow\infty$, the matrices $\tilde{Q}$, $\tilde{R}$ and $\tilde{G}$ become diagnolized indicating independence of actions among players, who nevertheless remain heterogeneous.
4 Characterization of Differentiable Potential Game
In this section, we propose, instead of Theorem 2.2, to construct potential functions based on derivatives of value functions, assuming their regularities with respect to policies.
Let us start by a precise definition of the differentiability of a (scalar-valued) function with respect to unilateral deviations of policies. Recall that for each $i\in I_{N}$, $\pi_{i}$ is a subset of all measurable functions from $\mathcal{S}$ to $A_{i}$. We denote by $\operatorname*{span}{(\pi_{i})}$ the vector space of all linear combinations of policies in $\pi_{i}$:
$\operatorname*{span}{(\pi_{i})}=\left\{\sum_{\ell=1}^{m}\alpha_{\ell}{\phi}^{(\ell)}\bigg{|}m\in{\mathbb{N}},(\alpha_{\ell})_{\ell=1}^{m}\subset{\mathbb{R}},(\phi^{(\ell)})_{\ell=1}^{m}\subset\pi_{i}\right\}.$
Definition 4.1.
Let $\pi^{(N)}=\prod_{i\in I_{N}}\pi_{i}$ be a convex set and $f:\pi^{(N)}\rightarrow{\mathbb{R}}$. For each $i\in I_{N}$, we say $f$ has a linear derivative with respect to $\pi_{i}$, if there exists $\frac{\delta f}{\delta\phi_{i}}:\pi^{(N)}\times\operatorname*{span}(\pi_{i})\rightarrow{\mathbb{R}}$, such that for all $\phi=(\phi_{i},\phi_{-i})\in\pi^{(N)}$, $\frac{\delta f}{\delta\phi_{i}}(\phi;\cdot)$ is linear and
$\lim_{\varepsilon\downarrow 0}\frac{f\big{(}(\phi_{i}+\varepsilon(\phi^{\prime}_{i}-\phi_{i}),\phi_{-i})\big{)}-f(\phi)}{\varepsilon}=\frac{\delta f}{\delta\phi_{i}}(\phi;\phi^{\prime}_{i}-\phi_{i}),\quad\forall\phi^{\prime}_{i}\in\pi_{i}.$ (4.1)
For each $i,j\in I_{N}$, we say $f$ has second-order linear derivatives with respect to $\pi_{i}\times\pi_{j}$, if (i) for all $k\in\{i,j\}$, $f$ has a linear derivative $\frac{\delta f}{\delta\phi_{k}}$ with respect to $\pi_{k}$, and (ii) for all $(k,\ell)\in\{(i,j),(j,i)\}$, there exists $\frac{\delta^{2}f}{\delta\phi_{k}\delta\phi_{\ell}}:\pi^{(N)}\times\operatorname*{span}(\pi_{k})\times\operatorname*{span}(\pi_{\ell})\rightarrow{\mathbb{R}}$ such that for all $\phi\in\pi^{(N)}$, $\frac{\delta^{2}f}{\delta\phi_{k}\delta\phi_{\ell}}(\phi,\cdot,\cdot)$ is bilinear and for all $\phi^{\prime}_{k}\in\operatorname*{span}(\pi_{k})$, $\frac{\delta^{2}f}{\delta\phi_{k}\delta\phi_{\ell}}(\cdot;\phi^{\prime}_{k},\cdot)$ is a linear derivative of $\frac{\delta f}{\delta\phi_{k}}(\cdot;\phi^{\prime}_{k})$ with respect to $\pi_{\ell}$. We refer to $\frac{\delta^{2}f}{\delta\phi_{i}\delta\phi_{j}}$ and $\frac{\delta^{2}f}{\delta\phi_{j}\delta\phi_{i}}$ as second-order linear derivatives of $f$ with respect to $\pi_{i}\times\pi_{j}$.
Remark 4.1.
Definition 4.1 allows $(\pi_{i})_{i\in I_{N}}$ to be generic convex sets of measurable functions, which may not be vector spaces. This is important for games with stochastic/mixed policies, whose policy class $\pi_{i}$ consists of functions mapping the system state to a probability measure over the action set. In general, $f:\pi^{(N)}\rightarrow{\mathbb{R}}$ may have multiple linear derivatives since the condition (4.1) may not uniquely determine the values of $\frac{\delta f}{\delta\phi_{i}}(\phi;\cdot)$ on $\operatorname*{span}{(\pi_{i})}$, especially when $\pi_{i}$ is not a vector space. This non-uniqueness of linear derivatives will not affect our subsequent analysis, as our results hold for any choices of $\frac{\delta f}{\delta\phi_{i}}$ satisfying the properties outlined in Definition 4.1. For a specific game with sufficiently regular coefficients, there exist natural choices of linear derivatives for value functions, as discussed in Sections 4.1 and 4.2. The same remark also applies to the second-order derivatives of $f$.
We then present two lemmas regarding the linear derivative of $f:\pi^{(N)}\rightarrow{\mathbb{R}}$, which are crucial for Theorem 4.3. Lemma 4.1 shows that if $f$ has a linear derivative in $\pi_{i}$, then $f$ is differentiable along any line segment within $\pi_{i}$. Additionally, Lemma 4.2 shows that if $f$ is differentiable concerning all unilateral deviations, then $f$ is differentiable with respect to perturbations of all players’ policies, and the derivative can be computed via a chain rule. The proofs are given in Appendix B.2.
Lemma 4.1.
Suppose $\pi^{(N)}$ is convex, $i\in I_{N}$, and $f:\pi^{(N)}\rightarrow{\mathbb{R}}$ has a linear derivative $\frac{\delta f}{\delta\phi_{i}}$ with respect to $\pi_{i}$. Let $\phi\in\pi^{(N)}$, $\phi^{\prime}_{i}\in\pi_{i}$, and for each $\varepsilon\in[0,1]$, let $\phi^{\varepsilon}=(\phi_{i}+\varepsilon(\phi^{\prime}_{i}-\phi_{i}),\phi_{-i})$. Then the map $[0,1]\ni\varepsilon\mapsto f(\phi^{\varepsilon})\in{\mathbb{R}}$ is differentiable and $\frac{{\mathrm{d}}}{{\mathrm{d}}\varepsilon}f(\phi^{\varepsilon})=\frac{\delta f}{\delta\phi_{i}}(\phi^{\varepsilon};\phi^{\prime}_{i}-\phi_{i})$ for all $\varepsilon\in[0,1]$.
Lemma 4.2.
Suppose $\pi^{(N)}$ is convex and for all $i\in I_{N}$, $f:\pi^{(N)}\rightarrow{\mathbb{R}}$ has a linear derivative $\frac{\delta f}{\delta\phi_{i}}$ with respect to $\pi_{i}$ such that for all $z,\phi\in\pi^{(N)}$ and $\phi^{\prime}_{i}\in\pi_{i}$ , $[0,1]^{N}\ni{\varepsilon}\mapsto\frac{\delta f}{\delta\phi_{i}}(z+\varepsilon\cdot(\phi-z);\phi^{\prime}_{i})$ is continuous at 00, where $z+\varepsilon\cdot(\phi-z)\coloneqq(z_{i}+\varepsilon_{i}({\phi}_{i}-z_{i}))_{i\in I_{N}}$. Then for all $z,\phi\in\pi^{(N)}$, the map $[0,1]\ni r\mapsto f(z+r(\phi-z))\in{\mathbb{R}}$ is differentiable and $\frac{{\mathrm{d}}}{{\mathrm{d}}r}f(z+r(\phi-z))=\sum_{j=1}^{N}\frac{\delta f}{\delta\phi_{j}}(z+r(\phi-z);\phi_{j}-z_{j})$.
We now show that if the value functions of a dynamics game are sufficiently regular and have a symmetric Jacobian, then this game is a potential game, and its dynamic potential function can be constructed via the linear derivative of its value function.
Theorem 4.3.
Let $\mathcal{G}=(I_{N},\mathcal{S},(A_{i})_{i\in I_{N}},\pi^{(N)},(V_{i})_{i\in I_{N}})$ be a game whose set of policy profiles $\pi^{(N)}$ is convex. Suppose that for some $s_{0}\in\mathcal{S}$ and for all $i,j\in I_{N}$, the value function $V^{s_{0}}_{i}$ has second-order linear derivatives with respect to $\pi_{i}\times\pi_{j}$ such that for all $z=(z_{j})_{j\in I_{N}},\phi=(\phi_{j})_{j\in I_{N}}\in\pi^{(N)}$, $\phi^{\prime}_{i},\tilde{\phi}^{\prime}_{i}\in\pi_{i}$ and ${\phi}^{\prime\prime}_{j}\in\pi_{j}$,
  1. (1)
    (Boundedness.) $\sup_{r,\varepsilon\in[0,1]}\left|\frac{\delta^{2}V_{i}^{s_{0}}}{\delta\phi_{i}\delta\phi_{j}}\big{(}z+r(\phi^{\varepsilon}-z);{\phi}^{\prime}_{i},{\phi}^{\prime\prime}_{j}\big{)}\right|<\infty$, where $\phi^{\varepsilon}\coloneqq(\phi_{i}+\varepsilon(\tilde{\phi}^{\prime}_{i}-\phi_{i}),\phi_{-i})$;
  2. (2)
    (Continuity.) $[0,1]^{N}\ni{\varepsilon}\mapsto\frac{\delta^{2}V_{i}^{s_{0}}}{\delta\phi_{i}\phi_{j}}(z+\varepsilon\cdot(\phi-z);\phi^{\prime}_{i},\phi^{\prime\prime}_{j})$ is continuous at 00, where $z+\varepsilon\cdot(\phi-z)\coloneqq(z_{i}+\varepsilon_{i}({\phi}_{i}-z_{i}))_{i\in I_{N}}$;
  3. (3)
    (Symmetric Jacobian.) $\frac{\delta^{2}V_{i}^{s_{0}}}{\delta\phi_{i}\delta\phi_{j}}(\phi;\phi^{\prime}_{i},\phi^{\prime\prime}_{j})=\frac{\delta^{2}V_{j}^{s_{0}}}{\delta\phi_{j}\delta\phi_{i}}(\phi;\phi^{\prime\prime}_{j},\phi^{\prime}_{i}).$
Then $\mathcal{G}$ with initial state $s_{0}$ is a CLPG and for any $z\in\pi^{(N)}$, $\Phi^{s_{0}}:\pi^{(N)}\rightarrow{\mathbb{R}}$ defined by
$\Phi^{s_{0}}(\phi)=\int_{0}^{1}\sum_{j=1}^{N}\frac{\delta V_{j}^{s_{0}}}{\delta\phi_{j}}(z+r(\phi-z);\phi_{j}-z_{j}){\mathrm{d}}r$ (4.2)
is a potential function. If the above conditions hold for all $s_{0}\in\mathcal{S}$, then $\mathcal{G}$ is an MPG with a potential function $(s,\phi)\mapsto\Phi^{s}(\phi)$ defined as in (4.2).
Remark 4.2.
Theorem 4.3 generalizes the differential characterization for static potential games in [37] to dynamic games. Indeed, according to [37, Theorem 4.5], a static game is a static potential game if ${\partial^{2}_{a_{i}a_{j}}V_{i}}={\partial^{2}_{a_{j}a_{i}}V_{j}}$ for all $i,j\in I_{N}$, and for any $z\in\prod_{i=1}^{N}A_{i}$, a potential function is given by $\Phi(a)=\int_{0}^{1}\sum_{i=1}^{N}({\partial_{a_{i}}V_{i}})(z+r(a-z))(a_{i}-z_{i}){\mathrm{d}}r,$ with ${\partial_{a_{i}}V_{i}}$ and ${\partial^{2}_{a_{i}a_{j}}V_{i}}$ being the classical derivatives of $V_{i}$ when the action sets are either intervals [37, 45] or finite-dimensional spaces [29, 2]. In comparison, Theorem 4.3 extends the characterization by replacing the classical derivative by the linear derivative with respect to policies, given by $\frac{\delta V_{i}}{\delta a_{i}}(a;a^{\prime}_{i})\coloneqq({\partial_{a_{i}}V_{i}})(a)a^{\prime}_{i}$ and $\frac{\delta^{2}V_{i}}{\delta a_{i}\delta a_{j}}(a;a^{\prime}_{i},a^{\prime\prime}_{j})\coloneqq({\partial^{2}_{a_{i}a_{j}}V_{i}})(a)a^{\prime}_{i}a^{\prime\prime}_{j}$.
In fact, the main technical challenge in establishing an analogous differential characterization for the dynamic game is defining an appropriate notion of the derivative of the value function with respect to policies, when the set of admissible policies $\pi^{(N)}$ may not form a vector space and lacks natural norms to define differentiability. Consequently, one cannot straightforwardly replace the classical derivatives in [37] with Fréchet derivatives to characterize dynamic potential games. For instance, in continuous-time problems with continuous state spaces, determining the differentiability of value functions in Markov policies is a well-known challenge (see [41, 42] for some results on continuity).
Our approach is to propose the linear derivatives of value functions, defined along a line segment connecting two policies. The advantage of this linear derivative for dynamic games is to avoid introducing a topology on $\pi^{(N)}$ and accommodate any convex policy classes.
For continuous-time stochastic games with state dynamics governed by controlled diffusions, linear derivatives exist and can be easily computed under mild regularity assumptions on the policy classes and model coefficients. For such games, we will provide detailed analysis for differentiable policies and coefficients in Section 4.1 and for nondifferentiable policies and coefficients in Section 4.2.
4.1 Probabilistic Characterization of Continuous-time Potential Game
Consider the game $\mathcal{G}_{\rm prob}=(I_{N},\mathcal{S},(A_{i})_{i\in I_{N}},\pi^{(N)},(V_{i})_{i\in I_{N}})$, where $I_{N}=\{1,\ldots,N\}$, $\mathcal{S}=[0,T]\times{\mathbb{R}}^{n_{x}}$ with $T>0$ and $n_{x}\in{\mathbb{N}}$; $A_{i}\subset{\mathbb{R}}^{k_{i}}$ with $k_{i}\in{\mathbb{N}}$ is player $i$’s action set; $\pi^{(N)}=\prod_{i\in I_{N}}\pi_{i}$ is the set of admissible policy profiles, where player $i$’s policy class $\pi_{i}$ is defined by:
$\pi_{i}\coloneqq\left\{\phi:[0,T]\times{\mathbb{R}}^{n_{x}}\rightarrow A_{i}\,\middle|\,\begin{aligned} &\textnormal{$\phi$ is measurable and for all $t\in[0,T]$, $x\mapsto\phi(t,x)$ is}\\ &\textnormal{twice continuously differentiable and }\\ &\textnormal{$\|\phi(\cdot,0)\|_{0}+\|\partial_{x}\phi\|_{0}+\|\partial_{xx}\phi\|_{0}<\infty$}\end{aligned}\right\}.$ (4.3)
Let $n_{a}=\sum_{i\in I_{N}}k_{i}$ and $A^{(N)}=\prod_{i\in I_{N}}A_{i}\subset{\mathbb{R}}^{n_{a}}$ be the set of action profiles of all players. Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space supporting an $n_{w}$-dimensional Brownian motion $W$. For each $i\in I_{N}$, player $i$’s value function $V_{i}:[0,T]\times{\mathbb{R}}^{n_{x}}\times\pi^{(N)}\rightarrow{\mathbb{R}}$ is given by: for all $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$ and $\phi\in\pi^{(N)}$,
$V_{i}^{t,x}(\phi)\coloneqq{\mathbb{E}}\left[\int_{t}^{T}f_{i}(s,X^{t,x,\phi}_{s},\phi(s,X^{t,x,\phi}_{s})){\mathrm{d}}s+g_{i}(X^{t,x,\phi}_{T})\right],$ (4.4)
where $f_{i}:[0,T]\times{\mathbb{R}}^{n_{x}}\times{\mathbb{R}}^{n_{a}}\rightarrow{\mathbb{R}}$ and $g_{i}:{\mathbb{R}}^{n_{x}}\rightarrow{\mathbb{R}}$ are given cost functions, $X^{t,x,\phi}$ is the state process governed by the following dynamics:
${\mathrm{d}}X_{s}=B(t,X_{s},\phi(s,X_{s})){\mathrm{d}}s+\Sigma(s,X_{s},\phi(s,X_{s})){\mathrm{d}}W_{s},\quad s\in[t,T];\quad X_{t}=x,$ (4.5)
and $(B,\Sigma):[0,T]\times{\mathbb{R}}^{n_{x}}\times{\mathbb{R}}^{n_{a}}\rightarrow{\mathbb{R}}^{n_{x}}\times{\mathbb{R}}^{n_{x}\times n_{w}}$ are given coefficients. Note that in (4.4), we write $V_{i}^{t,x}(\phi)\coloneqq V_{i}(t,x,\phi)$ , as we will analyze the derivatives of value functions for each fixed initial condition $(t,x)$ (see Theorem 4.4).
Remark 4.3.
It has been shown in [42] that under suitable assumptions, smooth policies are nearly optimal, thus justifying the restriction of the optimization problem to the space of smooth policies. Consequently, one can characterize value function’s linear derivatives via suitable sensitivity processes derived from a chain rule. Moreover, to facilitate algorithm design, most existing works restrict admissible policies to some parameterized families of smooth functions (see [29] and references therein). Finally, this differentiability restriction in (4.3) can be removed when characterizing the linear derivatives via a PDE approach, as will be shown in Section 4.2.
Throughout this section, the following regularity assumptions are imposed on the action sets $(A_{i})_{i\in I_{N}}$ and coefficients $(B,\Sigma,(f_{i})_{i\in I_{N}},(g_{i})_{i\in I_{N}})$.
H. 1.
$A_{i}\subset{\mathbb{R}}^{k_{i}}$, $i\in I_{N}$, is nonempty and convex. The functions $(B,\Sigma):[0,T]\times{\mathbb{R}}^{n_{x}}\times{\mathbb{R}}^{n_{a}}\rightarrow{\mathbb{R}}^{n_{x}}\times{\mathbb{R}}^{n_{x}\times n_{w}}$, $f_{i}:[0,T]\times{\mathbb{R}}^{n_{x}}\times{\mathbb{R}}^{n_{a}}\rightarrow{\mathbb{R}}$ and $g_{i}:{\mathbb{R}}^{n_{x}}\rightarrow{\mathbb{R}}$, $i\in I_{N}$, are measurable and satisfy the following properties: for all $i\in I_{N}$ and $t\in[0,T]$,
  1. (1)
    $(x,a)\mapsto\big{(}B(t,x,a),\Sigma(t,x,a),f_{i}(t,x,a),g_{i}(x)\big{)}$ are twice continuously differentiable.
  2. (2)
    $(x,a)\mapsto(B(t,x,a),\Sigma(t,x,a))$ is of linear growth and their first and second derivatives are bounded (uniformly in $t$).
  3. (3)
    $(x,a)\mapsto(f_{i}(t,x,a),g_{i}(x))$ and their first and second derivatives are of polynomial growth (uniformly in $t$).
Under Condition (H.1), for all $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$ and $\phi\in\operatorname*{span}(\pi^{(N)})$, (4.5) admits a unique strong solution $X^{t,x,\phi}\in\mathcal{S}^{\infty}([t,T];{\mathbb{R}}^{n_{x}})$ and the value function $V^{t,x}_{i}(\phi)$ in (4.4) is well-defined (see e.g., [47, Theorem 3.3.1]). We then introduce several stochastic processes, which will be used to describe the conditions under which $\mathcal{G}_{\rm prob}$ is a potential game.
Fix $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$ and $\phi\in\pi^{(N)}$. For each $i\in I_{N}$ and $\phi^{\prime}_{i}\in\operatorname*{span}(\pi_{i})$, let $\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi;\phi^{\prime}_{i})\in\mathcal{S}^{\infty}([t,T];{\mathbb{R}}^{n_{x}})$ be the solution to the following equation:
$\begin{split}{\mathrm{d}}Y_{s}&=(\partial_{x}B^{\phi})(s,X^{t,x,\phi}_{s})Y_{s}{\mathrm{d}}s+\big{(}(\partial_{x}\Sigma^{\phi})(s,X^{t,x,\phi}_{s})Y_{s}\big{)}{\mathrm{d}}W_{s}\\ &\quad+(\partial_{a_{i}}B^{\phi})[\phi^{\prime}_{i}](s,X^{t,x,\phi}_{s}){\mathrm{d}}s+(\partial_{a_{i}}\Sigma^{\phi})[\phi^{\prime}_{i}](s,X^{t,x,\phi}_{s}){\mathrm{d}}W_{s}\quad\forall s\in[t,T];\quad Y_{t}=0,\end{split}$ (4.6)
where $B^{\phi}(t,x)\coloneqq B(t,x,\phi(t,x))$, $(\partial_{a_{i}}B^{\phi})[\phi^{\prime}_{i}](t,x)\coloneqq(\partial_{a_{i}}B)(t,x,\phi(t,x))\phi^{\prime}_{i}(t,x)$, and $\Sigma^{\phi}(t,x)$ and $(\partial_{a_{i}}\Sigma^{\phi})[\phi^{\prime}_{i}](t,x)$ are defined similarly. Here the differentiations are taken componentwise, i.e., $((\partial_{x}B^{\phi})(s,X^{t,x,\phi}_{s})Y_{s})_{\ell}=\sum_{j=1}^{n_{x}}\partial_{x_{j}}B^{\phi}_{\ell}(s,X^{t,x,\phi}_{s})(Y_{s})_{j}$ for all $\ell=1,\ldots,n_{x}$. Equation (4.6) is the sensitivity equation of (4.5) with respect to $\phi_{i}$ (see Lemma A.1).
In addition, for each $\phi^{\prime}_{i}\in\operatorname*{span}(\pi_{i})$ and $\phi^{\prime\prime}_{j}\in\operatorname*{span}(\pi_{j})$, let $\frac{\delta^{2}X^{t,x}}{\delta\phi_{i}\delta\phi_{j}}(\phi;\phi^{\prime}_{i},\phi^{\prime\prime}_{j})\in\mathcal{S}^{\infty}([t,T];{\mathbb{R}}^{n_{x}})$ be the solution to the following equation:
$\begin{split}{\mathrm{d}}Z_{s}&=(\partial_{x}B^{\phi})(s,X^{t,x,\phi}_{s})Z_{s}{\mathrm{d}}s+\big{(}(\partial_{x}\Sigma^{\phi})(s,X^{t,x,\phi}_{s})Z_{s}\big{)}{\mathrm{d}}W_{s}\\ &\quad+F_{B}\left(s,X^{t,x,\phi}_{s},\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi;\phi^{\prime}_{i}),\frac{\delta X^{t,x}}{\delta\phi_{j}}(\phi;\phi^{\prime\prime}_{j}),\phi^{\prime}_{i},\phi^{\prime\prime}_{j}\right){\mathrm{d}}s\\ &\quad+F_{\Sigma}\left(s,X^{t,x,\phi}_{s},\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi;\phi^{\prime}_{i}),\frac{\delta X^{t,x}}{\delta\phi_{j}}(\phi;\phi^{\prime\prime}_{j}),\phi^{\prime}_{i},\phi^{\prime\prime}_{j}\right){\mathrm{d}}W_{s}\quad\forall s\in[t,T];\quad Z_{t}=0,\end{split}$ (4.7)
where
$\begin{split}&F_{B}(\cdot,y_{1},y_{2},\phi^{\prime}_{i},\phi^{\prime\prime}_{j})\\ &=y_{2}^{\top}(\partial_{xx}B^{\phi})(\cdot)y_{1}+\phi^{\prime\prime}_{j}(\cdot)^{\top}(\partial_{a_{i}a_{j}}B)(\cdot,\phi(\cdot))\phi^{\prime}_{i}(\cdot)\\ &\quad+\left[\phi^{\prime\prime}_{j}(\cdot)^{\top}(\partial_{xa_{j}}B)(\cdot,\phi(\cdot))+\phi^{\prime\prime}_{j}(\cdot)^{\top}(\partial_{aa_{j}}B)(\cdot,\phi(\cdot))\partial_{x}\phi(\cdot)\right]y_{1}+(\partial_{a_{j}}B)(\cdot,\phi(\cdot))\partial_{x}\phi^{\prime\prime}_{j}(\cdot)y_{1}\\ &\quad+y_{2}^{\top}\left[(\partial_{a_{i}x}B)(\cdot,\phi(\cdot))\phi^{\prime}_{i}(\cdot)+\partial_{x}\phi(\cdot)^{\top}(\partial_{a_{i}a}B)(\cdot,\phi(\cdot))\phi^{\prime}_{i}(\cdot)\right]+(\partial_{a_{i}}B)(\cdot,\phi(\cdot)\partial_{x}\phi^{\prime}_{i}(\cdot)y_{2},\end{split}$ (4.8)
and $F_{\Sigma}(\cdot,y_{1},y_{2},\phi^{\prime}_{i},\phi^{\prime\prime}_{j})$ is defined similarly. Equation (4.7) is the sensitivity equation of (4.5) with respect to $\phi_{i}$ and $\phi_{j}$ (see Lemma A.2).
Finally, let $\alpha^{t,x,\phi}_{s}=\phi(s,X^{t,x,\phi}_{s})$ for all $s\in[t,T]$, and for each $i,j\in I_{N}$, define $\frac{\delta\alpha^{t,x}}{\delta\phi_{i}}(\phi;\cdot):\operatorname*{span}(\pi_{i})\rightarrow\mathcal{H}^{\infty}([t,T];{\mathbb{R}}^{n_{a}})$ by
$\frac{\delta\alpha^{t,x}}{\delta\phi_{i}}(\phi;\phi^{\prime}_{i})_{s}$ $=(\partial_{x}\phi)(s,X^{t,x,\phi}_{s})\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi;\phi^{\prime}_{i})_{s}+E_{i}\phi^{\prime}_{i}(s,X^{t,x,\phi}_{s}),$ (4.9)
and define $\frac{\delta^{2}\alpha^{t,x}}{\delta\phi_{i}\delta\phi_{j}}(\phi;\cdot,\cdot):\operatorname*{span}(\pi_{i})\times\operatorname*{span}(\pi_{j})\rightarrow\mathcal{H}^{\infty}([t,T];{\mathbb{R}}^{n_{a}})$ by
$\begin{split}&\frac{\delta^{2}\alpha^{t,x}}{\delta\phi_{i}\delta\phi_{j}}(\phi;\phi^{\prime}_{i},\phi^{\prime\prime}_{j})_{s}\\ &=\left(\frac{\delta X^{t,x}}{\delta\phi_{j}}(\phi;\phi^{\prime\prime}_{j})_{s}\right)^{\top}(\partial_{xx}\phi)(s,X^{t,x,\phi}_{s})\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi;\phi^{\prime}_{i})_{s}+(\partial_{x}\phi)(s,X^{t,x,\phi}_{s})\frac{\delta^{2}X^{t,x}}{\delta\phi_{i}\delta\phi_{j}}(\phi;\phi^{\prime}_{i},\phi^{\prime\prime}_{j})_{s}\\ &\quad+E_{j}(\partial_{x}\phi^{\prime\prime}_{j})(s,X^{t,x,\phi}_{s})\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi;\phi^{\prime}_{i})_{s}+E_{i}(\partial_{x}\phi^{\prime}_{i})(s,X^{t,x,\phi}_{s})\frac{\delta X^{t,x}}{\delta\phi_{j}}(\phi;\phi^{\prime\prime}_{j})_{s},\end{split}$ (4.10)
where $E_{\ell}\in{\mathbb{R}}^{n_{a}\times k_{\ell}}$, $\ell\in\{i,j\}$, is a block row matrix whose $\ell$-th row is a $k_{\ell}$-by-$k_{\ell}$ identity matrix (player $\ell$’s action) and other rows are zero (other players’ actions).
The following theorem characterizes the linear derivatives of value functions in (4.4) using the sensitivity processes of the states and controls, and further constructs potential functions for the game $\mathcal{G}_{\rm prob}$ defined by (4.3) and (4.4).
Theorem 4.4.
Assume (H.1). For each $i\in I_{N}$, given $\pi_{i}$ and $V_{i}$ in (4.3) and (4.4), respectively.
  1. (1)
    For each $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$ and $i,j\in I_{N}$, define $\frac{\delta V^{t,x}_{i}}{\delta\phi_{j}}:\pi^{(N)}\times\operatorname*{span}(\pi_{k})\rightarrow{\mathbb{R}}$ such that for all $\phi\in\pi^{(N)}$ and $\phi^{\prime}_{j}\in\operatorname*{span}(\pi_{j})$,
    $\begin{split}\frac{\delta V^{t,x}_{i}}{\delta\phi_{j}}(\phi;\phi^{\prime}_{j})&={\mathbb{E}}\bigg{[}\int_{t}^{T}\begin{pmatrix}\partial_{x}f_{i}&\partial_{a}f_{i}\end{pmatrix}(s,X^{t,x,\phi}_{s},\alpha^{t,x,\phi}_{s})\begin{pmatrix}\frac{\delta X^{t,x}}{\delta\phi_{j}}(\phi;\phi^{\prime}_{j})_{s}\\ \frac{\delta\alpha^{t,x}}{\delta\phi_{j}}(\phi;\phi^{\prime}_{j})_{s}\end{pmatrix}{\mathrm{d}}s\bigg{]}\\ &\quad+{\mathbb{E}}\bigg{[}(\partial_{x}g_{i})(X^{t,x,\phi}_{T})\frac{\delta X^{t,x}}{\delta\phi_{j}}(\phi;\phi^{\prime}_{j})_{T}\bigg{]}.\end{split}$ (4.11)
    Then $\frac{\delta V^{t,x}_{i}}{\delta\phi_{j}}$ is a linear derivative of $\phi\mapsto V^{t,x}_{i}(\phi)$ with respect to $\pi_{j}$.
  2. (2)
    For each $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$ and $i,k,\ell\in I_{N}$, define $\frac{\delta^{2}V^{t,x}_{i}}{\delta\phi_{k}\delta\phi_{\ell}}:\pi^{(N)}\times\operatorname*{span}(\pi_{k})\times\operatorname*{span}(\pi_{\ell})\rightarrow{\mathbb{R}}$ such that for all $\phi\in\pi^{(N)}$, $\phi^{\prime}_{k}\in\operatorname*{span}(\pi_{k})$ and $\phi^{\prime\prime}_{\ell}\in\operatorname*{span}(\pi_{\ell})$,
    $\frac{\delta^{2}V^{t,x}_{i}}{\delta\phi_{k}\delta\phi_{\ell}}(\phi;\phi^{\prime}_{k},\phi^{\prime\prime}_{\ell})$
    $\quad={\mathbb{E}}\left[\int_{t}^{T}\begin{pmatrix}\frac{\delta X^{t,x}}{\delta\phi_{\ell}}(\phi;\phi^{\prime\prime}_{\ell})_{s}\\ \frac{\delta\alpha^{t,x}}{\delta\phi_{\ell}}(\phi;\phi^{\prime\prime}_{\ell})_{s}\end{pmatrix}^{\top}\begin{pmatrix}\partial_{xx}f_{i}&\partial_{ax}f_{i}\\ \partial_{xa}f_{i}&\partial_{aa}f_{i}\end{pmatrix}(s,X^{t,x,\phi}_{s},\alpha^{t,x,\phi}_{s})\begin{pmatrix}\frac{\delta X^{t,x}}{\delta\phi_{k}}(\phi;\phi^{\prime}_{k})_{s}\\ \frac{\delta\alpha^{t,x}}{\delta\phi_{k}}(\phi;\phi^{\prime}_{k})_{s}\end{pmatrix}{\mathrm{d}}s\right]$
    $\qquad+{\mathbb{E}}\bigg{[}\int_{t}^{T}\begin{pmatrix}\partial_{x}f_{i}&\partial_{a}f_{i}\end{pmatrix}(s,X^{t,x,\phi}_{s},\alpha^{t,x,\phi}_{s})\begin{pmatrix}\frac{\delta^{2}X^{t,x}}{\delta\phi_{k}\phi_{\ell}}(\phi;\phi^{\prime}_{k},\phi^{\prime\prime}_{\ell})_{s}\\ \frac{\delta^{2}\alpha^{t,x}}{\delta\phi_{k}\phi_{\ell}}(\phi;\phi^{\prime}_{k},\phi^{\prime\prime}_{\ell})_{s}\end{pmatrix}{\mathrm{d}}s\bigg{]}$
    $\qquad+{\mathbb{E}}\left[\left(\frac{\delta X^{t,x}}{\delta\phi_{\ell}}(\phi;\phi^{\prime\prime}_{\ell})_{T}\right)^{\top}(\partial_{xx}g_{i})(X^{t,x,\phi}_{T})\frac{\delta X^{t,x}}{\delta\phi_{k}}(\phi;\phi^{\prime}_{k})_{T}\right]$
    $\qquad+{\mathbb{E}}\left[(\partial_{x}g_{i})(X^{t,x,\phi}_{T})\frac{\delta^{2}X^{t,x}}{\delta\phi_{k}\phi_{\ell}}(\phi;\phi^{\prime}_{k},\phi^{\prime\prime}_{\ell})_{T}\right].$ (4.12)
    Then $\frac{\delta^{2}V^{t,x}_{i}}{\delta\phi_{i}\delta\phi_{j}}$ and $\frac{\delta^{2}V^{t,x}_{i}}{\delta\phi_{j}\delta\phi_{i}}$ are second-order linear derivatives of $\phi\mapsto V^{t,x}_{i}(\phi)$ with respect to $\pi_{i}\times\pi_{j}$.
  3. (3)
    If there exists $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$ such that for all $\phi\in\pi^{(N)}$, $i,j\in I_{N}$, $\phi^{\prime}_{i}\in\pi_{i}$ and ${\phi}^{\prime\prime}_{j}\in\pi_{j}$,
    $\frac{\delta^{2}V^{t,x}_{i}}{\delta\phi_{i}\delta\phi_{j}}(\phi;\phi^{\prime}_{i},\phi^{\prime\prime}_{j})=\frac{\delta^{2}V^{t,x}_{j}}{\delta\phi_{j}\delta\phi_{i}}(\phi;\phi^{\prime\prime}_{j},\phi^{\prime}_{i})$ (4.13)
    with $\frac{\delta^{2}V^{t,x}_{i}}{\delta\phi_{i}\delta\phi_{j}}$ and $\frac{\delta^{2}V^{t,x}_{j}}{\delta\phi_{j}\delta\phi_{i}}$ defined in Item 2, then $\mathcal{G}_{\rm prob}$ with initial condition $(t,x)$ is a CLPG and a potential function is given by (4.2) with $(\frac{\delta V^{t,x}_{i}}{\delta\phi_{i}})_{i\in I_{N}}$ defined in Item 1.
    Moreover, if (4.13) holds for all $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$, then $\mathcal{G}_{\rm prob}$ is an MPG.
Theorem 4.4 generalises Theorem 3.1 by accommodating interconnected states among players. Indeed, for the distributed game $\mathcal{G}_{\rm dist}$ in Definition 3.1, player $i$’s state and control processes depend solely on her own policy $\phi_{i}$, resulting in zero derivatives with respect to player $j$’s policy $\phi_{j}$ for all $j\neq i$. In this distributed game, (2) can be simplified and (4.13) follows from condition (3.6).
To prove Theorem 4.4, one first shows that (4.6) and (4.7) characterize the derivatives of $X^{t,x,\phi}$ with respect to policies, using the pathwise differentiability of solutions to SDEs [25, Theorem 4, p. 105]. Then, by the regularity assumptions and the chain rule [25, Theorem 9, p. 97], one can show that (4.9) and (4.10) characterize the derivatives of $\alpha^{t,x,\phi}$ with respect to policies, and further establish (4.11) and (2). The detailed proof is given in Appendix A.1.
4.2 PDE Characterization of Continuous-time Potential Games
This section studies the continuous-time games in Section 4.1 via a PDE approach. For general games with Hölder continuous policies and nondegenerate diffusion coefficients, potential functions and the linear derivatives of value functions are characterized using a system of linear PDEs.
Consider a game $\mathcal{G}_{\rm analyt}=(I_{N},\mathcal{S},(A_{i})_{i\in I_{N}},\pi^{(N)},(V_{i})_{i\in I_{N}})$, with $I_{N}=\{1,\ldots,N\}$, $\mathcal{S}=[0,T]\times{\mathbb{R}}^{n_{x}}$, $A_{i}\subset{\mathbb{R}}^{k_{i}}$ and $n_{a}=\sum_{i\in I_{N}}k_{i}$ as in Section 4.1. The set $\pi^{(N)}=\prod_{i\in I_{N}}\pi_{i}$ consists of all policy profiles, where player $i$’s policy class $\pi_{i}$ is defined by (cf. (4.3)):
$\pi_{i}=\{\phi:[0,T]\times{\mathbb{R}}^{n_{x}}\rightarrow A_{i}\mid\phi\in C^{\gamma/2,\gamma}([0,T]\times{\mathbb{R}}^{n_{x}};{\mathbb{R}}^{k_{i}})\}\quad\textnormal{for some $\gamma\in(0,1]$.}$ (4.14)
The value functions $(V_{i})_{i\in I_{N}}$ are defined in a manner analogous to (4.4). However, we impose a different set of regularity conditions compared to (H.1) to facilitate the subsequent PDE analysis.
H. 2.
$A_{i}\subset{\mathbb{R}}^{k_{i}}$, $i\in I_{N}$, is nonempty and convex. There exists $\beta\in(0,1]$ and $\kappa>0$ such that
  1. (1)
    $(B,\Sigma)\in C^{{\beta}/{2},\beta,\beta}([0,T]\times{\mathbb{R}}^{n_{x}}\times{\mathbb{R}}^{n_{a}};{\mathbb{R}}^{n_{x}}\times{\mathbb{R}}^{n_{x}\times n_{x}})$ and for all $i\in I_{N}$, $f_{i}\in C^{{\beta}/{2},\beta,2+\beta}([0,T]\times{\mathbb{R}}^{n_{x}}\times{\mathbb{R}}^{n_{a}};{\mathbb{R}})$ and $g_{i}\in C^{2+\beta}({\mathbb{R}}^{n_{x}};{\mathbb{R}})$;
  2. (2)
    For all $t\in[0,T]$ and $(x,a)\in{\mathbb{R}}^{n_{x}}\times{\mathbb{R}}^{n_{a}}$, $\Sigma(t,x,a)$ is symmetric and satisfies $\xi^{\top}\Sigma(t,x,a)\xi\geq\kappa|\xi|^{2}$ for all $\xi\in{\mathbb{R}}^{n_{x}}$.
Remark 4.4.
Condition (H.2) and (4.14) allow for Hölder continuous state coefficients and policies, relaxing the twice differentiability condition required in Section 4.1 (cf. (H.1) and (4.3)). The Hölder continuity of policies, along with the uniform ellipticity of the diffusion coefficient, ensures the uniqueness in law of the state process and facilitates characterizing the value functions’ linear derivatives by solutions of associated PDEs in Hölder spaces. If the diffusion coefficient of (4.5) is uncontrolled, one can in fact consider measurable policies and establish similar characterizations by analyzing PDEs in suitable Sobolev spaces (see e.g., [23, 43, 42]).
Under (H.2), by [36, Theorem 2], for each $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$ and $\phi\in\pi^{(N)}$, the state dynamics (4.5) admits a unique weak solution $X^{t,x,\phi}$ on a probability space $(\Omega,\mathcal{F},\bar{\mathbb{P}})$ and the value functions
$V^{\phi}_{i}(t,x)={\mathbb{E}}^{\bar{\mathbb{P}}}\left[\int_{t}^{T}f_{i}(s,X^{t,x,\phi}_{s},\phi(s,X^{t,x,\phi}_{s})){\mathrm{d}}s+g_{i}(X^{t,x,\phi}_{T})\right],\quad i\in I_{N}$ (4.15)
are well-defined. In contrast to (4.4), in (4.15), we use the notation $V^{\phi}_{i}(t,x)\coloneqq V_{i}(t,x,\phi)$ to analyze the value functions for all initial conditions $(t,x)$ simultaneously. By standard regularity results of linear PDEs (see e.g., [27, Theorem 5.1, p. 320]) and Itô’s formula, $(t,x)\mapsto V^{\phi}_{i}(t,x)$ is the unique classical solution to the following PDE: for all $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$,
$\begin{split}&\partial_{t}V(t,x)+\mathcal{L}^{\phi}V(t,x)+f_{i}(t,x,\phi(t,x))=0;\quad V(T,x)=g_{i}(x),\end{split}$ (4.16)
where $\mathcal{L}^{\phi}$ is the generator of (4.5) defined by
$\mathcal{L}^{\phi}u(t,x)=\frac{1}{2}\operatorname{tr}\big{(}\Sigma\Sigma^{\top}(t,x,\phi(t,x))(\partial_{xx}u)(t,x)\big{)}+B(t,x,\phi(t,x))^{\top}(\partial_{x}u)(t,x).$
We now introduce several equations associated with (4.16), whose solutions characterize the linear derivatives of $(V_{i})_{i\in I_{N}}$. Fix $i\in I_{N}$. For each $k\in I_{N}$, consider $\frac{\delta V_{i}}{\delta\phi_{k}}:[0,T]\times{\mathbb{R}}^{n_{x}}\times\pi^{(N)}\times\operatorname*{span}(\pi_{k})\rightarrow{\mathbb{R}}$:
$(t,x,\phi,\phi^{\prime}_{k})\mapsto\frac{\delta V_{i}}{\delta\phi_{k}}(t,x,\phi,\phi^{\prime}_{k})\coloneqq\frac{\delta V^{\phi}_{i}}{\delta\phi_{k}}(t,x;\phi^{\prime}_{k}),$ (4.17)
where $(t,x)\mapsto\frac{\delta V^{\phi}_{i}}{\delta\phi_{k}}(t,x;\phi^{\prime}_{k})$ solves the following equation: for all $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$,
[…]
with the function $H_{i}:[0,T]\times{\mathbb{R}}^{n_{x}}\times{\mathbb{R}}^{n_{x}}\times{\mathbb{R}}^{n_{x}\times n_{x}}\times{\mathbb{R}}^{n_{a}}\rightarrow{\mathbb{R}}$ defined as follows:
$H_{i}(t,x,y,z,a)=\frac{1}{2}\operatorname{tr}\big{(}(\Sigma\Sigma^{\top})(t,x,a)z\big{)}+B(t,x,a)^{\top}y+f_{i}(t,x,a).$ (4.19)
Equation (4.18) describes the sensitivity of (4.16) with respect to $\phi_{k}$. It is derived by considering the value function $V^{\phi^{\varepsilon}}$ with $\phi^{\varepsilon}=(\phi_{k}+\varepsilon(\phi^{\prime}_{k}-\phi_{k}),\phi_{-k})$ for $\varepsilon\in[0,1)$, and then differentiating the corresponding PDE (4.16) with respect to $\varepsilon$.
In addition, for each $k,\ell\in I_{N}$, consider the map $\frac{\delta^{2}V^{\phi}_{i}}{\delta\phi_{k}\phi_{\ell}}:[0,T]\times{\mathbb{R}}^{n_{x}}\times\pi^{(N)}\times\operatorname*{span}(\pi_{k})\times\operatorname*{span}(\pi_{\ell})\rightarrow{\mathbb{R}}$:
$(t,x,\phi,\phi^{\prime}_{k},\phi^{\prime\prime}_{\ell})\mapsto\frac{\delta^{2}V_{i}}{\delta\phi_{k}\phi_{\ell}}(t,x,\phi,\phi^{\prime}_{k},\phi^{\prime\prime}_{\ell})\coloneqq\frac{\delta^{2}V^{\phi}_{i}}{\delta\phi_{k}\phi_{\ell}}(t,x;\phi^{\prime}_{k},\phi^{\prime\prime}_{\ell}),$ (4.20)
where $(t,x)\mapsto\frac{\delta^{2}V^{\phi}_{i}}{\delta\phi_{k}\phi_{\ell}}(t,x;\phi^{\prime}_{k},\phi^{\prime\prime}_{\ell})$ solves the following equation: for all $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$,
[…]
with the function $L:[0,T]\times{\mathbb{R}}^{n_{x}}\times{\mathbb{R}}^{n_{x}}\times{\mathbb{R}}^{n_{x}\times n_{x}}\times{\mathbb{R}}^{n_{a}}\rightarrow{\mathbb{R}}$ defined by
$L(t,x,y,z,a)=\frac{1}{2}\operatorname{tr}\big{(}(\Sigma\Sigma^{\top})(t,x,a)z\big{)}+B(t,x,a)^{\top}y.$ (4.22)
Equation (4.21) describes the sensitivity of (4.16) with respect to $\phi_{k}$ and $\phi_{\ell}$.
Now, via the Schauder’s estimate [27, Theorem 5.1, p. 320] for (4.16), (4.18) and (4.21) in suitable Hölder spaces one can show that the maps (4.17) and (4.20) are well-defined, and are the linear derivatives of value function $V_{i}$.
Theorem 4.5.
Assume (H.2). For each $i\in I_{N}$, let $\pi_{i}$ and $V_{i}$ be defined by (4.14) and (4.15), respectively.
  1. (1)
    For all $i,k\in I_{N}$, the map $\frac{\delta V_{i}}{\delta\phi_{k}}$ in (4.17) is well-defined, and for all $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$, $(\phi,\phi^{\prime}_{k})\mapsto\frac{\delta V^{\phi}_{i}}{\delta\phi_{k}}(t,x;\phi^{\prime}_{k})$ is a linear derivative of $\phi\mapsto V^{\phi}_{i}(t,x)$ with respect to $\pi_{k}$.
  2. (2)
    For all $i,k,\ell\in I_{N}$, the map $\frac{\delta^{2}V_{i}}{\delta\phi_{k}\delta\phi_{\ell}}$ in (4.20) is well-defined, and for all $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$, $(\phi,\phi^{\prime}_{k},\phi^{\prime\prime}_{\ell})\mapsto\frac{\delta^{2}V^{\phi}_{i}}{\delta\phi_{k}\delta\phi_{\ell}}(t,x;\phi^{\prime}_{k},\phi^{\prime\prime}_{\ell})$ and $(\phi,\phi^{\prime\prime}_{\ell},\phi^{\prime}_{k})\mapsto\frac{\delta^{2}V^{\phi}_{i}}{\delta\phi_{\ell}\delta\phi_{k}}(t,x;\phi^{\prime\prime}_{\ell},\phi^{\prime}_{k})$ are second-order linear derivatives of $\phi\mapsto V^{\phi}_{i}(t,x)$ with respect to $\pi_{k}\times\pi_{\ell}$.
  3. (3)
    If there exists $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$ such that for all $\phi\in\pi^{(N)}$, $i,j\in I_{N}$, $\phi^{\prime}_{i}\in\pi_{i}$ and ${\phi}^{\prime\prime}_{j}\in\pi_{j}$,
    $\frac{\delta^{2}V^{\phi}_{i}}{\delta\phi_{i}\delta\phi_{j}}(t,x;\phi^{\prime}_{i},\phi^{\prime\prime}_{j})=\frac{\delta^{2}V^{\phi}_{j}}{\delta\phi_{j}\delta\phi_{i}}(t,x;\phi^{\prime\prime}_{j},\phi^{\prime}_{i}),$ (4.23)
    with $\frac{\delta^{2}V^{\phi}_{i}}{\delta\phi_{i}\delta\phi_{j}}$ and $\frac{\delta^{2}V^{\phi}_{j}}{\delta\phi_{j}\delta\phi_{i}}$ defined in Item 2, then $\mathcal{G}_{\rm analyt}$ with initial condition $(t,x)$ is a CLPG and a potential function is given by (4.2) with $(\frac{\delta V_{i}}{\delta\phi_{i}})_{i\in I_{N}}$ defined in Item 1. Moreover, if (4.23) holds for all $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$, then $\mathcal{G}_{\rm analyt}$ is an MPG.
Compared with the probabilistic characterizations in Theorem 4.4, (4.18) and (4.21) do not need the spatial derivatives of policies, allowing for characterizing the linear derivatives of value functions with a more general class of nondifferentiable policies.
4.3 Continuous-time Linear Quadratic Potential Game
Theorems 4.4 and 4.5 characterize continuous-time MPGs differently depending on the regularity of model coefficients and players’ admissible policies. This section shows that these two characterizations coincide for a class of linear-quadratic (LQ) games. Moreover, by leveraging the LQ structure of the problem, simpler characterizations of MPGs can be obtained through a system of ODEs. In this section, the time variable of all coefficients is dropped when there is no risk of confusion.
Consider the game $\mathcal{G}_{\rm LQ}=(I_{N},\mathcal{S},(A_{i})_{i\in I_{N}},\pi^{(N)},(V_{i})_{i\in I_{N}})$, where $I_{N}=\{1,\ldots,N\}$, $\mathcal{S}=[0,T]\times{\mathbb{R}}^{n_{x}}$, $A_{i}={\mathbb{R}}^{k_{i}}$, and $\pi^{(N)}=\prod_{i\in I_{N}}\pi_{i}$ is the set of policy profiles, where player $i$’s policy class $\pi_{i}$ contains linear policies defined by
$\pi_{i}=\{\phi:[0,T]\times{\mathbb{R}}^{n_{x}}\rightarrow{\mathbb{R}}^{k_{i}}\mid\phi(t,x)=K_{i}(t)x,\;K_{i}\in C([0,T];{\mathbb{R}}^{k_{i}\times n_{x}})\}.$ (4.24)
With an abuse of notation, we identify a policy $\phi_{i}\in\pi_{i}$ with the feedback map $K_{i}$, and identify a policy profile $\phi\in\pi^{(N)}$ with feedback maps $K=(K_{i})_{i\in I_{N}}$. For each $i\in I_{N}$, define player $i$’s value function $V_{i}:[0,T]\times{\mathbb{R}}^{n_{x}}\times\pi^{(N)}\rightarrow{\mathbb{R}}$ by
$\begin{split}V^{\phi}_{i}(t,x)&=\frac{1}{2}{\mathbb{E}}\bigg{[}\int_{t}^{T}\left(\big{(}X^{t,x,\phi}_{s}\big{)}^{\top}Q_{i}(t)X^{t,x,\phi}_{s}+\big{(}K(s)X^{t,x,\phi}_{s}\big{)}^{\top}R_{i}(t)K(s)X^{t,x,\phi}_{s}\right){\mathrm{d}}s\\ &\qquad+(X^{t,x,\phi}_{T})^{\top}G_{i}X^{t,x,\phi}_{T}\bigg{]},\end{split}$ (4.25)
where $X^{t,x,\phi}$ satisfies the following dynamics:
${\mathrm{d}}X_{s}=\left(A(s)X_{s}+B(s)K(s)X_{s}\right){\mathrm{d}}s+\sigma{\mathrm{d}}W_{s},\quad s\in[t,T];\quad X_{t}=x,$ (4.26)
and $W$ is a given $n_{w}$-dimensional standard Brownian motion. We assume the following standard conditions for the coefficients of (4.25) and (4.26): $A\in C([0,T];{\mathbb{R}}^{n_{x}\times n_{x}})$, $B\in C([0,T];{\mathbb{R}}^{n_{x}\times n_{a}})$, $\sigma\in{\mathbb{R}}^{n_{x}\times n_{w}}$, and for all $i\in I_{N}$, $Q_{i}\in C([0,T];{\mathbb{S}}^{n_{x}})$, $R_{i}\in C([0,T];{\mathbb{S}}^{n_{a}})$ and $G_{i}\in{\mathbb{S}}^{n_{x}}$. Here for ease of exposition and simple characterization, we focus on LQ games with (possibly degenerate) uncontrolled additive noises. As will be clear from the analysis that similar characterizations can be established for LQ games with controlled multiplicative noises.
Probabilistic characterization in LQ games.
To characterize the LQ games, write $B$ and $R_{i}$ in the following block form:
$B=(B_{1},\ldots,B_{N}),\quad R_{i}=((R_{i})_{1},\ldots,(R_{i})_{N})=((R_{i})_{h\ell})_{h,\ell\in I_{N}}$ (4.27)
with $B_{\ell}\in C([0,T];{\mathbb{R}}^{n_{x}\times k_{\ell}})$, $(R_{i})_{\ell}\in C([0,T];{\mathbb{R}}^{n_{a}\times k_{\ell}})$ and $(R_{i})_{h\ell}\in C([0,T];{\mathbb{R}}^{k_{h}\times k_{\ell}})$ for all $h,\ell\in I_{N}$. Fix $i,h,\ell\in I_{N}$ and $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$. The probabilistic characterization (4.11) of $\frac{\delta V^{t,x}_{i}}{\delta\phi_{h}}(\phi;\phi^{\prime}_{h})$ simplifies into
$\begin{split}\frac{\delta V^{t,x}_{i}}{\delta\phi_{h}}(\phi;\phi^{\prime}_{h})&={\mathbb{E}}\bigg{[}\int_{t}^{T}\bigg{(}X_{s}^{\top}Q_{i}Y^{h}_{s}+(KX_{s})^{\top}\Big{(}R_{i}KY^{h}_{s}+(R_{i})_{h}K^{\prime}_{h}X_{s}\Big{)}\bigg{)}{\mathrm{d}}s+X_{T}^{\top}G_{i}Y^{h}_{T}\bigg{]},\end{split}$ (4.28)
where $(R_{i})_{h}$ is defined in (4.27), $X=X^{t,x,\phi}$ is the state process satisfying (4.26), $Y^{h}$ is the sensitivity process of $X$ with respect $K^{\prime}_{h}$ satisfying (cf. (4.6)):
$\begin{split}{\mathrm{d}}Y^{h}_{s}&=\big{(}(A+BK)Y^{h}_{s}+B_{h}K^{\prime}_{h}X_{s}\big{)}{\mathrm{d}}s\quad\forall s\in[t,T];\quad Y^{h}_{t}=0.\end{split}$ (4.29)
Moreover, the probabilistic characterization (2) of $\frac{\delta^{2}V^{t,x}_{i}}{\delta\phi_{h}\delta\phi_{\ell}}(\phi;\phi^{\prime}_{h},\phi^{\prime\prime}_{\ell})$ simplifies into
$\begin{split}\frac{\delta^{2}V^{t,x}_{i}}{\delta\phi_{h}\delta\phi_{\ell}}(\phi;\phi^{\prime}_{h},\phi^{\prime\prime}_{\ell})&={\mathbb{E}}\left[\int_{t}^{T}\left(Y^{\ell}_{s}Q_{i}Y^{h}_{s}+\big{(}KY^{\ell}_{s}+E_{\ell}K^{\prime\prime}_{\ell}X_{s}\big{)}^{\top}R_{i}\big{(}KY^{h}_{s}+E_{h}K^{\prime}_{h}X_{s}\big{)}\right){\mathrm{d}}s\right]\\ &\quad+{\mathbb{E}}\bigg{[}\int_{t}^{T}\left(X^{\top}_{s}Q_{i}Z_{s}+(KX_{s})^{\top}R_{i}\big{(}KZ_{s}+E_{\ell}K^{\prime\prime}_{\ell}Y^{h}_{s}+E_{h}K^{\prime}_{h}Y^{\ell}_{s}\big{)}\right){\mathrm{d}}s\bigg{]}\\ &\quad+{\mathbb{E}}\left[(Y^{\ell}_{T})^{\top}G_{i}Y^{h}_{T}+X^{\top}_{T}G_{i}Z_{T}\right],\end{split}$ (4.30)
where $E_{j}\in{\mathbb{R}}^{n_{a}\times k_{j}}$, $j\in\{h,\ell\}$, is a block row matrix defined as in (4.10), $Z$ is the sensitivity process of $X$ with respect $K^{\prime}_{h}$ and $K^{\prime\prime}_{\ell}$ satisfying (cf. (4.7)):
$\begin{split}{\mathrm{d}}Z_{s}&=\left((A+BK)Z_{s}+B_{\ell}K^{\prime\prime}_{\ell}Y^{h}_{s}+B_{h}K^{\prime}_{h}Y^{\ell}_{s}\right){\mathrm{d}}s\quad\forall s\in[t,T];\quad Z_{t}=0.\end{split}$ (4.31)
PDE characterization in LQ games.
For the PDE characterization, define for all $\phi\in\pi^{(N)}$, $\phi^{\prime}_{h}=K^{\prime}_{h}\in\pi_{h}$ and $\phi^{\prime\prime}_{\ell}=K^{\prime\prime}_{\ell}\in\pi_{\ell}$,
${{V}^{\phi}_{i}}(t,x)$ $\coloneqq\frac{1}{2}x^{\top}\Psi_{i}(t)x+\psi_{i}(t)\quad\forall(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}},$ (4.32a)
${\frac{\delta\overline{V}^{\phi}_{i}}{\delta\phi_{h}}}(t,x;\phi^{\prime}_{h})$ $\coloneqq\frac{1}{2}x^{\top}\Theta^{h}_{i}(t)x+\theta^{h}_{i}(t)\quad\forall(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}},$ (4.32b)
${\frac{\delta^{2}\overline{V}^{\phi}_{i}}{\delta\phi_{h}\delta\phi_{\ell}}}(t,x;\phi^{\prime}_{h},\phi^{\prime\prime}_{\ell})$ $\coloneqq\frac{1}{2}x^{\top}\Lambda^{h,\ell}_{i}(t)x+\lambda^{h,\ell}_{i}(t)\quad\forall(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}},$ (4.32c)
where $\Psi_{i},\Theta^{h}_{i},\Lambda^{h,\ell}_{i}\in C([0,T];{\mathbb{S}}^{n_{x}})$ satisfies
$\dot{\Psi}_{i}$ $+\left(A+BK\right)^{\top}\Psi_{i}+\Psi_{i}\left(A+BK\right)+Q_{i}+K^{\top}R_{i}K=0\quad\forall t\in[0,T];\quad\Psi_{i}(T)=G_{i},$ (4.33a)
$\dot{\Theta}^{h}_{i}$ $+\left(A+BK\right)^{\top}\Theta^{h}_{i}+\Theta^{h}_{i}\left(A+BK\right)+\Psi_{i}B_{h}K^{\prime}_{h}+(B_{h}K^{\prime}_{h})^{\top}\Psi_{i}$
$+K^{\top}(R_{i})_{h}K^{\prime}_{h}+(K^{\prime}_{h})^{\top}\big{(}(R_{i})_{h}\big{)}^{\top}K=0\quad\forall t\in[0,T];\quad\Theta^{h}_{i}(T)=0,$ (4.33b)
$\dot{\Lambda}^{h,\ell}_{i}$ $+\left(A+BK\right)^{\top}\Lambda^{h,\ell}_{i}+\Lambda^{h,\ell}_{i}\left(A+BK\right)$
$+\left(\Theta^{h}_{i}B_{\ell}K^{\prime\prime}_{\ell}+(B_{\ell}K^{\prime\prime}_{\ell})^{\top}\Theta^{h}_{i}\right)+\left(\Theta^{\ell}_{i}B_{h}K^{\prime}_{h}+(B_{h}K^{\prime}_{h})^{\top}\Theta^{\ell}_{i}\right)$
$+(K^{\prime\prime}_{\ell})^{\top}(R_{i})_{\ell h}K^{\prime}_{h}+(K^{\prime}_{h})^{\top}\big{(}(R_{i})_{\ell h}\big{)}^{\top}K^{\prime\prime}_{\ell}=0\quad\forall t\in[0,T];\quad\Lambda^{h,\ell}_{i}(T)=0,$ (4.33c)
and $\psi_{i},\theta^{h}_{i},\lambda^{h,\ell}_{i}\in C([0,T];{\mathbb{R}})$ satisfies
$&\dot{\psi}_{i}+\frac{1}{2}\textnormal{tr}\left(\sigma\sigma^{\top}\Psi_{i}\right)=0\quad\forall t\in[0,T];\quad\psi_{i}(T)=0,\\ &\dot{\theta}^{h}_{i}+\frac{1}{2}\textnormal{tr}\left(\sigma\sigma^{\top}\Theta^{h}_{i}\right)=0\quad\forall t\in[0,T];\quad\theta^{h}_{i}(T)=0,\\ &\dot{\lambda}^{h,\ell}_{i}+\frac{1}{2}\textnormal{tr}\left(\sigma\sigma^{\top}\Lambda^{h,\ell}_{i}\right)=0\quad\forall t\in[0,T];\quad\lambda^{h,\ell}_{i}(T)=0.\end{aligned}\right.$ (4.34)
In (4.33) and (4.34), we use a dot to represent the derivative with respect to time, and use $\textnormal{tr}(\cdot)$ to denote the trace of a matrix. Additionally, for the sake of notational simplicity, we omit the dependence of $\Theta^{h}_{i},\theta^{h}_{i}$ on $K^{\prime}_{h}$, and the dependence of $\Lambda^{h,\ell}_{i},\lambda^{h,\ell}_{i}$ on $K^{\prime}_{h}$ and $K^{\prime\prime}_{\ell}$. One can easily verify that (4.32a), (4.32b) and (4.32c) are the unique solutions to the PDEs (4.16), (4.18) and (4.21), respectively.
The following theorem proves the equivalence of Theorems 4.4 and 4.5 for LQ games. The proof follows from a direct application of Itô’s formula, whose detail is given in Appendix A.2.
Theorem 4.6.
For all $i,h,\ell\in I_{N}$, $\phi\in\pi^{(N)}$, $\phi^{\prime}_{h}\in\pi_{h}$, $\phi^{\prime\prime}_{\ell}\in\pi_{\ell}$ and $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$,
$\frac{\delta V^{t,x}_{i}}{\delta\phi_{h}}(\phi;\phi^{\prime}_{h})={\frac{\delta\overline{V}^{\phi}_{i}}{\delta\phi_{h}}}(t,x;\phi^{\prime}_{h}),\quad\frac{\delta^{2}V^{t,x}_{i}}{\delta\phi_{h}\delta\phi_{\ell}}(\phi;\phi^{\prime}_{h},\phi^{\prime\prime}_{\ell})={\frac{\delta^{2}\overline{V}^{\phi}_{i}}{\delta\phi_{h}\delta\phi_{\ell}}}(t,x;\phi^{\prime}_{h},\phi^{\prime\prime}_{\ell}).$ (4.35)
That is, Condition (4.13) in Theorem 4.4 is equivalent to Condition (4.23) in Theorem 4.5.
Consequently, if there exists $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$ such that for all $i,j\in I_{N}$ with $i\not=j$, $\phi\in\pi^{(N)}$, $\phi^{\prime}_{i}\in\pi_{i}$ and $\phi^{\prime\prime}_{j}\in\pi_{j}$,
$\frac{1}{2}x^{\top}\Lambda^{i,j}_{i}(t)x+\lambda^{i,j}_{i}(t)=\frac{1}{2}x^{\top}\Lambda^{j,i}_{j}(t)x+\lambda^{j,i}_{j}(t),$ (4.36)
then $\mathcal{G}_{\rm LQ}$ with initial condition $(t,x)$ is a CLPG. Moreover, $\mathcal{G}_{\rm LQ}$ is an MPG if and only if for all $i,j\in I_{N}$ with $i\not=j$, $\phi\in\pi^{(N)}$, $\phi^{\prime}_{i}\in\pi_{i}$ and $\phi^{\prime\prime}_{j}\in\pi_{j}$,
$\Lambda^{i,j}_{i}(t)=\Lambda^{j,i}_{j}(t),\quad\forall t\in[0,T].$ (4.37)
Theorem 4.6 characterizes LQ MPGs using a system of ODEs parametrized by all players’ policies. In particular, an LQ MPG with the policy class (4.24) can be directly characterized by the first-order derivatives $\Theta^{i}_{j}$ of value functions.
Corollary 4.7.
$\mathcal{G}_{\rm LQ}$ is an MPG if and only if for all $i,j\in I_{N}$, $(R_{i})_{ij}=(R_{j})_{ij}$, and for all $\phi\in\pi^{(N)}$ and $\phi^{\prime}_{i}\in\pi_{i}$, $(\Theta^{i}_{i}-\Theta^{i}_{j})B_{j}=0$, where $\Theta^{i}_{i}$ and $\Theta^{i}_{j}$ are defined by (4.33b) and $B_{j}$ is given in (4.27).
By Corollary 4.7, one can establish, under suitable controllability assumptions of the system, the equivalence of the LQ MPG game and the team Markov game (introduced in Example 2.1).
Proposition 4.8.
Suppose $B=(B_{1},\ldots,B_{N})$ satisfies for all $i\in I_{N}$ and $t\in[0,T]$, $B_{i}(t)$ has full row rank. Then $\mathcal{G}_{\rm LQ}$ is an MPG if and only if for all $i,j\in I_{N}$, $Q_{i}=Q_{j}$, $R_{i}=R_{j}$ and $G_{i}=G_{j}$.
Remark 4.5.
Proposition 4.8 holds for multi-dimensional LQ games with a general state coefficient $A$, a general noise coefficient $\sigma$, and symmetric cost coefficients $(Q_{i},R_{i},G_{i})_{i\in I_{N}}$. The proof relies on the full rank condition of $(B_{i})_{i\in I_{N}}$, which by the Hautus lemma ensures each player’s controllability of the system.
This controllability allows players to drive the system to any desired state regardless of the policies chosen by other players. It implies a strong coupling between the players’ actions and hence requires the homogeneity of cost functions for a potential game. It suggests that the strength of interactions among players through dynamics and policies is crucial in determining whether a dynamic game is a potential game. These are intriguing insights about the structure of dynamic potential games. Quantifying the intricate interplay between the potential structure and player interactions for general dynamic games will be an intriguing research topic.
5 Proofs of Main Results
5.1 Proofs of Theorems 2.2, 3.1 and 3.2
Proof of Theorem 2.2.
We only prove the characterization of CLPGs with a fixed initial state $s_{0}\in\mathcal{G}$. The characterization of MPGs holds by repeating the argument for all initial states.
It is clear that if for all $i\in I_{N}$, the value function $V^{s_{0}}_{i}$ admits the decomposition (2.2), then $\mathcal{G}$ with initial state $s_{0}$ is a CLPG with potential $\Phi^{s_{0}}$. Hence it remains to prove that (2.2) is a necessary condition for $\mathcal{G}$ being a CLPG with potential $\Phi^{s_{0}}$. Let $i\in I_{N}$, $\phi_{-i}\in\pi^{(N)}_{-i}$ and $\phi_{i},\phi_{i}^{\prime},\phi_{i}^{\prime\prime}\in\pi_{i}$ be arbitrary policies. Then by Definition 2.2,
$V^{s_{0}}_{i}((\phi_{i},\phi_{-i}))$ $=\Phi^{s_{0}}((\phi_{i},\phi_{-i}))-\Phi^{s_{0}}((\phi_{i}^{\prime},\phi_{-i}))+V^{s_{0}}_{i}((\phi_{i}^{\prime},\phi_{-i})),$
$V^{s_{0}}_{i}((\phi_{i},\phi_{-i}))$ $=\Phi^{s_{0}}((\phi_{i},\phi_{-i}))-\Phi^{s_{0}}((\phi_{i}^{\prime\prime},\phi_{-i}))+V^{s_{0}}_{i}((\phi_{i}^{\prime\prime},\phi_{-i})).$
This shows that for any pair of policies $(\phi^{\prime}_{i},\phi^{\prime\prime}_{i})\in\pi_{i}$,
$-\Phi^{s_{0}}((\phi_{i}^{\prime},\phi_{-i}))+V^{s_{0}}_{i}((\phi_{i}^{\prime},\phi_{-i}))=-\Phi^{s_{0}}((\phi_{i}^{\prime\prime},\phi_{-i}))+V^{s_{0}}_{i}((\phi_{i}^{\prime\prime},\phi_{-i})).$
Consequently, $U^{s_{0}}_{i}(\phi_{-i})\coloneqq-\Phi^{s_{0}}((\phi_{i}^{\prime},\phi_{-i}))+V^{s_{0}}_{i}((\phi_{i}^{\prime},\phi_{-i}))$ is well-defined, as it only depends on $\phi_{-i}$ but is independent of $\phi^{\prime}_{i}\in\pi_{i}$. ∎
Proof of Theorem 3.1.
Observe that for all $i\in I_{N}$ and $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$, by (3.2) and (3.3), $V^{t,x}_{i}(\phi)=\Phi^{t,x}(\phi)+U^{t,x}_{i}(\phi_{-i}),$ where
$U^{t,x}_{i}(\phi_{-i})$ $={\mathbb{E}}^{\mathbb{P}}\left[\int_{t}^{T}U_{f_{i}}(s,X^{t,x,\phi,-i}_{s},\alpha^{t,x,\phi,-i}_{s}){\mathrm{d}}s+U_{g_{i}}(X^{t,x,\phi,-i}_{T})\right],$
with $X^{t,x,\phi,-i}=(X^{t,x_{j},\phi_{j}})_{j\not=i}$ and $\alpha^{t,x,\phi,-i}=(\phi_{j}(\cdot,X^{t,x_{j},\phi_{j}}))_{j\not=i}$ being the state and control processes of other players. By (3.1) and the definition of distributed policies, the states $(X^{t,x,\phi,\ell})_{\ell\in I_{N}}$ are decoupled and hence the term $U^{t,x}_{i}(\phi_{-i})$ depends only on $\phi_{-i}$ and is independent of $\phi_{i}$. This along with Theorem 2.2 implies that $\Phi$ is a potential function for $\mathcal{G}_{\rm dist}$. ∎
Proof of Theorem 3.2.
Recall that by [13, Theorem 1.3.1], for a given continuously differentiable function $H:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n}$, there exists a function $\Phi:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ such that $H(z)=\nabla\Phi(z)$ for all $z\in{\mathbb{R}}^{n}$ if and only if the Jacobian matrix of $H$ is symmetric for all $z\in{\mathbb{R}}^{n}$. In this case, $\Phi:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ can be chosen as
$\Phi(z)\coloneqq\int_{0}^{1}H(z^{0}+t(z-z^{0}))^{\top}(z-z^{0}){\mathrm{d}}t,$
for any given $z^{0}\in{\mathbb{R}}^{n}$. The desired conclusion follows by applying this theorem with different choices of $H$.
First, recall that $n_{x}+n_{a}=\sum_{i\in I_{N}}(n_{i}+k_{i})$ and for each $t\in[0,T]$, define $H_{f,t}:{\mathbb{R}}^{n_{x}+n_{a}}\rightarrow{\mathbb{R}}^{n_{x}+n_{a}}$ such that $(H_{f,t}(x,a))_{i}\coloneqq(\partial_{(x_{i},a_{i})}f_{i})(t,x,a)$ for all $i\in I_{N}$. The symmetry condition (3.6) implies that the Jacobian matrix of $H_{f,t}$ is symmetric for all $(x,a)\in{\mathbb{R}}^{n_{x}+n_{a}}$. Hence applying [13, Theorem 1.3.1] with $H=H_{f,t}$ implies that $H_{f,t}=\nabla F(t,\cdot)$ with $\nabla F$ defined in (3.7). In particular, for all $i\in I_{N}$, $(\partial_{(x_{i},a_{i})}f_{i})(t,x,a)=(\partial_{(x_{i},a_{i})}F)(t,x,a)$ for all $(t,x,a)\in[0,T]\times{\mathbb{R}}^{n_{x}}\times{\mathbb{R}}^{n_{a}}$. Thus the function $U_{f_{i}}\coloneqq f_{i}-F$ satisfies $\partial_{(x_{i},a_{i})}U_{f_{i}}\equiv 0$, which along the the mean-value theorem implies that $U_{f_{i}}$ is independent of $(x_{i},a_{i})$. Hence the function $F$ in (3.7) satisfies the condition (3.3). The claim that $G$ in (3.7) satisfies the condition (3.3) follows from similar arguments. ∎
5.2 Proof of Theorem 4.3
Proof of Theorem 4.3.
We only prove the characterization of CLPGs, as the characterization of MPGs follows by repeating the argument for all initial states. Throughout this proof, we denote by $s$ the fixed initial state $s_{0}$ for simplicity.
For each $i\in I_{N}$ and $\phi^{\prime}_{i}\in\pi_{i}$, using Condition 2 and applying Lemma 4.2 to $\frac{\delta V_{i}^{s}}{\delta\phi_{i}}(\cdot;{\phi}^{\prime}_{i})$ yield that for all $\phi,z\in\pi^{(N)}$, $[0,1]\ni r\mapsto\frac{\delta V_{i}^{s}}{\delta\phi_{i}}(z+r(\phi-z);{\phi}^{\prime}_{i})\in{\mathbb{R}}$ is differentiable and
$\frac{{\mathrm{d}}}{{\mathrm{d}}r}\frac{\delta V_{i}^{s}}{\delta\phi_{i}}(z+r(\phi-z);{\phi}^{\prime}_{i})=\sum_{j=1}^{N}\frac{\delta^{2}V_{i}^{s}}{\delta\phi_{i}\delta\phi_{j}}(z+r(\phi-z);{\phi}^{\prime}_{i},\phi_{j}-z_{j}).$ (5.1)
Moreover, as the second-order derivatives are bilinear with respect to the last two arguments, one can assume without loss of generality that Conditions 1 and 3 and (5.1) hold for all ${\phi}^{\prime}_{i}\in\operatorname*{span}(\pi_{i})$ and ${\phi}^{\prime\prime}_{j}\in\operatorname*{span}(\pi_{j})$.
Note that $\Phi^{s}:\pi^{(N)}\rightarrow{\mathbb{R}}$ in (4.2) is well-defined, as $r\mapsto\frac{\delta V_{j}^{s}}{\delta\phi_{j}}(z+r(\phi-z);\phi_{j}-z_{j})$ is continuous on $[0,1]$ for all $j\in I_{N}$. We first prove that for all $\phi=(\phi_{i})_{i\in I_{N}}\in\pi^{(N)}$, $i\in I_{N}$ and $\phi^{\prime}_{i}\in\pi_{i}$,
$\frac{{\mathrm{d}}}{{\mathrm{d}}\varepsilon}\Phi^{s}(\phi^{\varepsilon})\bigg{|}_{\varepsilon=0}=\frac{\delta V^{s}}{\delta\phi_{i}}(\phi;\phi^{\prime}_{i}-\phi_{i})\quad\textnormal{with $\phi^{\varepsilon}\coloneqq(\phi_{i}+\varepsilon(\phi^{\prime}_{i}-\phi_{i}),\phi_{-i})$.}$
Observe that by (4.2), for all $\varepsilon\in(0,1]$,
$\Phi^{s}(\phi^{\varepsilon})-\Phi^{s}(\phi)$ $=\int_{0}^{1}\sum_{j=1}^{N}\frac{\delta V_{j}^{s}}{\delta\phi_{j}}(z+r(\phi^{\varepsilon}-z);\phi_{j}+\varepsilon\delta_{ji}(\phi^{\prime}_{i}-\phi_{i})-z_{j}){\mathrm{d}}r$
$\quad-\int_{0}^{1}\sum_{j=1}^{N}\frac{\delta V_{j}^{s}}{\delta\phi_{j}}(z+r(\phi-z);\phi_{j}-z_{j}){\mathrm{d}}r,$
where $\delta_{ji}=0$ if $j\not=i$ and $\delta_{ii}=1$. Then by $z+r(\phi^{\varepsilon}-z)\in\pi^{(N)}$, for all $\varepsilon\in(0,1]$,
$\begin{split}\frac{\Phi^{s}(\phi^{\varepsilon})-\Phi^{s}(\phi)}{\varepsilon}&=\frac{1}{\varepsilon}\int_{0}^{1}\sum_{j=1}^{N}\left(\frac{\delta V_{j}^{s}}{\delta\phi_{j}}(z+r(\phi^{\varepsilon}-z);\phi_{j}-z_{j})-\frac{\delta V_{j}^{s}}{\delta\phi_{j}}(z+r(\phi-z);\phi_{j}-z_{j})\right){\mathrm{d}}r\\ &\quad+\frac{1}{\varepsilon}\int_{0}^{1}\sum_{j=1}^{N}\varepsilon\delta_{ji}\frac{\delta V_{j}^{s}}{\delta\phi_{j}}(z+r(\phi^{\varepsilon}-z);\phi^{\prime}_{i}-\phi_{i}){\mathrm{d}}r\\ &=\int_{0}^{1}\sum_{j=1}^{N}\frac{1}{\varepsilon}\left(\frac{\delta V_{j}^{s}}{\delta\phi_{j}}(z+r(\phi^{\varepsilon}-z);\phi_{j}-z_{j})-\frac{\delta V_{j}^{s}}{\delta\phi_{j}}(z+r(\phi-z);\phi_{j}-z_{j})\right){\mathrm{d}}r\\ &\quad+\int_{0}^{1}\frac{\delta V_{i}^{s}}{\delta\phi_{i}}(z+r(\phi^{\varepsilon}-z);\phi^{\prime}_{i}-\phi_{i}){\mathrm{d}}r.\end{split}$ (5.2)
To send $\varepsilon\rightarrow 0$ in (5.2), note that for all $\varepsilon\in[0,1]$, $(z+r(\phi^{\varepsilon}-z))_{-i}=z_{-i}+r(\phi_{-i}-z_{-i})$ and
$\begin{split}(z+r(\phi^{\varepsilon}-z))_{i}&=z_{i}+r(\phi_{i}+\varepsilon(\phi^{\prime}_{i}-\phi_{i})-z_{i})\\ &=z_{i}+r(\phi_{i}-z_{i})+\varepsilon\big{(}(z_{i}+r(\phi^{\prime}_{i}-z_{i}))-(z_{i}+r(\phi_{i}-z_{i}))\big{)}\end{split}$ (5.3)
with $z_{i}+r(\phi_{i}-z_{i}),z_{i}+r(\phi^{\prime}_{i}-z_{i})\in\pi_{i}$. Thus for all $j\in I_{N}$, as $\phi_{j}-z_{j}\in\operatorname*{span}(\pi_{j})$, the twice differentiability of $V^{s}_{j}$ and Lemma 4.1 imply that $\varepsilon\mapsto\frac{\delta V_{j}^{s}}{\delta\phi_{j}}(z+r(\phi^{\varepsilon}-z);\phi_{j}-z_{j})$ is differentiable on $[0,1]$ and
$\frac{{\mathrm{d}}}{{\mathrm{d}}\varepsilon}\frac{\delta V_{j}^{s}}{\delta\phi_{j}}(z+r(\phi^{\varepsilon}-z);\phi_{j}-z_{j})$ $=\frac{\delta^{2}V_{j}^{s}}{\delta\phi_{j}\delta\phi_{i}}(z+r(\phi^{\varepsilon}-z);\phi_{j}-z_{j},r(\phi^{\prime}_{i}-\phi_{i}))$
$=\frac{\delta^{2}V_{j}^{s}}{\delta\phi_{j}\delta\phi_{i}}(z+r(\phi^{\varepsilon}-z);\phi_{j}-z_{j},\phi^{\prime}_{i}-\phi_{i})r,$
where the last identity used the linearity of $\frac{\delta^{2}V_{j}^{s}}{\delta\phi_{j}\delta\phi_{i}}$ in its last component. Hence, by the mean value theorem, for all $\ \varepsilon\in(0,1]$,
$\begin{split}&\left|\frac{1}{\varepsilon}\left(\frac{\delta V_{j}^{s}}{\delta\phi_{j}}(z+r(\phi^{\varepsilon}-z);\phi_{j}-z_{j})-\frac{\delta V_{j}^{s}}{\delta\phi_{j}}(z+r(\phi-z);\phi_{j}-z_{j})\right)\right|\\ &\quad\leq\sup_{r,\varepsilon\in[0,1]}\left|\frac{\delta^{2}V_{j}^{s}}{\delta\phi_{j}\delta\phi_{i}}(z+r(\phi^{\varepsilon}-z);\phi_{j}-z_{j},\phi^{\prime}_{i}-\phi_{i})r\right|\\ &\quad=\sup_{r,\varepsilon\in[0,1]}\left|\frac{\delta^{2}V_{i}^{s}}{\delta\phi_{i}\delta\phi_{j}}(z+r(\phi^{\varepsilon}-z);\phi^{\prime}_{i}-\phi_{i},\phi_{j}-z_{j})r\right|<\infty,\end{split}$ (5.4)
where the last inequality follows from Conditions 1 and 3. Similarly, as $\phi^{\prime}_{i}-\phi_{i}\in\operatorname*{span}(\pi_{i})$, by the twice differentiability of $V^{s}_{i}$ (see (5.3)), for all $r\in(0,1)$,
$\lim_{\varepsilon\downarrow 0}\frac{\delta V_{i}^{s}}{\delta\phi_{i}}(z+r(\phi^{\varepsilon}-z);\phi^{\prime}_{i}-\phi_{i})=\frac{\delta V_{i}^{s}}{\delta\phi_{i}}(z+r(\phi-z);\phi^{\prime}_{i}-\phi_{i}),$
and for all $r,\varepsilon\in[0,1]$, by the mean value theorem,
$\left|\frac{\delta V_{i}^{s}}{\delta\phi_{i}}(z+r(\phi^{\varepsilon}-z);\phi^{\prime}_{i}-\phi_{i})\right|$
$\quad\leq\left|\frac{\delta V_{i}^{s}}{\delta\phi_{i}}(z+r(\phi-z);\phi^{\prime}_{i}-\phi_{i})\right|+\left|\frac{\delta^{2}V_{i}^{s}}{\delta\phi_{i}\delta\phi_{i}}(z+r(\phi^{\varepsilon}-z);\phi^{\prime}_{i}-\phi_{i},\phi^{\prime}_{i}-\phi_{i})r\right|,$
which is uniformly bounded with respect to $(r,\varepsilon)\in[0,1]^{2}$ due to (5.1) and Condition 1. Hence, letting $\varepsilon\rightarrow 0$ in (5.2) and using Lebesgue’s dominated convergence theorem give
$\begin{split}\frac{{\mathrm{d}}}{{\mathrm{d}}\varepsilon}\Phi^{s}(\phi^{\varepsilon})\bigg{|}_{\varepsilon=0}&=\int_{0}^{1}\sum_{j=1}^{N}\frac{\delta^{2}V_{j}^{s}}{\delta\phi_{j}\delta\phi_{i}}(z+r(\phi-z);\phi_{j}-z_{j},\phi^{\prime}_{i}-\phi_{i})r{\mathrm{d}}r+\int_{0}^{1}\frac{\delta V_{i}^{s}}{\delta\phi_{i}}(z+r(\phi-z);\phi^{\prime}_{i}-\phi_{i}){\mathrm{d}}r\\ &=\int_{0}^{1}\sum_{j=1}^{N}\frac{\delta^{2}V_{i}^{s}}{\delta\phi_{i}\delta\phi_{j}}(z+r(\phi-z);\phi^{\prime}_{i}-\phi_{i},\phi_{j}-z_{j})r{\mathrm{d}}r+\int_{0}^{1}\frac{\delta V_{i}^{s}}{\delta\phi_{i}}(z+r(\phi-z);\phi^{\prime}_{i}-\phi_{i}){\mathrm{d}}r\\ &=\int_{0}^{1}r\frac{{\mathrm{d}}}{{\mathrm{d}}r}\left(\frac{\delta V_{i}^{s}}{\delta\phi_{i}}(z+r(\phi-z);\phi^{\prime}_{i}-\phi_{i})\right){\mathrm{d}}r+\int_{0}^{1}\frac{\delta V_{i}^{s}}{\delta\phi_{i}}(z+r(\phi-z);\phi^{\prime}_{i}-\phi_{i}){\mathrm{d}}r,\end{split}$
where the second to last identity used Condition 3 and the last identity used (5.1). The integration by parts formula then yields
$\begin{split}&\frac{{\mathrm{d}}}{{\mathrm{d}}\varepsilon}\Phi^{s}(\phi^{\varepsilon})\bigg{|}_{\varepsilon=0}=\frac{\delta V_{i}^{s}}{\delta\phi_{i}}(\phi;\phi^{\prime}_{i}-\phi_{i}).\end{split}$ (5.5)
Now let $i\in I_{N}$, $\phi^{\prime}_{i}\in\pi_{i}$ and $\phi\in\pi^{(N)}$. For each $\varepsilon\in[0,1]$, let $\phi^{\varepsilon}=(\phi_{i}+\varepsilon(\phi^{\prime}_{i}-\phi_{i}),\phi_{-i})\in\pi^{(N)}$. By the differentiability of $V_{i}^{s}$ and Lemma 4.1, $\frac{{\mathrm{d}}}{{\mathrm{d}}\varepsilon}V_{i}^{s}(\phi^{\varepsilon})=\frac{\delta V_{i}^{s}}{\delta\phi_{i}}(\phi^{\varepsilon};\phi^{\prime}_{i}-\phi_{i})$ for all $\varepsilon\in[0,1]$, and $\varepsilon\mapsto\frac{\delta V_{i}^{s}}{\delta\phi_{i}}(\phi^{\varepsilon};\phi^{\prime}_{i}-\phi_{i})$ is differentiable on $[0,1]$. This implies that $\varepsilon\mapsto V_{i}^{s}(\phi^{\varepsilon})$ is continuously differentiable on $[0,1]$. Then by Lemma 4.1 and (5.5), $[0,1]\ni\varepsilon\mapsto\Phi^{s}(\phi^{\varepsilon})\in{\mathbb{R}}$ is also continuously differentiable with $\frac{{\mathrm{d}}}{{\mathrm{d}}\varepsilon}\Phi^{s}(\phi^{\varepsilon})=\frac{{\mathrm{d}}}{{\mathrm{d}}\varepsilon}V_{i}^{s}(\phi^{\varepsilon})$ for all $\varepsilon\in[0,1]$. Hence by the fundamental theorem of calculus,
$V_{i}^{s}((\phi^{\prime}_{i},\phi_{-i}))-V_{i}^{s}((\phi_{i},\phi_{-i}))$ $=\int_{0}^{1}\frac{\delta V_{i}^{s}}{\delta\phi_{i}}(\phi^{\varepsilon};\phi^{\prime}_{i}-\phi_{i}){\mathrm{d}}\varepsilon=\int_{0}^{1}\frac{\delta\Phi^{s}}{\delta\phi_{i}}(\phi^{\varepsilon};\phi^{\prime}_{i}-\phi_{i}){\mathrm{d}}\varepsilon$
$=\Phi^{s}((\phi^{\prime}_{i},\phi_{-i}))-\Phi^{s}((\phi_{i},\phi_{-i})).$
This proves that $\Phi^{s}$ defined in (4.2) is a potential function of the game $\mathcal{G}$ with initial state $s$. ∎
5.3 Proof of Theorem 4.5
Proof of Theorem 4.5 Item 1.
Fix $i,k\in I_{N}$, $\phi\in\pi^{(N)}$ and $\phi^{\prime}_{k}\in\operatorname*{span}(\pi_{k})$. Let $\eta=\beta\gamma$. By (H.2), the Hölder continuity of $\phi,\phi^{\prime}_{k}$ and $V^{\phi}_{i}$, it is clear that all coefficients of (4.18) are in $C^{{\eta}/{2},\eta}([0,T]\times{\mathbb{R}}^{n_{x}};{\mathbb{R}})$. Thus by [27, Theorem 5.1, p. 320], (4.18) admits a unique classical solution in $C^{1+{\eta}/{2},2+\eta}([0,T]\times{\mathbb{R}}^{n_{x}};{\mathbb{R}})$. This shows that the map $\frac{\delta V_{i}}{\delta\phi_{k}}$ in (4.17) is well-defined. Moreover, as (4.18) is linear in terms of $\phi^{\prime}_{k}$, $\operatorname*{span}(\pi_{k})\ni\phi^{\prime}_{k}\mapsto\frac{\delta V^{\phi}_{i}}{\delta\phi_{k}}(t,x;\phi^{\prime}_{k})\in{\mathbb{R}}$ is linear.
Now fix $\phi^{\prime}_{k}\in\pi_{k}$. For each $\varepsilon\in[0,1]$, let $\phi^{\varepsilon}=(\phi_{k}+\varepsilon(\phi^{\prime}_{k}-\phi_{k}),\phi_{-k})$. We claim that
$\lim_{\varepsilon\downarrow 0}\bigg{\|}\frac{V^{\phi^{\varepsilon}}_{i}(\cdot)-V^{\phi}_{i}(\cdot)}{\varepsilon}-\frac{\delta V^{\phi}_{i}}{\delta\phi_{k}}(\cdot;\phi^{\prime}_{k}-\phi_{k})\bigg{\|}_{1+\beta\gamma/2,2+\beta\gamma}=0.$ (5.6)
Recall that $V^{\phi^{\varepsilon}}_{i}$, $\varepsilon\in(0,1]$, satisfies the following PDE: for all $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$,
$\begin{split}&\partial_{t}V^{\phi^{\varepsilon}}_{i}(t,x)+\mathcal{L}^{\phi^{\varepsilon}}V^{\phi^{\varepsilon}}_{i}(t,x)+f_{i}(t,x,\phi^{\varepsilon}(t,x))=0;\quad V^{\phi^{\varepsilon}}_{i}(T,x)=g_{i}(x).\end{split}$ (5.7)
Then by (4.18), $U^{\varepsilon}(\cdot)=\frac{V^{\phi^{\varepsilon}}_{i}(\cdot)-V^{\phi}_{i}(\cdot)}{\varepsilon}-\frac{\delta V^{\phi}_{i}}{\delta\phi_{k}}(\cdot;\phi^{\prime}_{k}-\phi_{k})$ satisfies for all $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$,
$\begin{split}&\partial_{t}U^{\varepsilon}(t,x)+\mathcal{L}^{\phi^{\varepsilon}}U^{\varepsilon}(t,x)+F^{\varepsilon}(t,x)=0;\quad U^{\varepsilon}(T,x)=0,\end{split}$ (5.8)
where
$\begin{split}F^{\varepsilon}(t,x)&=\frac{(\mathcal{L}^{\phi^{\varepsilon}}-\mathcal{L}^{\phi})V(t,x)+f_{i}(t,x,\phi^{\varepsilon}(t,x))-f_{i}(t,x,\phi(t,x))}{\varepsilon}\\ &\quad-(\partial_{a_{k}}H_{i})\big{(}t,x,(\partial_{x}V^{\phi}_{i})(t,x),(\partial_{xx}V^{\phi}_{i})(t,x),\phi(t,x)\big{)}^{\top}(\phi^{\prime}_{k}-\phi_{k})(t,x)\\ &\quad+(\mathcal{L}^{\phi^{\varepsilon}}-\mathcal{L}^{\phi})\frac{\delta V^{\phi}_{i}}{\delta\phi_{k}}(t,x;\phi^{\prime}_{k}-\phi_{k}).\end{split}$ (5.9)
By Taylor’s theorem, for all $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$,
$\frac{f_{i}(t,x,\phi^{\varepsilon}(t,x))-f_{i}(t,x,\phi(t,x))}{\varepsilon}-(\partial_{a_{k}}f_{i})\big{(}t,x,\phi(t,x)\big{)}^{\top}(\phi^{\prime}_{k}-\phi_{k})(t,x)$
$=\varepsilon\int_{0}^{1}(\phi^{\prime}_{k}-\phi_{k})(t,x)^{\top}(\partial_{a_{k}a_{k}}f_{i})(t,x,\phi^{r\varepsilon}(t,x))(\phi^{\prime}_{k}-\phi_{k})(t,x){\mathrm{d}}r,$
which along with the Hölder continuity of $\partial_{a_{k}a_{k}}f_{i}$, $\phi_{k}$ and $\phi^{\prime}_{k}$ shows that there exists $C\geq 0$ such that $\|\frac{f_{i}(\cdot,\phi^{\varepsilon}(\cdot))-f_{i}(\cdot,\phi(\cdot))}{\varepsilon}-(\partial_{a_{k}}f_{i})\big{(}\cdot,\phi(\cdot)\big{)}^{\top}(\phi^{\prime}_{k}-\phi_{k})(\cdot)\|_{\eta/2,\eta}\leq C\varepsilon$ for all $\varepsilon\in(0,1]$. Similarly, by using $V,\frac{\delta V^{\phi}_{i}}{\delta\phi_{k}}(\cdot;\phi^{\prime}_{k}-\phi_{k})\in C^{1+{\eta}/{2},2+\eta}([0,T]\times{\mathbb{R}}^{n_{x}};{\mathbb{R}})$, one can prove the $\|\cdot\|_{\eta/2,\eta}$-norm of the remaining terms in (5.9) converges to zero as $\varepsilon\rightarrow 0$. As $\phi^{\varepsilon}$ is a convex combination of $\phi$ and $(\phi^{\prime}_{k},\phi_{-k})$ (see (4.14)), $\sup_{\varepsilon\in[0,1]}\|\phi^{\varepsilon}\|_{\gamma/2,\gamma}<\infty$. Hence applying the Schauder a-priori estimate [27, Theorem 5.1, p. 320] to (5.8) implies that there exists $C\geq 0$ such that for all $\varepsilon\in(0,1]$, $\|U^{\varepsilon}\|_{1+\eta/2,2+\eta}\leq C\|F^{\varepsilon}\|_{\eta/2,\eta}\leq C\varepsilon$. This proves (5.6), which implies that $\frac{{\mathrm{d}}}{{\mathrm{d}}\varepsilon}{V^{\phi^{\varepsilon}}_{i}}(t,x)\big{|}_{\varepsilon=0}=\frac{\delta V^{\phi}_{i}}{\delta\phi_{k}}(t,x;\phi^{\prime}_{k}-\phi_{k})$. ∎
Proof of Theorem 4.5 Item 2.
Fix $i,k,\ell\in I_{N}$, $\phi\in\pi^{(N)}$, $\phi^{\prime}_{k}\in\operatorname*{span}(\pi_{k})$ and $\phi^{\prime\prime}_{\ell}\in\operatorname*{span}(\pi_{\ell})$. Let $\eta=\beta\gamma$. Recall that $V^{\phi}_{i}(\cdot)$, $\frac{\delta V^{\phi}_{i}}{\delta\phi_{k}}(\cdot;\phi^{\prime}_{k})$ and $\frac{\delta V^{\phi}_{i}}{\delta\phi_{\ell}}(\cdot;\phi^{\prime\prime}_{\ell})$ are in $C^{1+{\eta}/{2},2+\eta}([0,T]\times{\mathbb{R}}^{n_{x}};{\mathbb{R}})$, as shown in the proof of Theorem 4.5 Item 1. Hence the well-posedness of (4.21) in $C^{1+{\eta}/{2},2+\eta}([0,T]\times{\mathbb{R}}^{n_{x}};{\mathbb{R}})$ follows from a similar argument as that for (4.18), which subsequently implies that the map $\frac{\delta^{2}V_{i}}{\delta\phi_{k}\delta\phi_{\ell}}$ in (4.20) is well-defined. Moreover, by (4.18), for all $k\in I_{N}$, $\phi^{\prime}_{k}\mapsto\big{(}\partial_{x}\frac{\delta V^{\phi}_{i}}{\delta\phi_{k}}\big{)}(t,x;\phi^{\prime}_{k})$ and $\phi^{\prime}_{k}\mapsto\big{(}\partial_{xx}\frac{\delta V^{\phi}_{i}}{\delta\phi_{k}}\big{)}(t,x;\phi^{\prime}_{k})$ are linear. This shows that (4.21) is bilinear in terms of $\phi^{\prime}_{k}$ and $\phi^{\prime\prime}_{\ell}$, and hence the bilinearity of $(\phi^{\prime}_{k},\phi^{\prime\prime}_{\ell})\mapsto\frac{\delta^{2}V^{\phi}_{i}}{\delta\phi_{k}\phi_{\ell}}(t,x;\phi^{\prime}_{k},\phi^{\prime\prime}_{\ell})$.
Now fix $\phi^{\prime}_{k}\in\operatorname*{span}(\pi_{k})$ and $\phi^{\prime\prime}_{\ell}\in\pi_{\ell}$. For all $\varepsilon\in[0,1]$, let $\phi^{\varepsilon}=(\phi_{\ell}+\varepsilon(\phi^{\prime\prime}_{\ell}-\phi_{\ell}),\phi_{-\ell})$. We aim to prove
$\lim_{\varepsilon\downarrow 0}\bigg{\|}\frac{1}{\varepsilon}\bigg{(}\frac{\delta V^{\phi^{\varepsilon}}_{i}}{\delta\phi_{k}}(\cdot;\phi^{\prime}_{k})-\frac{\delta V^{\phi}_{i}}{\delta\phi_{k}}(\cdot;\phi^{\prime}_{k})\bigg{)}-\frac{\delta^{2}V^{\phi}_{i}}{\delta\phi_{k}\delta\phi_{\ell}}(\cdot;\phi^{\prime}_{k},\phi^{\prime\prime}_{\ell}-\phi_{\ell})\bigg{\|}_{0}=0.$ (5.10)
By (4.18) and (4.21), for all $\varepsilon\in(0,1]$,
$W^{\varepsilon}(\cdot)=\frac{1}{\varepsilon}\left(\frac{\delta V^{\phi^{\varepsilon}}_{i}}{\delta\phi_{k}}(\cdot;\phi^{\prime}_{k})-\frac{\delta V^{\phi}_{i}}{\delta\phi_{k}}(\cdot;\phi^{\prime}_{k})\right)-\frac{\delta^{2}V^{\phi}_{i}}{\delta\phi_{k}\delta\phi_{\ell}}(\cdot;\phi^{\prime}_{k},\phi^{\prime\prime}_{\ell}-\phi_{\ell})$
satisfies for all $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$,
$\partial_{t}W^{\varepsilon}(t,x)+\mathcal{L}^{\phi^{\varepsilon}}W^{\varepsilon}(t,x)+G^{\varepsilon}(t,x)=0;\ \ \ \ W^{\varepsilon}(T,x)=0,$ (5.11)
where $G^{\varepsilon}(t,x)=G^{\varepsilon}_{1}(t,x)+G^{\varepsilon}_{2}(t,x)+G^{\varepsilon}_{3}(t,x)$ satisfies for all $\bar{x}=(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$,
$G^{\varepsilon}_{1}(\bar{x})$ $=\frac{1}{\varepsilon}(\mathcal{L}^{\phi^{\varepsilon}}-\mathcal{L}^{\phi})\frac{\delta V^{\phi}_{i}}{\delta\phi_{k}}(\bar{x};\phi^{\prime}_{k})$ (5.12)
$\quad-(\partial_{a_{\ell}}L)\left(\bar{x},\left(\partial_{x}\frac{\delta V^{\phi}_{i}}{\delta\phi_{k}}\right)(\bar{x};\phi^{\prime}_{k}),\left(\partial_{xx}\frac{\delta V^{\phi}_{i}}{\delta\phi_{k}}\right)(\bar{x};\phi^{\prime}_{k}),\phi(\bar{x})\right)^{\top}(\phi^{\prime\prime}_{\ell}-\phi_{\ell})(\bar{x}),$
$G^{\varepsilon}_{2}(\bar{x})$ $=\frac{1}{\varepsilon}\bigg{[}(\partial_{a_{k}}H_{i})\big{(}\bar{x},(\partial_{x}V^{\phi^{\varepsilon}}_{i})(\bar{x}),(\partial_{xx}V^{\phi^{\varepsilon}}_{i})(\bar{x}),{\phi^{\varepsilon}}(\bar{x})\big{)}^{\top}\phi^{\prime}_{k}(\bar{x})$ (5.13)
$\quad-(\partial_{a_{k}}H_{i})\big{(}\bar{x},(\partial_{x}V^{\phi}_{i})(\bar{x}),(\partial_{xx}V^{\phi}_{i})(\bar{x}),\phi(\bar{x})\big{)}^{\top}\phi^{\prime}_{k}(\bar{x})\bigg{]}$
$\quad-\bigg{[}(\partial_{a_{k}}L)\left(\bar{x},\left(\partial_{x}\frac{\delta V^{\phi}_{i}}{\delta\phi_{\ell}}\right)(\bar{x};\phi^{\prime\prime}_{\ell}-\phi_{\ell}),\left(\partial_{xx}\frac{\delta V^{\phi}_{i}}{\delta\phi_{\ell}}\right)(\bar{x};\phi^{\prime\prime}_{\ell}-\phi_{\ell}),\phi(\bar{x})\right)^{\top}\phi^{\prime}_{k}(\bar{x})$
$\quad+(\phi^{\prime\prime}_{\ell}-\phi_{\ell})(\bar{x})^{\top}(\partial_{a_{k}a_{\ell}}H_{i})\big{(}t,x,(\partial_{x}V^{\phi}_{i})(\bar{x}),(\partial_{xx}V^{\phi}_{i})(\bar{x}),\phi(\bar{x})\big{)}\phi^{\prime}_{k}(\bar{x})\bigg{]},$
$G^{\varepsilon}_{3}(\bar{x})$ $=(\mathcal{L}^{\phi^{\varepsilon}}-\mathcal{L}^{\phi})\frac{\delta^{2}V^{\phi}_{i}}{\delta\phi_{k}\delta\phi_{\ell}}(\bar{x};\phi^{\prime}_{k},\phi^{\prime\prime}_{\ell}-\phi_{\ell}).$ (5.14)
We claim that $\lim_{\varepsilon\downarrow 0}\|G^{\varepsilon}\|_{0}=0$. The convergences of $G^{\varepsilon}_{1}$ and $G^{\varepsilon}_{3}$ follow directly from the Hölder continuity of $B$ and $\Sigma$. For $G^{\varepsilon}_{2}$, using $H_{i}(t,x,y,z,a)-H_{i}(t,x,y^{\prime},z^{\prime},a)=L(t,x,y,z,a)-L(t,x,y^{\prime},z^{\prime},a)$ (see (4.19) and (4.22)),
$\|G^{\varepsilon}_{2}\|_{0}$ $\leq\bigg{\|}\frac{1}{\varepsilon}\bigg{[}(\partial_{a_{k}}L)\big{(}\cdot,\partial_{x}V^{\phi^{\varepsilon}}_{i},\partial_{xx}V^{\phi^{\varepsilon}}_{i},{\phi^{\varepsilon}}\big{)}^{\top}\phi^{\prime}_{k}-(\partial_{a_{k}}L)\big{(}\cdot,\partial_{x}V^{\phi}_{i},\partial_{xx}V^{\phi}_{i},{\phi^{\varepsilon}}\big{)}^{\top}\phi^{\prime}_{k}\bigg{]}$
$\quad-(\partial_{a_{k}}L)\left(\cdot,\partial_{x}\frac{\delta V^{\phi}_{i}}{\delta\phi_{\ell}},\partial_{xx}\frac{\delta V^{\phi}_{i}}{\delta\phi_{\ell}},\phi^{\varepsilon}\right)^{\top}\phi^{\prime}_{k}\bigg{\|}_{0}$
$\quad+\bigg{\|}\frac{1}{\varepsilon}\left((\partial_{a_{k}}H_{i})\big{(}\cdot,\partial_{x}V^{\phi}_{i},\partial_{xx}V^{\phi}_{i},{\phi^{\varepsilon}}\big{)}-(\partial_{a_{k}}H_{i})\big{(}\cdot,(\partial_{x}V^{\phi}_{i}),(\partial_{xx}V^{\phi}_{i}),\phi\big{)}\right)^{\top}\phi^{\prime}_{k}$
$\quad-(\phi^{\prime\prime}_{\ell}-\phi_{\ell})^{\top}(\partial_{a_{k}a_{\ell}}H_{i})\big{(}\cdot,(\partial_{x}V^{\phi}_{i}),(\partial_{xx}V^{\phi}_{i}),\phi\big{)}\phi^{\prime}_{k}\bigg{\|}_{0}$
$\quad+\left\|(\partial_{a_{k}}L)\left(\cdot,\partial_{x}\frac{\delta V^{\phi}_{i}}{\delta\phi_{\ell}},\partial_{xx}\frac{\delta V^{\phi}_{i}}{\delta\phi_{\ell}},\phi^{\varepsilon}\right)-(\partial_{a_{k}}L)\left(\cdot,\partial_{x}\frac{\delta V^{\phi}_{i}}{\delta\phi_{\ell}},\partial_{xx}\frac{\delta V^{\phi}_{i}}{\delta\phi_{\ell}},\phi\right)\right\|_{0}\|\phi^{\prime}_{k}\|_{0}$
$\coloneqq G^{\varepsilon}_{2,1}+G^{\varepsilon}_{2,2}+G^{\varepsilon}_{2,3},$
where we wrote $\frac{\delta V^{\phi}_{i}}{\delta\phi_{\ell}}$ for $\frac{\delta V^{\phi}_{i}}{\delta\phi_{\ell}}(\cdot;\phi^{\prime\prime}_{\ell}-\phi_{\ell})$ to simplify the notation. By (5.6), $\lim_{\varepsilon\downarrow 0}\Big{\|}\frac{V^{\phi^{\varepsilon}}_{i}(\cdot)-V^{\phi}_{i}(\cdot)}{\varepsilon}-\frac{\delta V^{\phi}_{i}}{\delta\phi_{\ell}}(\cdot;\phi^{\prime\prime}_{\ell}-\phi_{\ell})\Big{\|}_{1+\eta/2,2+\eta}=0$, which along with the linearity of $(y,z)\mapsto L(t,x,y,z,a)$ and (H.2) shows $\lim_{\varepsilon\downarrow 0}\|G^{\varepsilon}_{2,1}\|_{0}=0$. For the term $G^{\varepsilon}_{2,2}$, for all $\bar{x}\in[0,T]\times{\mathbb{R}}^{n_{x}}$, writing $\partial_{x}V^{\phi}_{i}=(\partial_{x}V^{\phi}_{i})(\bar{x})$ and $\partial_{xx}V^{\phi}_{i}=(\partial_{xx}V^{\phi}_{i})(\bar{x})$ and applying the fundamental theorem of calculus yield
$\frac{1}{\varepsilon}\left((\partial_{a_{k}}H_{i})\big{(}\bar{x},\partial_{x}V^{\phi}_{i},\partial_{xx}V^{\phi}_{i},{\phi^{\varepsilon}}(\bar{x})\big{)}-(\partial_{a_{k}}H_{i})\big{(}\bar{x},\partial_{x}V^{\phi}_{i},\partial_{xx}V^{\phi}_{i},\phi(\bar{x})\big{)}\right)$
$\quad-(\partial_{a_{k}a_{\ell}}H_{i})\big{(}\bar{x},\partial_{x}V^{\phi}_{i},\partial_{xx}V^{\phi}_{i},\phi(\bar{x})\big{)}^{\top}(\phi^{\prime\prime}_{\ell}-\phi_{\ell})(\bar{x})$
$=\int_{0}^{1}\left((\partial_{a_{k}a_{\ell}}H_{i})\big{(}\bar{x},\partial_{x}V^{\phi}_{i},\partial_{xx}V^{\phi}_{i},\phi^{r\varepsilon}(\bar{x})\big{)}-(\partial_{a_{k}a_{\ell}}H_{i})\big{(}\bar{x},\partial_{x}V^{\phi}_{i},\partial_{xx}V^{\phi}_{i},\phi(\bar{x})\big{)}\right)^{\top}(\phi^{\prime\prime}_{\ell}-\phi_{\ell})(\bar{x}){\mathrm{d}}r$
$\leq C\varepsilon^{\beta}\|\phi^{\prime\prime}_{\ell}-\phi_{\ell}\|_{0}^{2},$
where the last inequality used the $\beta$-Hölder continuity of $a\mapsto(\partial_{a_{k}a_{\ell}}H_{i})\big{(}\bar{x},\partial_{x}V^{\phi}_{i},\partial_{xx}V^{\phi}_{i},a\big{)}$. This proves $\lim_{\varepsilon\downarrow 0}\|G^{\varepsilon}_{2,2}\|_{0}=0$. The Hölder continuity of coefficients gives $\lim_{\varepsilon\downarrow 0}\|G^{\varepsilon}_{2,3}\|_{0}=0$ and $\lim_{\varepsilon\downarrow 0}\|G^{\varepsilon}_{3}\|_{0}=0$. Applying the maximum principle [27, Theorem 2.5, p. 18] to (5.11) yields that there exists $C\geq 0$ such that $\|W^{\varepsilon}\|_{0}\leq C\|G^{\varepsilon}\|_{0}$ for all $\varepsilon\in(0,1]$. This along with the convergence of $G^{\varepsilon}$ implies $\lim_{\varepsilon\downarrow 0}\|W^{\varepsilon}\|_{0}=0$ and proves (5.10). Consequently, for all $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$, $\frac{{\mathrm{d}}}{{\mathrm{d}}\varepsilon}\frac{\delta V^{\phi^{\varepsilon}}_{i}}{\delta\phi_{k}}(t,x;\phi^{\prime}_{k})\big{|}_{\varepsilon=0}=\frac{\delta^{2}V^{\phi}_{i}}{\delta\phi_{k}\delta\phi_{\ell}}(t,x;\phi^{\prime}_{k},\phi^{\prime\prime}_{\ell}-\phi_{\ell})$. Interchanging the roles of $k$ and $\ell$ in the above analysis proves the desired statement. ∎
Proof of Theorem 4.5 Item 3.
By Theorem 4.5 Items 1 and 2, it remains to verify Theorem 4.3 Conditions 1 and 2 for fixed $i,j\in I_{N}$, $z,\phi\in\pi^{(N)}$, $\phi^{\prime}_{i},\tilde{\phi}^{\prime}_{i}\in\pi_{i}$ and ${\phi}^{\prime\prime}_{j}\in\pi_{j}$.
For Condition 1, for all $r,\varepsilon\in[0,1]$, let $\phi^{r,\varepsilon}=z+r((\phi_{i}+\varepsilon(\tilde{\phi}^{\prime}_{i}-\phi_{i}),\phi_{-i})-z)$. As $\sup_{r,\varepsilon\in[0,1]}\|\phi^{r,\varepsilon}\|_{\gamma/2,\gamma}<\infty$, by applying the Schauder a-priori estimate [27, Theorem 5.1, p. 320] to (4.16), (4.18) and (4.21), there exists $\eta\in(0,1]$ such that $\sup_{r,\varepsilon\in[0,1]}\Big{\|}\frac{\delta^{2}V^{\phi^{r,\varepsilon}}_{i}}{\delta\phi_{i}\delta\phi_{j}}(\cdot;\phi^{\prime}_{i},\phi^{\prime\prime}_{j})\Big{\|}_{\gamma/2,\gamma}<\infty$.
For Condition 2, for all $\varepsilon=(\varepsilon_{i})_{i=1}^{N}\in[0,1]^{N}$, let $\phi^{\varepsilon}=(z_{i}+\varepsilon_{i}({\phi}_{i}-z_{i}))_{i\in I_{N}}$ and observe that $U^{\varepsilon}=V^{\phi^{\varepsilon}}-V^{\phi^{0}}$ satisfies for all $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$,
[…]
As $\sup_{\varepsilon\in[0,1]^{N}}\|\phi^{\varepsilon}\|_{\gamma/2,\gamma}<\infty$, by (H.2), the Schauder a-priori estimate and the fundamental theorem of calculus, there exists $C\geq 0$ such that for all $\varepsilon\in[0,1]^{N}$, it holds with $\eta=\beta\gamma$ that
$\|U^{\varepsilon}\|_{1+\eta/2,2+\eta}$ $\leq C\|(\mathcal{L}^{\phi^{\varepsilon}}-\mathcal{L}^{\phi^{0}})V^{\phi^{0}}+f_{i}(\cdot,\phi^{\varepsilon}(\cdot))-f_{i}(\cdot,{\phi^{0}}(\cdot))\|_{\eta/2,\eta}$
$\leq C\left\|\int_{0}^{1}\sum_{j=1}^{N}(\partial_{a_{j}}H_{i})\big{(}\cdot,(\partial_{x}V^{\phi^{0}}_{i})(\cdot),(\partial_{xx}V^{\phi^{0}}_{i})(\cdot),\phi^{r\varepsilon}(\cdot)\big{)}^{\top}\varepsilon_{j}(\phi_{j}-z_{j})(\cdot){\mathrm{d}}r\right\|_{\eta/2,\eta},$
where $\phi^{r\varepsilon}=(z_{i}+r\varepsilon_{i}({\phi}_{i}-z_{i}))_{i\in I_{N}}$. This along with the Hölder continuity of $\partial_{a}H_{i}$ implies $\lim_{\varepsilon\downarrow 0}\|V^{\phi^{\varepsilon}}-V^{\phi^{0}}\|_{1+\beta\gamma/2,2+\beta\gamma}=0$. Applying similar arguments to (4.18) yields $\lim_{\varepsilon\downarrow 0}\frac{\delta V^{\phi^{\varepsilon}}_{i}}{\delta\phi_{i}}(\cdot;\phi^{\prime}_{i})=\frac{\delta V^{\phi^{0}}_{i}}{\delta\phi_{i}}(\cdot;\phi^{\prime}_{i})$ and $\lim_{\varepsilon\downarrow 0}\frac{\delta V^{\phi^{\varepsilon}}_{i}}{\delta\phi_{j}}(\cdot;\phi^{\prime\prime}_{j})=\frac{\delta V^{\phi^{0}}_{i}}{\delta\phi_{j}}(\cdot;\phi^{\prime\prime}_{j})$ with respect to the $\|\cdot\|_{1+\beta\gamma/2,2+\beta\gamma}$-norm. Finally, by (4.21) and the maximum principle, $\lim_{\varepsilon\downarrow 0}\big{\|}\frac{\delta^{2}V^{\phi^{\varepsilon}}_{i}}{\delta\phi_{i}\delta\phi_{j}}(\cdot;\phi^{\prime}_{i},\phi^{\prime\prime}_{j})-\frac{\delta^{2}V^{\phi^{0}}_{i}}{\delta\phi_{i}\delta\phi_{j}}(\cdot;\phi^{\prime}_{i},\phi^{\prime\prime}_{j})\big{\|}_{0}=0$. This verifies Condition 2. ∎
5.4 Proofs of Corollary 4.7 and Proposition 4.8
Proof of Corollary 4.7.
For each $\pi\in\pi^{(N)}$ with feedback maps $K$, define $F\coloneqq A+BK$. For each $i,j\in I_{N}$ with $i\not=j$, $\phi^{\prime}_{i}\in\pi_{i}$ and $\phi^{\prime\prime}_{j}\in\pi_{j}$, define $\Delta\Theta^{i}\coloneqq\Theta^{i}_{i}-\Theta^{i}_{j}$ and $\Delta\Theta^{j}\coloneqq\Theta^{j}_{i}-\Theta^{j}_{j}$. For any matrix $M\in{\mathbb{R}}^{n_{x}\times n_{x}}$, define $M^{S}\coloneqq M+M^{\top}$.
Fix $\pi\in\pi^{(N)}$. Observe from (4.33c) that for all $\phi^{\prime}_{i}\in\pi_{i}$ and $\phi^{\prime\prime}_{j}\in\pi_{j}$, $M\coloneqq\Lambda^{i,j}_{i}-\Lambda^{j,i}_{j}$ satisfies the ODE:
$\dot{M}$ $+F^{\top}M+MF+\left(\Delta\Theta^{i}B_{j}K^{\prime\prime}_{j}\right)^{S}+\left(\Delta\Theta^{j}B_{i}K^{\prime}_{i}\right)^{S}$
$+\left((K^{\prime\prime}_{j})^{\top}(R_{i}-R_{j})_{ji}K^{\prime}_{i}\right)^{S}=0,\quad t\in[0,T];\quad M(T)=0,$
where we used $\big{(}(R_{i})_{\ell h}\big{)}^{\top}=(R_{i})_{h\ell}$ for all $i,h,\ell\in I_{N}$, as $R_{i}$ takes values in ${\mathbb{S}}^{n_{a}}$. By Condition (4.37) and the continuity of coefficients, $\mathcal{G}_{\rm LQ}$ is an MPG if and only if for all $i,j\in I_{N}$ with $i\not=j$, $K_{i}\in C([0,T];{\mathbb{R}}^{k_{i}\times n_{x}})$ and $K_{j}\in C([0,T];{\mathbb{R}}^{k_{j}\times n_{x}})$,
$\left(\Delta\Theta^{i}B_{j}K^{\prime\prime}_{j}\right)^{S}+\left(\Delta\Theta^{j}B_{i}K^{\prime}_{i}\right)^{S}+\left((K^{\prime\prime}_{j})^{\top}(R_{i}-R_{j})_{ji}K^{\prime}_{i}\right)^{S}=0,$ (5.16)
Note that by (4.33c), $\Delta\Theta^{i}(T)=\Delta\Theta^{j}(T)=0$, and $\Delta\Theta^{i}(t)$ (resp. $\Delta\Theta^{j}(t)$) depends linearly on the integral of $K_{i}^{\prime}$ (resp. $K^{\prime\prime}_{j}$) for all $t\in[0,T]$. Hence for each $t\in[0,T]$, choosing a sequence of $K_{i}^{\prime}$ (resp. $K^{\prime\prime}_{j}$) that converges to 00 except at $t$ shows that (5.16) is equivalent to for all $K_{i}\in C([0,T];{\mathbb{R}}^{k_{i}\times n_{x}})$, $K_{j}\in C([0,T];{\mathbb{R}}^{k_{j}\times n_{x}})$, $M_{i}\in{\mathbb{R}}^{k_{i}\times n_{x}}$, and $M_{j}\in{\mathbb{R}}^{k_{i}\times n_{x}}$, $\left(\Delta\Theta^{i}B_{j}M_{j}\right)^{S}=0$, $\left(\Delta\Theta^{j}B_{i}M_{i}\right)^{S}=0$, and $\left(M_{j}^{\top}(R_{i}-R_{j})_{ji}M_{i}\right)^{S}=0$. Since $\textnormal{tr}(M^{S})=2\textnormal{tr}(M)$ for any $M\in{\mathbb{R}}^{n_{x}\times n_{x}}$, and $(M_{1},M_{2})\mapsto\textnormal{tr}(M^{\top}_{1}M_{2})$ is the Frobenius inner product, we see $\mathcal{G}_{\rm LQ}$ is an MPG if and only if for all $i,j\in I_{N}$ with $i\not=j$, $(R_{i})_{ji}=(R_{j})_{ji}$, and for all $K_{i}\in C([0,T];{\mathbb{R}}^{k_{i}\times n_{x}})$ and $K_{j}\in C([0,T];{\mathbb{R}}^{k_{j}\times n_{x}})$, $\Delta\Theta^{i}B_{j}=0$ and $\Delta\Theta^{j}B_{i}=0$. This finishes the proof. ∎
Proof of Proposition 4.8.
Since for all $t\in[0,T]$, $B_{i}(t)$ has full row rank and hence has a right inverse, by Corollary 4.7, $\mathcal{G}_{\rm LQ}$ is an MPG if and only if for all $i,j\in I_{N}$ with $i\not=j$, all $K\in C([0,T];{\mathbb{R}}^{n_{a}\times n_{x}})$, and all $K_{i}\in C([0,T];{\mathbb{R}}^{k_{i}\times n_{x}})$,
$(R_{i})_{ij}=(R_{j})_{ij}\quad\textnormal{and}\quad\Theta^{i}_{i}-\Theta^{i}_{j}=0.$ (5.17)
By (4.33b), $M\coloneqq\Theta^{i}_{i}-\Theta^{i}_{j}$ satisfies the ODE: for all $t\in[0,T]$,
$\dot{M}$ $+F^{\top}M+MF+\left((\Psi_{i}-\Psi_{j})B_{i}K^{\prime}_{i}\right)^{S}+\left(K^{\top}(R_{i}-R_{j})_{i}K^{\prime}_{i}\right)^{S}=0;\quad M(T)=0,$
where $A^{S}\coloneqq A+A^{\top}$ for any $A\in{\mathbb{R}}^{n_{x}\times n_{x}}$. Since $\Psi_{i}-\Psi_{j}$ is independent of $K^{\prime}_{i}$, it holds that (5.17) is equivalent to for all $i,j\in I_{N}$ with $i\not=j$, $(R_{i})_{ij}=(R_{j})_{ij}$ and for all $K\in C([0,T];{\mathbb{R}}^{n_{a}\times n_{x}})$, $(\Psi_{i}-\Psi_{j})B_{i}+K^{\top}(R_{i}-R_{j})_{i}=0$. Observe from (4.33a) that modifying $K$ at any time point will not change the value of $\Psi_{i}-\Psi_{j}$. This, along with a limiting argument and the full row rank of $B_{i}$, shows that $\mathcal{G}_{\rm LQ}$ is an MPG if and only if $R_{i}=R_{j}$ and $\Psi_{i}=\Psi_{j}$ for all $K\in C([0,T];{\mathbb{R}}^{n_{a}\times n_{x}})$, which is equivalent to that for all $i,j\in I_{N}$, $R_{i}=R_{j}$, $Q_{i}=Q_{j}$ and $G_{i}=G_{j}$ by (4.33a). This completes the proof. ∎
Appendix A Proofs of Theorems 4.4 and 4.6
A.1 Proof of Theorem 4.4
We start by presenting Lemmas A.1, A.2 and A.3, which characterize the derivatives of $X^{t,x,\phi}$ and $\alpha^{t,x,\phi}$ with respect to policies. The proofs are given in Appendix B.2.
The following notation will be used in the subsequent analysis: for each $t\in[0,T]$, $n\in{\mathbb{R}}$ and $(X^{n})_{n=0}^{\infty}\in\mathcal{S}^{\infty}([t,T];{\mathbb{R}}^{n})$, we write $\mathcal{S}^{\infty}\text{-}\lim_{n\rightarrow\infty}X^{n}=X^{0}$ if for all $q\geq 1$, ${\mathbb{E}}[\sup_{s\in[t,T]}|X^{n}_{s}-X^{0}_{s}|^{q}]<\infty$. Let $X^{\varepsilon}\in\mathcal{S}^{\infty}([t,T];{\mathbb{R}}^{n})$ for all $\varepsilon\in[0,1)$, we write $Y=\mathcal{S}^{\infty}\text{-}\partial_{\varepsilon}X^{\varepsilon}|_{\varepsilon=0}$ if $\mathcal{S}^{\infty}\text{-}\lim_{r\downarrow 0}\frac{1}{r}(X^{r}-X^{0})=Y$. Higher order derivatives of a process can be defined similarly. The continuity and differentiability for processes in $\mathcal{H}^{\infty}([t,T];{\mathbb{R}}^{n})$ are defined similarly.
Lemma A.1.
Assume (H.1). Let $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$, $\phi\in\pi^{(N)}$ and $i\in I_{N}$. For all $\phi^{\prime}_{i}\in\operatorname*{span}(\pi_{i})$, (4.6) admits a unique solution $\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi;\phi^{\prime}_{i})\in\mathcal{S}^{\infty}([t,T];{\mathbb{R}}^{n_{x}})$ satisfying the following properties:
  1. (1)
    The map $\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi;\cdot):\operatorname*{span}(\pi_{i})\rightarrow\mathcal{S}^{\infty}([t,T];{\mathbb{R}}^{n_{x}})$ is linear;
  2. (2)
    For all $\phi^{\prime}_{i}\in\pi_{i}$, $\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi;\phi^{\prime}_{i}-\phi_{i})=\mathcal{S}^{\infty}\text{-}\partial_{\varepsilon}X^{t,x,(\phi_{i}+\varepsilon(\phi_{i}^{\prime}-\phi_{i}),\phi_{-i})}\big{|}_{\varepsilon=0}$.
Lemma A.2.
Suppose (H.1) holds. Let $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$, $\phi\in\pi^{(N)}$ and $i,j\in I_{N}$. For all $\phi^{\prime}_{i}\in\operatorname*{span}(\pi_{i})$ and $\phi^{\prime\prime}_{j}\in\operatorname*{span}(\pi_{j})$, (4.7) admits a unique solution $\frac{\delta^{2}X^{t,x}}{\delta\phi_{i}\delta\phi_{j}}(\phi;\phi^{\prime}_{i},\phi^{\prime\prime}_{j})\in\mathcal{S}^{\infty}([t,T];{\mathbb{R}}^{n_{x}})$ satisfying the following properties:
  1. (1)
    The map $\frac{\delta^{2}X^{t,x}}{\delta\phi_{i}\delta\phi_{j}}(\phi;\cdot,\cdot):\operatorname*{span}(\pi_{i})\times\operatorname*{span}(\pi_{j})\rightarrow\mathcal{S}^{\infty}([t,T];{\mathbb{R}}^{n_{x}})$ is bilinear;
  2. (2)
    For all $\phi^{\prime}_{i}\in\pi_{i}$ and $\phi^{\prime\prime}_{j}\in\pi_{j}$, $\frac{\delta^{2}X^{t,x}}{\delta\phi_{i}\delta\phi_{j}}(\phi;\phi^{\prime}_{i},\phi^{\prime\prime}_{j})=\frac{\delta^{2}X^{t,x}}{\delta\phi_{j}\delta\phi_{i}}(\phi;\phi^{\prime\prime}_{j},\phi^{\prime}_{i})$;
  3. (3)
    For all $\phi^{\prime}_{i}\in\operatorname*{span}(\pi_{i})$ and $\phi^{\prime\prime}_{j}\in\pi_{j}$, $\frac{\delta^{2}X^{t,x}}{\delta\phi_{i}\delta\phi_{j}}(\phi;\phi^{\prime}_{i},\phi^{\prime\prime}_{j}-\phi_{j})=\mathcal{S}^{\infty}\text{-}\partial_{\varepsilon}\frac{\delta X^{t,x}}{\delta\phi_{i}}((\phi_{j}+\varepsilon(\phi^{\prime\prime}_{j}-\phi_{j}),\phi_{-j});\phi^{\prime}_{i})\big{|}_{\varepsilon=0}.$
Lemma A.3.
Suppose (H.1) holds. Let $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$ and $\phi\in\pi^{(N)}$. For each $s\in[t,T]$, let $\alpha^{t,x,\phi}_{s}=\phi(s,X^{t,x,\phi}_{s})$. Then for all $i,j\in I_{N}$,
  1. (1)
    $\frac{\delta\alpha^{t,x}}{\delta\phi_{i}}(\phi;\cdot):\operatorname*{span}(\pi_{i})\rightarrow\mathcal{H}^{\infty}([t,T];{\mathbb{R}}^{n_{a}})$ is linear, and $\frac{\delta\alpha^{t,x}}{\delta\phi_{i}}(\phi;\phi^{\prime}_{i}-\phi_{i})=\mathcal{H}^{\infty}\text{-}\partial_{\varepsilon}\alpha^{t,x,(\phi_{i}+\varepsilon(\phi_{i}^{\prime}-\phi_{i}),\phi_{-i})}\big{|}_{\varepsilon=0}$ for all $\phi^{\prime}_{i}\in\pi_{i}$;
  2. (2)
    $\frac{\delta^{2}\alpha^{t,x}}{\delta\phi_{i}\delta\phi_{j}}(\phi;\cdot,\cdot):\operatorname*{span}(\pi_{i})\times\operatorname*{span}(\pi_{j})\rightarrow\mathcal{H}^{\infty}([t,T];{\mathbb{R}}^{n_{a}})$ is bilinear, for all $\phi^{\prime}_{i}\in\pi_{i}$ and $\phi^{\prime\prime}_{j}\in\pi_{j}$, $\frac{\delta^{2}\alpha^{t,x}}{\delta\phi_{i}\delta\phi_{j}}(\phi;\phi^{\prime}_{i},\phi^{\prime\prime}_{j})=\frac{\delta^{2}\alpha^{t,x}}{\delta\phi_{j}\delta\phi_{i}}(\phi;\phi^{\prime\prime}_{j},\phi^{\prime}_{i})$ and for all $\phi^{\prime}_{i}\in\operatorname*{span}(\pi_{i})$ and $\phi^{\prime\prime}_{j}\in\pi_{j}$, $\frac{\delta^{2}X^{t,x}}{\delta\phi_{i}\delta\phi_{j}}(\phi;\phi^{\prime}_{i},\phi^{\prime\prime}_{j}-\phi_{j})=\mathcal{H}^{\infty}\text{-}\partial_{\varepsilon}\frac{\delta\alpha^{t,x}}{\delta\phi_{i}}((\phi_{j}+\varepsilon(\phi^{\prime\prime}_{j}-\phi_{j}),\phi_{-j});\phi^{\prime}_{i})\big{|}_{\varepsilon=0}$.
Proofs of Theorem 4.4 Items 1 and 2.
Fix $\phi\in\pi^{(N)}$. The linearity of $\frac{\delta V^{t,x}_{i}}{\delta\phi_{j}}(\phi;\cdot):\operatorname*{span}(\pi_{j})\rightarrow{\mathbb{R}}$ follows from the linearity of $\frac{\delta X^{t,x}}{\delta\phi_{j}}(\phi;\cdot)$ and $\frac{\delta\alpha^{t,x}}{\delta\phi_{j}}(\phi;\cdot)$ shown in Lemmas A.1 and A.3. Let $\phi^{\prime}_{j}\in\pi_{j}$ and for each $\varepsilon\in[0,1)$, let $\phi^{\varepsilon}=(\phi_{j}+\varepsilon(\phi^{\prime}_{j}-\phi_{j}),\phi_{-j})$. Recall that
$V^{t,x}_{i}(\phi^{\varepsilon})={\mathbb{E}}\left[\int_{t}^{T}f_{i}(s,X^{t,x,\phi^{\varepsilon}}_{s},\alpha^{t,x,\phi^{\varepsilon}}_{s}){\mathrm{d}}s+g_{i}(X^{t,x,\phi^{\varepsilon}}_{T})\right].$
Then [25, Theorem 9, p. 97], the $\mathcal{S}^{\infty}$-differentiability of $X^{t,x,\phi^{\varepsilon}}$ and the $\mathcal{H}^{\infty}$-differentiability of $\alpha^{t,x,\phi^{\varepsilon}}$ (see Lemmas A.1, A.2 and A.3) imply that $\frac{{\mathrm{d}}}{{\mathrm{d}}\varepsilon}V^{t,x}_{i}(\phi^{\varepsilon})|_{\varepsilon=0}=\frac{\delta V^{t,x}_{i}}{\delta\phi_{j}}(\phi;\phi^{\prime}_{j}-\phi_{j})$, hence Theorem 4.4 Item 1.
To prove Theorem 4.4 Item 2, for all $(k,\ell)\in\{(i,j),(j,i)\}$, observe that the bilinearity of $\frac{\delta^{2}V^{t,x}_{i}}{\delta\phi_{k}\delta\phi_{\ell}}:\pi^{(N)}\times\operatorname*{span}(\pi_{k})\times\operatorname*{span}(\pi_{\ell})\rightarrow{\mathbb{R}}$ follows from (2), the linearity of $\frac{\delta X^{t,x}}{\delta\phi_{k}}(\phi;\cdot)$ and $\frac{\delta\alpha^{t,x}}{\delta\phi_{\ell}}(\phi;\cdot)$, and the bilinearity of $\frac{\delta^{2}X^{t,x}}{\delta\phi_{k}\phi_{\ell}}(\phi;\cdot,\cdot)$ and $\frac{\delta^{2}\alpha^{t,x}}{\delta\phi_{k}\phi_{\ell}}(\phi;\cdot,\cdot)$. The facts that $\frac{\delta^{2}V^{t,x}_{i}}{\delta\phi_{i}\delta\phi_{j}}$ is a linear derivative of $\frac{\delta V^{t,x}_{i}}{\delta\phi_{i}}$ in $\pi_{j}$ and that $\frac{\delta^{2}V^{t,x}_{i}}{\delta\phi_{j}\delta\phi_{i}}$ is a linear derivative of $\frac{\delta V^{t,x}_{i}}{\delta\phi_{j}}$ in $\pi_{i}$ can be proved by similar arguments as those employed for the first-order derivatives. This finishes the proof. ∎
Proof of Theorem 4.4 Item 3.
We only prove the characterization of CLPGs, which follows by applying Theorem 4.3 to the game $\mathcal{G}_{\rm prob}$ with fixed $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$. The convexity of $\pi_{i}$ follows from the convexity of $A_{i}$ in (H.1), the linear differentiability of $(V_{i})_{i\in I_{N}}$ has been proved in Theorem 4.4 Items 1 and 2, and Condition 3 in Theorem 4.3 has been imposed in Condition (4.13). Hence it remains to verify Conditions 1 and 2 in Theorem 4.3 for given $i,j\in I_{N}$, $z,\phi\in\pi^{(N)}$, $\phi^{\prime}_{i},\tilde{\phi}^{\prime}_{i}\in\pi_{i}$ and ${\phi}^{\prime\prime}_{j}\in\pi_{j}$.
To verify Condition 1, for all $r,\varepsilon\in[0,1]$, let $\phi^{r,\varepsilon}=z+r((\phi_{i}+\varepsilon(\tilde{\phi}^{\prime}_{i}-\phi_{i}),\phi_{-i})-z)$. To prove $\sup_{r,\varepsilon\in[0,1]}\left|\frac{\delta^{2}V_{i}^{t,x}}{\delta\phi_{i}\delta\phi_{j}}\big{(}\phi^{r,\varepsilon};{\phi}^{\prime}_{i},{\phi}^{\prime\prime}_{j}\big{)}\right|<\infty$, by (2) and the polynomial growth of $f_{i}$ and $g_{i}$ and their derivatives (see (H.1)), it suffices to show for all $q\geq 1$, there exists $C_{q}$ such that for all $r,\varepsilon\in[0,1]$,
$\|X^{t,x,\phi^{r,\varepsilon}}\|_{\mathcal{S}^{q}},\left\|\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi^{r,\varepsilon};\phi^{\prime}_{i})\right\|_{\mathcal{S}^{q}},\left\|\frac{\delta X^{t,x}}{\delta\phi_{j}}(\phi^{r,\varepsilon};\phi^{\prime\prime}_{j})\right\|_{\mathcal{S}^{q}},\left\|\frac{\delta^{2}X^{t,x}}{\delta\phi_{i}\phi_{j}}(\phi^{r,\varepsilon};\phi^{\prime}_{i},\phi^{\prime\prime}_{j})\right\|_{\mathcal{S}^{q}}\leq C_{q},$ (A.1)
and
$\|\alpha^{t,x,\phi^{r,\varepsilon}}\|_{\mathcal{H}^{q}},\left\|\frac{\delta\alpha^{t,x}}{\delta\phi_{i}}(\phi^{r,\varepsilon};\phi^{\prime}_{i})\right\|_{\mathcal{H}^{q}},\left\|\frac{\delta\alpha^{t,x}}{\delta\phi_{j}}(\phi^{r,\varepsilon};\phi^{\prime\prime}_{j})\right\|_{\mathcal{H}^{q}},\left\|\frac{\delta^{2}\alpha^{t,x}}{\delta\phi_{i}\phi_{j}}(\phi^{r,\varepsilon};\phi^{\prime}_{i},\phi^{\prime\prime}_{j})\right\|_{\mathcal{H}^{q}}\leq C_{q}.$ (A.2)
Since $\phi^{r,\varepsilon}$ is a convex combination of policies $\phi$, $z$ and $(\tilde{\phi}^{\prime}_{i},\phi_{-i})$, there exists $L\geq 0$ such that for all $r,\varepsilon\in[0,1]$,
$\sup_{(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}}({|\phi^{r,\varepsilon}(t,0)|}+|(\partial_{x}\phi^{r,\varepsilon})(t,x)|+|(\partial_{xx}\phi^{r,\varepsilon})(t,x)|)\leq L.$
Hence, the uniform moment estimate (A.1) follows by applying standard moment estimates of SDEs (see [47, Theorem 3.4.3]) to (4.5), (4.6) and (4.7). The estimate (A.2) then follows from (4.9), (4.10) and the Cauchy-Schwarz inequality. This verifies Condition 1.
To verify Condition 2, for all $\varepsilon=(\varepsilon_{i})_{i=1}^{N}\in[0,1]^{N}$, let $\phi^{\varepsilon}=(z_{i}+\varepsilon_{i}({\phi}_{i}-z_{i}))_{i\in I_{N}}$. Observe that $\phi^{\varepsilon}_{i}$ is a convex combination of $\phi_{i}$ and $z_{i}$, there exists $L\geq 0$ such that for all $\varepsilon\in[0,1]^{N}$,
$\sup_{(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}}({|\phi^{\varepsilon}(t,0)|}+|(\partial_{x}\phi^{\varepsilon})(t,x)|+|(\partial_{xx}\phi^{\varepsilon})(t,x)|)\leq L,$
from which one can deduce uniform moment estimates of the state and control processes as in (A.1) and (A.2). As $\lim_{\varepsilon_{i}\downarrow 0}\phi^{\varepsilon}_{i}=\phi^{0}_{i}$ a.e., by stability results of SDEs (e.g., [47, Theorem 3.4.2]),
$\lim_{\varepsilon\rightarrow 0}\left(X^{t,x,\phi^{\varepsilon}},\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi^{\varepsilon};\phi^{\prime}_{i}),\frac{\delta X^{t,x}}{\delta\phi_{j}}(\phi^{\varepsilon};\phi^{\prime\prime}_{j}),\frac{\delta^{2}X^{t,x}}{\delta\phi_{i}\phi_{j}}(\phi^{\varepsilon};\phi^{\prime}_{i},\phi^{\prime\prime}_{j})\right)$
$\quad=\left(X^{t,x,\phi^{0}},\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi^{0};\phi^{\prime}_{i}),\frac{\delta X^{t,x}}{\delta\phi_{j}}(\phi^{0};\phi^{\prime\prime}_{j}),\frac{\delta^{2}X^{t,x}}{\delta\phi_{i}\phi_{j}}(\phi^{0};\phi^{\prime}_{i},\phi^{\prime\prime}_{j})\right)\quad\textnormal{a.e.~{}$[t,T]\times\Omega$.}$
By (4.9), (4.10) and the convergence of $\phi^{\varepsilon}$, similar convergence results also hold for the control processes. Hence by Lebesgue’s dominated convergence theorem, $\lim_{\varepsilon\rightarrow 0}\frac{\delta^{2}V_{i}^{t,x}}{\delta\phi_{i}\phi_{j}}(\phi^{\varepsilon};\phi^{\prime}_{i},\phi^{\prime\prime}_{j})=\frac{\delta^{2}V_{i}^{t,x}}{\delta\phi_{i}\phi_{j}}(\phi^{0};\phi^{\prime}_{i},\phi^{\prime\prime}_{j})$. This verifies Condition 2. ∎
A.2 Proof of Theorem 4.6
Proof of Theorem 4.6.
Fix $i,h,\ell\in I_{N}$, $\phi\in\pi^{(N)}$, $\phi^{\prime}_{h}\in\pi_{h}$, $\phi^{\prime\prime}_{\ell}\in\pi_{\ell}$ and $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$. Denote by $X$ the state process $X^{t,x,\phi}$. We first prove the equivalence of first-order linear derivatives. Applying Itô’s formula to $s\mapsto\frac{1}{2}X_{s}^{\top}\Theta^{h}_{i}(s)X_{s}+\theta^{h}_{i}(s)$ gives that
$\left(\frac{1}{2}X_{T}^{\top}\Theta^{h}_{i}(T)X_{T}+\theta^{h}_{i}(T)\right)-\left(\frac{1}{2}x^{\top}\Theta^{h}_{i}(t)x+\theta^{h}_{i}(t)\right)$
$\quad={\mathbb{E}}\left[\int_{t}^{T}\left(\frac{1}{2}X_{s}^{\top}\dot{\Theta}^{h}_{i}X_{s}+\dot{\theta}^{h}+\frac{1}{2}\langle(\Theta^{h}_{i}+(\Theta^{h}_{i})^{\top})X_{s},(A+BK)X_{s}\rangle+\frac{1}{2}\textnormal{tr}\left(\sigma\sigma^{\top}\Theta^{h}_{i}\right)\right){\mathrm{d}}s\right],$
which along with (4.32b) and (4.33b) shows that
${\frac{\delta\overline{V}^{\phi}_{i}}{\delta\phi_{h}}}(t,x;\phi^{\prime}_{h})$ $=\frac{1}{2}{\mathbb{E}}\left[\int_{t}^{T}X_{s}^{\top}\left(\Psi_{i}B_{h}K^{\prime}_{h}+(B_{h}K^{\prime}_{h})^{\top}\Psi_{i}+K^{\top}(R_{i})_{h}K^{\prime}_{h}+(K^{\prime}_{h})^{\top}\big{(}(R_{i})_{h}\big{)}^{\top}K\right)X_{s}{\mathrm{d}}s\right]$
$={\mathbb{E}}\left[\int_{t}^{T}\left(X_{s}^{\top}\Psi_{i}B_{h}K^{\prime}_{h}X_{s}+(KX_{s})^{\top}(R_{i})_{h}K^{\prime}_{h}X_{s}\right){\mathrm{d}}s\right],$
where we used the symmetry of $\Psi_{i}(\cdot)$. Then by (4.28),
$\begin{split}&\frac{\delta V^{t,x}_{i}}{\delta\phi_{h}}(\phi;\phi^{\prime}_{h})-{\frac{\delta\overline{V}^{\phi}_{i}}{\delta\phi_{h}}}(t,x;\phi^{\prime}_{h})\\ &\quad={\mathbb{E}}\bigg{[}\int_{t}^{T}\bigg{(}X_{s}^{\top}Q_{i}Y^{h}_{s}+(KX_{s})^{\top}R_{i}KY^{h}_{s}-X_{s}^{\top}\Psi_{i}B_{h}K^{\prime}_{h}X_{s}\bigg{)}{\mathrm{d}}s+X_{T}^{\top}G_{i}Y^{h}_{T}\bigg{]}.\end{split}$ (A.3)
Applying Itô’s formula to $s\mapsto X_{s}^{\top}\Psi_{i}(s)Y^{h}_{s}$ and using (4.29) and (4.33a) imply that
$\begin{split}{\mathbb{E}}[X_{T}^{\top}G_{i}Y^{h}_{T}]&={\mathbb{E}}\bigg{[}\int_{t}^{T}\bigg{(}X_{s}^{\top}\dot{\Psi}_{i}Y^{h}_{s}+\langle\Psi_{i}X_{s},(A+BK)Y^{h}_{s}+B_{h}K^{\prime}_{h}X_{s}\rangle+\langle\Psi_{i}Y^{h}_{s},(A+BK)X_{s}\rangle\bigg{)}{\mathrm{d}}s\bigg{]}\\ &=-{\mathbb{E}}\bigg{[}\int_{t}^{T}\bigg{(}X_{s}^{\top}(Q_{i}+K^{\top}R_{i}K)Y^{h}_{s}-X_{s}^{\top}\Psi_{i}B_{h}K^{\prime}_{h}X_{s}\bigg{)}{\mathrm{d}}s\bigg{]}.\end{split}$
This along with (A.3) proves $\frac{\delta V^{t,x}_{i}}{\delta\phi_{h}}(\phi;\phi^{\prime}_{h})-{\frac{\delta\overline{V}^{\phi}_{i}}{\delta\phi_{h}}}(t,x;\phi^{\prime}_{h})=0$.
Next we prove the equivalence of second-order linear derivatives. By (4.30) and $R_{i}E_{h}=(R_{i})_{h}$,
$\begin{split}\frac{\delta^{2}V^{t,x}_{i}}{\delta\phi_{h}\delta\phi_{\ell}}(\phi;\phi^{\prime}_{h},\phi^{\prime\prime}_{\ell})&={\mathbb{E}}\bigg{[}\int_{t}^{T}\bigg{(}Y^{\ell}_{s}Q_{i}Y^{h}_{s}+(KY^{\ell}_{s})^{\top}R_{i}KY^{h}_{s}\\ &\quad+(KY^{h}_{s})^{\top}(R_{i})_{\ell}K^{\prime\prime}_{\ell}X_{s}+(KY^{\ell}_{s})^{\top}(R_{i})_{h}K^{\prime}_{h}X_{s}+(K^{\prime\prime}_{\ell}X_{s})^{\top}(R_{i})_{\ell h}K^{\prime}_{h}X_{s}\\ &\quad+X^{\top}_{s}Q_{i}Z_{s}+(KX_{s})^{\top}\Big{(}R_{i}KZ_{s}+(R_{i})_{\ell}K^{\prime\prime}_{\ell}Y^{h}_{s}+(R_{i})_{h}K^{\prime}_{h}Y^{\ell}_{s}\Big{)}\bigg{)}{\mathrm{d}}s\bigg{]}\\ &\quad+{\mathbb{E}}\left[(Y^{\ell}_{T})^{\top}G_{i}Y^{h}_{T}+X^{\top}_{T}G_{i}Z_{T}\right].\end{split}$ (A.4)
Applying Itô’s formula to $s\mapsto(Y^{\ell}_{s})^{\top}\Psi_{i}(s)Y^{h}_{s}+X^{\top}_{s}\Psi_{i}(s)Z_{s}$ and using (4.29), (4.31) and (4.33a),
${\mathbb{E}}\left[(Y^{\ell}_{T})^{\top}G_{i}Y^{h}_{T}+X^{\top}_{T}G_{i}Z_{T}\right]$
$\quad={\mathbb{E}}\left[\int_{t}^{T}\left(\langle\Psi_{i}Y^{\ell}_{s},{\mathrm{d}}Y^{h}_{s}\rangle+\langle{\mathrm{d}}Y^{\ell}_{s},\Psi_{i}Y^{h}_{s}\rangle+\langle\Psi_{i}X_{s},{\mathrm{d}}Z_{s}\rangle+\langle{\mathrm{d}}X_{s},\Psi_{i}Z_{s}\rangle\right){\mathrm{d}}s\right]$
$\qquad+{\mathbb{E}}\left[\int_{t}^{T}\left((Y^{\ell}_{s})^{\top}\dot{\Psi}_{i}Y^{h}_{s}+X_{s}^{\top}\dot{\Psi}_{i}Z_{s}\right){\mathrm{d}}s\right]$
$\quad={\mathbb{E}}\left[\int_{t}^{T}\left((Y^{\ell}_{s})^{\top}\Psi_{i}B_{h}K^{\prime}_{h}X_{s}+(Y^{h}_{s})^{\top}\Psi_{i}B_{\ell}K^{\prime\prime}_{\ell}X_{s}+X_{s}^{\top}\Psi_{i}(B_{\ell}K^{\prime\prime}_{\ell}Y^{h}_{s}+B_{h}K^{\prime}_{h}Y^{\ell}_{s})\right){\mathrm{d}}s\right]$
$\qquad-{\mathbb{E}}\left[\int_{t}^{T}\left((Y^{\ell}_{s})^{\top}(Q_{i}+K^{\top}R_{i}K)Y^{h}_{s}+X_{s}^{\top}(Q_{i}+K^{\top}R_{i}K)Z_{s}\right){\mathrm{d}}s\right].$
Substituting this into (A.4) implies that
$\begin{split}&\frac{\delta^{2}V^{t,x}_{i}}{\delta\phi_{h}\delta\phi_{\ell}}(\phi;\phi^{\prime}_{h},\phi^{\prime\prime}_{\ell})\\ &\quad={\mathbb{E}}\bigg{[}\int_{t}^{T}\bigg{(}(Y^{\ell}_{s})^{\top}\left(\Psi_{i}B_{h}K^{\prime}_{h}+(B_{h}K^{\prime}_{h})^{\top}\Psi_{i}+K^{\top}(R_{i})_{h}K^{\prime}_{h}+(K^{\prime}_{h})^{\top}((R_{i})_{h})^{\top}K\right)X_{s}\\ &\qquad+(Y^{h}_{s})^{\top}\left(\Psi_{i}B_{\ell}K^{\prime\prime}_{\ell}+(B_{\ell}K^{\prime\prime}_{\ell})^{\top}\Psi_{i}+K^{\top}(R_{i})_{\ell}K^{\prime\prime}_{\ell}+(K^{\prime\prime}_{\ell})^{\top}((R_{i})_{\ell})^{\top}K\right)X_{s}\\ &\qquad+(K^{\prime\prime}_{\ell}X_{s})^{\top}(R_{i})_{\ell h}K^{\prime}_{h}X_{s}\bigg{)}{\mathrm{d}}s\bigg{]}.\end{split}$ (A.5)
Meanwhile, applying Itô’s formula to $s\mapsto\frac{1}{2}X_{s}^{\top}\Lambda^{h,\ell}_{i}(s)X_{s}+\lambda^{h,\ell}_{i}(s)$ and using (4.32c) and (4.33c) give
$\begin{split}{\frac{\delta^{2}\overline{V}^{\phi}_{i}}{\delta\phi_{h}\delta\phi_{\ell}}}(t,x;\phi^{\prime}_{h},\phi^{\prime\prime}_{\ell})&={\mathbb{E}}\left[\int_{t}^{T}\left(X_{s}^{\top}\Theta^{h}_{i}B_{\ell}K^{\prime\prime}_{\ell}X_{s}+X_{s}^{\top}\Theta^{\ell}_{i}B_{h}K^{\prime}_{h}X_{s}+(K^{\prime\prime}_{\ell}X_{s})^{\top}(R_{i})_{\ell h}K^{\prime}_{h}X_{s}\right){\mathrm{d}}s\right].\end{split}$ (A.6)
Finally, applying Itô’s formula to $s\mapsto(Y^{h}_{s})^{\top}\Theta^{\ell}_{i}(s)X_{s}+(Y^{\ell}_{s})^{\top}\Theta^{h}_{i}(s)X_{s}$ and using (4.29) yield
00 0 $={\mathbb{E}}\bigg{[}\int_{t}^{T}\bigg{(}(Y^{h}_{s})^{\top}\dot{\Theta}^{\ell}_{i}X_{s}+(Y^{\ell}_{s})^{\top}\dot{\Theta}^{h}_{i}X_{s}+\langle\Theta^{\ell}_{i}X_{s},(A+BK)Y^{h}_{s}+B_{h}K^{\prime}_{h}X_{s}\rangle$
$\quad+\langle\Theta^{\ell}_{i}Y^{h}_{s},(A+BK)X_{s}\rangle+\langle\Theta^{h}_{i}X_{s},(A+BK)Y^{\ell}_{s}+B_{\ell}K^{\prime\prime}_{\ell}X_{s}\rangle+\langle\Theta^{h}_{i}Y^{\ell}_{s},(A+BK)X_{s}\rangle\bigg{)}{\mathrm{d}}s\bigg{]}$
$={\mathbb{E}}\bigg{[}\int_{t}^{T}\bigg{(}(Y^{h}_{s})^{\top}\left(\dot{\Theta}^{\ell}_{i}+(A+BK)^{\top}\Theta^{\ell}_{i}+\Theta^{\ell}_{i}(A+BK)\right)X_{s}$
$\quad+(Y^{\ell}_{s})^{\top}\left(\dot{\Theta}^{h}_{i}+(A+BK)^{\top}\Theta^{h}_{i}+\Theta^{h}_{i}(A+BK)\right)X_{s}+X^{\top}_{s}\Theta^{\ell}_{i}B_{h}K^{\prime}_{h}X_{s}+X^{\top}_{s}\Theta^{h}_{i}B_{\ell}K^{\prime\prime}_{\ell}X_{s}\bigg{)}{\mathrm{d}}s\bigg{]},$
which along with (4.33b), (A.5) and (A.6) implies that $\frac{\delta^{2}V^{t,x}_{i}}{\delta\phi_{h}\delta\phi_{\ell}}(\phi;\phi^{\prime}_{h},\phi^{\prime\prime}_{\ell})={\frac{\delta^{2}\overline{V}^{\phi}_{i}}{\delta\phi_{h}\delta\phi_{\ell}}}(t,x;\phi^{\prime}_{h},\phi^{\prime\prime}_{\ell})$.
We then prove the equivalence of Conditions (4.13) and (4.36). By (4.35), it is clear that (4.36) implies (4.13). On the other hand, if (4.13) holds, by (4.34) and (4.35),
$x^{\top}(\Lambda^{i,j}_{i}(t)-\Lambda^{j,i}_{j}(t))x+\int_{t}^{T}\textnormal{tr}\left(\sigma\sigma^{\top}(\Lambda^{i,j}_{i}(s)-\Lambda^{j,i}_{j}(s))\right){\mathrm{d}}s=0\quad\forall(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}},$
which along with $\operatorname*{span}\{xx^{\top}|x\in{\mathbb{R}}^{n_{x}}\}={\mathbb{S}}^{n_{x}}$ (which can be shown by the eigenvalue decomposition) implies that $\Lambda^{i,j}_{i}(T)-\Lambda^{j,i}_{j}(T)=0$ and $\dot{\Lambda}^{i,j}_{i}-\dot{\Lambda}^{j,i}_{j}+\Lambda^{i,j}_{i}-\Lambda^{j,i}_{j}=0$. This shows that (4.36) holds.
Finally, by (4.34), it is clear that (4.37) implies (4.36) for all $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$, which shows that $\mathcal{G}_{\rm LQ}$ is an MPG. Conversely, suppose that $\mathcal{G}_{\rm LQ}$ is an MPG. Using the fact the policy class $\pi_{i}$ in (4.24) is a vector space and the mean-value theorem, one can show that (4.36) must hold for all $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$, which implies that $x^{\top}(\Lambda^{i,j}_{i}(t)-\Lambda^{j,i}_{j}(t))x=0$ for all $(t,x)\in[0,T]\times{\mathbb{R}}^{n_{x}}$. This along with $\Lambda^{i,j}_{i}(t)-\Lambda^{j,i}_{j}(t)\in{\mathbb{S}}^{n_{x}}$ implies that $\Lambda^{i,j}_{i}(t)=\Lambda^{j,i}_{j}(t)$ for all $t\in[0,T]$. ∎
Appendix B Proofs of Technical Lemmas
B.1 Proofs of Lemmas 4.1 and 4.2
Proof of Lemma 4.1.
By Definition 4.1 and the linear differentiability of $f$, $[0,1]\ni\varepsilon\mapsto f(\phi^{\varepsilon})\in{\mathbb{R}}$ admits a right-hand derivative $\frac{\delta f}{\delta\phi_{i}}(\phi;\phi^{\prime}_{i}-\phi_{i})$ at $\varepsilon=0$. We now prove the differentiability of $\varepsilon\mapsto f(\phi^{\varepsilon})$ at $\varepsilon_{0}\in(0,1)$. For all $\varepsilon>0$ such that $\varepsilon_{0}+\varepsilon<1$,
$\phi^{\varepsilon_{0}+\varepsilon}=(\phi^{\varepsilon_{0}}_{i}+\varepsilon(\phi^{\prime}_{i}-\phi_{i}),\phi_{-i})=\left(\phi^{\varepsilon_{0}}_{i}+\frac{\varepsilon}{1-\varepsilon_{0}}(\phi^{\prime}_{i}-\phi^{\varepsilon_{0}}_{i}),\phi_{-i}\right).$
Thus by the linear differentiability of $f$,
$\lim_{\varepsilon\downarrow 0}\frac{f(\phi^{\varepsilon_{0}+\varepsilon})-f(\phi^{\varepsilon_{0}})}{\varepsilon}$ $=\frac{1}{1-\varepsilon_{0}}\lim_{\varepsilon\downarrow 0}\frac{f\left(\left(\phi^{\varepsilon_{0}}_{i}+\frac{\varepsilon}{1-\varepsilon_{0}}(\phi^{\prime}_{i}-\phi^{\varepsilon_{0}}_{i}),\phi_{-i}\right)\right)-f(\phi^{\varepsilon_{0}})}{\varepsilon/(1-\varepsilon_{0})}$
$=\frac{1}{1-\varepsilon_{0}}\frac{\delta f}{\delta\phi_{i}}(\phi^{\varepsilon_{0}};\phi^{\prime}_{i}-\phi^{\varepsilon_{0}}_{i})=\frac{\delta f}{\delta\phi_{i}}(\phi^{\varepsilon_{0}};\phi^{\prime}_{i}-\phi_{i}),$
where the last identity used the linearity of $\frac{\delta f}{\delta\phi_{i}}(\phi^{\varepsilon_{0}};\cdot):\operatorname*{span}(\pi_{i})\rightarrow{\mathbb{R}}$. On the other hand, for all $\varepsilon>0$ such that $\varepsilon_{0}-\varepsilon>0$, $\phi^{\varepsilon_{0}-\varepsilon}=(\phi^{\varepsilon_{0}}_{i}-\varepsilon(\phi^{\prime}_{i}-\phi_{i}),\phi_{-i})=\left(\phi^{\varepsilon_{0}}_{i}+\frac{\varepsilon}{\varepsilon_{0}}(\phi_{i}-\phi^{\varepsilon_{0}}_{i}),\phi_{-i}\right).$ Hence by the linear differentiability of $f$,
$\lim_{\varepsilon\downarrow 0}\frac{f(\phi^{\varepsilon_{0}-\varepsilon})-f(\phi^{\varepsilon_{0}})}{-\varepsilon}$ $=-\frac{1}{\varepsilon_{0}}\lim_{\varepsilon\downarrow 0}\frac{f\left(\left(\phi^{\varepsilon_{0}}_{i}+\frac{\varepsilon}{\varepsilon_{0}}(\phi_{i}-\phi^{\varepsilon_{0}}_{i}),\phi_{-i}\right)\right)-f(\phi^{\varepsilon_{0}})}{\varepsilon/\varepsilon_{0}}$
$=-\frac{1}{\varepsilon_{0}}\frac{\delta f}{\delta\phi_{i}}(\phi^{\varepsilon_{0}};\phi_{i}-\phi^{\varepsilon_{0}}_{i})=\frac{\delta f}{\delta\phi_{i}}(\phi^{\varepsilon_{0}};\phi^{\prime}_{i}-\phi_{i}).$
This proves the differentiability of $\varepsilon\mapsto f(\phi^{\varepsilon})$ at $\varepsilon_{0}\in(0,1)$. Similarly, $\varepsilon\mapsto f(\phi^{\varepsilon})$ admits a left-hand derivative at $1$. ∎
Proof of Lemma 4.2.
Let $z,\phi\in\pi^{(N)}$ be fixed. For simplicity, we assume $N=2$ and $I_{N}=\{i,j\}$, as the same argument can be easily extended to $N\geq 3$.
We first prove the differentiability of $r\mapsto f(z+r(\phi-z))$ admits a right-hand derivative at 00. For each $r\in[0,1]$, let $\phi^{r}=z+r(\phi-z)$. Recall that by Lemma 4.1, for all $r\in[0,1]$, $\varepsilon\mapsto f((\phi^{r}_{-j},z_{j}+\varepsilon(\phi_{j}-z_{j})))$ is differentiable on $[0,1]$ and
$\frac{{\mathrm{d}}}{{\mathrm{d}}\varepsilon}f((\phi^{r}_{-j},z_{j}+\varepsilon(\phi_{j}-z_{j})))=\frac{\delta f}{\delta\phi_{j}}((\phi^{r}_{-j},z_{j}+\varepsilon(\phi_{j}-z_{j}));\phi_{j}-z_{j})\quad\forall\varepsilon\in(0,1).$
Then for all $r\in(0,1)$, using the fact that $\phi^{r}_{-i}=\phi^{r}_{j}$ (note that $I_{N}=\{i,j\}$),
$f(\phi^{r})-f(z)$ $=f(\phi^{r})-f((\phi^{r}_{i},z_{-i}))+f((\phi^{r}_{i},z_{-i}))-f(z)$
$=\frac{\delta f}{\delta\phi_{j}}((z_{i}+r(\phi_{i}-z_{i}),z_{j}+r\varepsilon_{r}(\phi_{j}-z_{j}));\phi_{j}-z_{j})r+f((\phi^{r}_{i},z_{-i}))-f(z),$
for some $\varepsilon_{r}\in(0,1)$, where the last identity used the mean value theorem. Dividing both sides of the above identity by $r$ and letting $r\rightarrow 0$ yield
$\lim_{r\downarrow 0}\frac{f(\phi^{r})-f(z)}{r}$ $=\lim_{r\downarrow 0}\frac{\delta f}{\delta\phi_{j}}((z_{i}+r(\phi_{i}-z_{i}),z_{j}+r\varepsilon_{r}(\phi_{j}-z_{j}));\phi_{j}-z_{j})+\frac{\delta f}{\delta\phi_{i}}(z;\phi_{i}-z_{i})$
$=\frac{\delta f}{\delta\phi_{j}}(z;\phi_{j}-z_{j})+\frac{\delta f}{\delta\phi_{i}}(z;\phi_{i}-z_{i}),$
where the last identity used the continuity of ${\varepsilon}\mapsto\frac{\delta f}{\delta\phi_{j}}(z+\varepsilon\cdot(\phi-z);\phi_{j}-z_{j})$ at 00, which holds due to the continuity assumption of $\frac{\delta f}{\delta\phi_{j}}$ and the linearity of $\frac{\delta f}{\delta\phi_{j}}$ with respect to the last argument. This proves the desired differentiability of $r\mapsto f(\phi^{r})$ at 00. Similarly, $r\mapsto f(\phi^{r})$ is differentiable at $r\in(0,1]$. ∎
B.2 Proofs of Lemmas A.1, A.2 and A.3
Proof of Lemma A.1.
Since $B$, $\Sigma$ and $\phi$ have bounded first-order derivatives and $\phi^{\prime}_{i}\in\operatorname*{span}(\pi_{i})$ is of linear growth in $x$, standard a-priori estimates of SDEs (see e.g., [47, Theorem 3.4.3]) show that (4.6) admits a unique solution $\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi;\phi^{\prime}_{i})\in\mathcal{S}^{\infty}([t,T];{\mathbb{R}}^{n_{x}})$. As the terms $(\partial_{a_{i}}B^{\phi})[\phi^{\prime}_{i}](s,X^{t,x,\phi}_{s})$ and $(\partial_{a_{i}}\Sigma^{\phi})[\phi^{\prime}_{i}](s,X^{t,x,\phi}_{s})$ are linear with respect to $\phi^{\prime}_{i}$, one can easily verify by the uniqueness of solutions that for all $\alpha,\beta\in{\mathbb{R}}$ and $\phi^{\prime}_{i},\phi_{i}^{\prime\prime}\in\operatorname*{span}(\pi_{i})$, $\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi;\alpha{\phi_{i}^{\prime}}+\beta\phi_{i}^{\prime\prime})=\alpha\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi;{\phi_{i}^{\prime}})++\beta\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi;\phi_{i}^{\prime\prime}).$ This proves the linearity of $\operatorname*{span}(\pi_{i})\ni\phi^{\prime}_{i}\mapsto\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi;\phi^{\prime}_{i})\in\mathcal{S}^{\infty}([t,T];{\mathbb{R}}^{n_{x}})$ in Item 1. Finally, observe that for all $\varepsilon\in[0,1)$, $X^{t,x,(\phi_{i}+\varepsilon(\phi_{i}^{\prime}-\phi_{i}),\phi_{-i})}$ satisfies
${\mathrm{d}}X_{s}$ $=B(s,X_{s},(\phi_{i}+\varepsilon(\phi^{\prime}_{i}-\phi_{i}),\phi_{-i})(s,X_{s})){\mathrm{d}}s$
$\quad+\Sigma(s,X_{s},(\phi_{i}+\varepsilon(\phi^{\prime}_{i}-\phi_{i}),\phi_{-i})(s,X_{s})){\mathrm{d}}W_{s},\quad s\in[t,T];\quad X_{t}=x.$
As $B,\Sigma$, $\phi$ and $\phi^{\prime}_{i}$ are continuous differentiable in $(x,a)$, $\mathcal{S}^{\infty}\text{-}\partial_{\varepsilon}X^{t,x,(\phi_{i}+\varepsilon(\phi_{i}^{\prime}-\phi_{i}),\phi_{-i})}\big{|}_{\varepsilon=0}$ exists due to [25, Theorem 4, p. 105] and satisfies (4.6) due to [25, Remark 5, p. 108]. The uniqueness of solutions to (4.6) then yields Item 2. ∎
Proof of Lemma A.2.
Standard well-posedness results of SDEs and the regularity conditions of $B$, $\Sigma$, $\phi$, $\phi^{\prime}_{i}$ and $\phi^{\prime\prime}_{j}$ show that (4.7) admits a unique solution $\frac{\delta^{2}X^{t,x}}{\delta\phi_{i}\delta\phi_{j}}(\phi;\phi^{\prime}_{i},\phi^{\prime\prime}_{j})$.
For Item 1, observe that by Lemma A.1 and (4.8), for any given $\phi^{\prime\prime}_{j}\in\operatorname*{span}(\pi_{j})$, the map $\phi^{\prime}_{i}\mapsto\big{(}F_{B},F_{\Sigma}\big{)}\left(s,X^{t,x,\phi}_{s},\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi;\phi^{\prime}_{i}),\frac{\delta X^{t,x}}{\delta\phi_{j}}(\phi;\phi^{\prime\prime}_{j}),\phi^{\prime}_{i},\phi^{\prime\prime}_{j}\right)$ is linear. This along with the linearity of the SDE (4.7) shows that $\phi^{\prime}_{i}\mapsto\frac{\delta^{2}X^{t,x}}{\delta\phi_{i}\delta\phi_{j}}(\phi;\phi^{\prime}_{i},\phi^{\prime\prime}_{j})$ is linear. Similar arguments yield the linearity of $\phi^{\prime\prime}_{j}\mapsto\frac{\delta^{2}X^{t,x}}{\delta\phi_{i}\delta\phi_{j}}(\phi;\phi^{\prime}_{i},\phi^{\prime\prime}_{j})$.
For Item 2, observe that $F_{B}(t,x,y_{1},y_{2},\phi^{\prime}_{i},\phi^{\prime\prime}_{j})=F_{B}(t,x,y_{2},y_{1},\phi^{\prime\prime}_{j},\phi^{\prime}_{i})$, as $B$ is twice continuously differentiable with respect to $(x,a)$ and hence has a symmetric Jacobian. Similarly, we have $F_{\Sigma}\left(t,x,y_{1},y_{2},\phi^{\prime}_{i},\phi^{\prime\prime}_{j}\right)=F_{\Sigma}\left(t,x,y_{2},y_{1},\phi^{\prime\prime}_{j},\phi^{\prime}_{i}\right)$. This proves that both $\frac{\delta^{2}X^{t,x}}{\delta\phi_{i}\delta\phi_{j}}(\phi;\phi^{\prime}_{i},\phi^{\prime\prime}_{j})$ and $\frac{\delta^{2}X^{t,x}}{\delta\phi_{j}\delta\phi_{i}}(\phi;\phi^{\prime\prime}_{j},\phi^{\prime}_{i})$ satisfy (4.7), which along with the uniqueness of solutions to (4.7) yields Item 2.
Finally, for all $\varepsilon\in[0,1)$, let $\phi^{\varepsilon}=(\phi_{j}+\varepsilon(\phi^{\prime\prime}_{j}-\phi_{j}),\phi_{-j})$. By Lemma A.1, $\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi^{\varepsilon};\phi^{\prime}_{i})$ satisfies
$\begin{split}{\mathrm{d}}Y_{s}&=(\partial_{x}B^{\phi^{\varepsilon}})(s,X^{t,x,\phi^{\varepsilon}}_{s})Y_{s}{\mathrm{d}}s+\big{(}(\partial_{x}\Sigma^{\phi^{\varepsilon}})(s,X^{t,x,{\phi^{\varepsilon}}}_{s})Y_{s}\big{)}{\mathrm{d}}W_{s}\\ &\quad+(\partial_{a_{i}}B^{\phi^{\varepsilon}})[\phi^{\prime}_{i}](s,X^{t,x,{\phi^{\varepsilon}}}_{s}){\mathrm{d}}s+(\partial_{a_{i}}\Sigma^{\phi^{\varepsilon}})[\phi^{\prime}_{i}](s,X^{t,x,{\phi^{\varepsilon}}}_{s}){\mathrm{d}}W_{s}\quad\forall s\in[t,T].\end{split}$
By [25, Theorem 4, p. 105], $\mathcal{S}^{\infty}\text{-}\partial_{\varepsilon}\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi^{\varepsilon};\phi^{\prime}_{i})\big{|}_{\varepsilon=0}$ exists and satisfies (4.7), which along with the uniqueness of solutions to (4.7) yields Item 3. This finishes the proof. ∎
Proof of Lemma A.3.
For Item 1, the linearity of $\frac{\delta\alpha^{t,x}}{\delta\phi_{i}}(\phi;\cdot)$ follows from (4.9) and the linearity of $\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi;\cdot)$ (see Lemma A.1). For each $\varepsilon\in[0,1)$, let $\phi^{\varepsilon}=(\phi_{i}+\varepsilon(\phi^{\prime}_{i}-\phi_{i}),\phi_{-i})$ and observe that
$\alpha^{t,x,\phi^{\varepsilon}}_{s}=\phi^{\varepsilon}(s,X^{t,x,\phi^{\varepsilon}}_{s})=(\phi_{i}+\varepsilon(\phi^{\prime}_{i}-\phi_{i}),\phi_{-i})(s,X^{t,x,\phi^{\varepsilon}}_{s}).$
Hence [25, Theorem 9, p. 97], the differentiability of $\phi$ and $\phi^{\prime}_{i}$ and the $\mathcal{S}^{\infty}$-differentiability of $(X^{t,x,\psi^{\varepsilon}})_{\varepsilon\in[0,1)}$ imply the existence and the desired formula of -$\mathcal{H}^{\infty}\text{-}\partial_{\varepsilon}\alpha^{t,x,(\phi_{i}+\varepsilon(\phi_{i}^{\prime}-\phi_{i}),\phi_{-i})}\big{|}_{\varepsilon=0}$.
For Item 2, the bilinearity and symmetry of $\frac{\delta^{2}\alpha^{t,x}}{\delta\phi_{i}\delta\phi_{j}}(\phi;\cdot,\cdot)$ follow directly from (4.10) and Lemmas A.1 and A.2. For each $\varepsilon\in[0,1)$, let $\phi^{\varepsilon}=(\phi_{j}+\varepsilon(\phi^{\prime}_{j}-\phi_{j}),\phi_{-j})$ and
$\frac{\delta\alpha^{t,x}}{\delta\phi_{i}}(\phi^{\varepsilon};\phi^{\prime}_{i})_{s}$ $=(\partial_{x}\phi^{\varepsilon})(s,X^{t,x,\phi^{\varepsilon}}_{s})\frac{\delta X^{t,x}}{\delta\phi_{i}}(\phi^{\varepsilon};\phi^{\prime}_{i})_{s}+E_{i}\phi^{\prime}_{i}(s,X^{t,x,\phi^{\varepsilon}}_{s}),$ (B.1)
Applying [25, Theorem 9, p. 97] again yields the expression of -$\mathcal{H}^{\infty}\text{-}\partial_{\varepsilon}\frac{\delta\alpha^{t,x}}{\delta\phi_{i}}((\phi_{j}+\varepsilon(\phi^{\prime\prime}_{j}-\phi_{j}),\phi_{-j});\phi^{\prime}_{i})\big{|}_{\varepsilon=0}$. ∎